A Complete Bibliography of Publications in IEEE Transactions on Big Data

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 August 2024 Version 1.16

Title word cross-reference

3 [CZLL19]. c [ZYCL22]. d [CGH^{+}22]. K [GQP^{+}23, GDY24, YWW^{+}21a, CWS^{+}21, LBL^{+}18, XLP^{+}18, ZWLN22]. LU [LZSL21]. N [DZPM20]. QR [LZSL21]. r [LYLJ22].

-Location [LBL^{+}18]. -Means [ZYCL22]. -Nearest [GDY24, YWW^{+}21a]. -NN [CWS^{+}21, ZWLN22]. -Simplexed [CGH^{+}22].

1 [CCD19b, SLV^{+}20, SCPS20, ZBB^{+}20]. 19 [CZQ^{+}21, DZ21, FRS^{+}22, LMPS21, LZS^{+}21, LCWL21, MSRS21, PTL^{+}21, SPYS23, WVD^{+}21, YWG^{+}21].

2 [CCD19c]. 2-Dimension [ZZL^{+}22].

3D [XYX^{+}23].

5G [HWB^{+}20].

8 [CWGA17].

Accelerated [CXT^{+}18, FZC20, Mat16, QLZ24, ZWD^{+}23]. Accelerating [HKKC22, WJW^{+}18]. Acceleration [YJH21]. Access [APK22, BSL^{+}22, HLC^{+}18, XCV22]. Accessible [PTL^{+}21]. Accuracy [XLL^{+}22, XJQ^{+}21]. Accurate [ONP^{+}24]. Achieving [YLC^{+}22]. Acquire [CCD^{+}19a]. Across [SZJ^{+}22]. Action [YLLC18]. Active [XLQ^{+}22, ZYW^{+}23]. Activities [LCS18]. Activity [JFG17, LSH^{+}23, MSRS21, MLK21, YLLC18]. Activity-Based [JFG17]. Ad [SWC^{+}22, TAL^{+}17]. Ad-Hoc [TAL^{+}17]. Adaptable [SGMB17]. Adaptation [YZW^{+}23a]. Adaptive [CGH^{+}22, HVVP21, KKKK16, SZXL22, XWI^{+}21, Yan15a, Yan23, ZGS^{+}22]. Adaptively [QLZ24]. Adaptively-Accelerated [QLZ24]. Addition [ZCSW23]. Addressing [RJS22]. Adjacent [HX22, ZZ23]. Adjoint [LQHC21]. Adopters [WGG^{+}23]. Advanced [ZBB^{+}20]. Adversarial [FWWC23, JLX^{+}23, QYC^{+}23, YZW^{+}23b, ZLZL22, ZLZ^{+}22, ZWD^{+}23, ZYW^{+}23]. Advertising [JFZ^{+}21]. Aesthetic [CLGX23]. AF [YXZL23]. AF-GCN [YXZL23]. Affordance [CKL^{+}23]. Age [LLS^{+}20, MCA^{+}20]. Agent [WHGW23]. Agents [Sak20]. Aggregate [DLZ^{+}19, GLZ^{+}24]. Aggregation [PB23, STM22]. Aggregators [JDC^{+}22]. Agile [CKH16]. Agnostic [YSM^{+}23]. Agricultural [DZS20]. AI [DZ21]. Aid [FZ18]. Air [HLLZ22a, YDL^{+}22, YLL22, YLLC23, YLL23, ZSL17, ZZZ^{+}18]. Aleatoric [WFZ^{+}23]. Algorithm [ADHS21, BLYM19, CWPW20, FLYL21, GSS23, HYLZ22, SPN^{+}22, WL22a, WLD^{+}23, ZYCL22, ZSJ23, ZLT20, ZHKL24]. Algorithms [ADF^{+}17, BTM16, ESC20, Gil16, JY22, LZL^{+}21, SAJP20, SZG^{+}23]. AlgorithmSeer [Gil16]. Alignment [GZS23, WHX^{+}23]. All-Flash [YAG^{+}22]. Allocate [XSYW22]. Allocation [DQWQ18, JHT18, QSC19, SX21, WCW^{+}22, ZCSW23]. Alternate [YCYL24]. Alternating [FZC20]. Alternative [SHL^{+}22]. Altruism [BS21]. ALTRUIST [BED^{+}24]. Altrumetrics [BS21]. AMDECDA [WWYH24]. AMIC [HVVP21]. AMS [Yan15a]. Analyses [RBM^{+}22]. Analysis [CSZ^{+}21, CLZ22, CLGX23, CSW23, DHQ^{+}23, DH18, FMD18, GCA^{+}20, GSS23, HLL^{+}23a, HH18, ICZ20, KAG^{+}22, LCS18, LZMZ20, LTG^{+}24, LZW^{+}23a, LLY^{+}17, LZL^{+}21, MXL^{+}20, MBS^{+}19, MG23, MSC19, PWW^{+}23, PRR22, PIMP17, SHH^{+}21, SHL^{+}22, WCYL21, WOD^{+}18, WWK^{+}21, ZLZ^{+}20, ZCH^{+}18, ZYZ^{+}23]. Analysis-Based [LTG^{+}24, WOD^{+}18]. Analytic [DEG^{+}22, ZSWZ19]. Analytical [RBM^{+}22]. Analytics [ADF^{+}17, CKH16, CGH^{+}22, CQH^{+}18, DLD^{+}20, FRS^{+}22, JGS^{+}19, JHT18, KAG^{+}22, LZH^{+}24, MLQ^{+}19, SJS^{+}21, SVYY16, SLT^{+}22, SWDX20, SZZ^{+}23, WZY^{+}18, WXZ^{+}19, WTM18, Yao15, ZLY^{+}22, ZWC^{+}16]. Analyzing [DLD^{+}17, LWH18, PWH16]. Anchor [FWHM22]. Android [AK19, HZZ^{+}21, ZWH^{+}22]. Annotation [Zhu15]. Anomalous [HHX^{+}19, LZZF23, SLT^{+}22]. Anomaly [HMW^{+}22, LQHC21, MSC19, TLS24, YDM^{+}23, YLB^{+}19, ZLY^{+}22, ZDZC23]. Anonymisation [ZQD^{+}22]. Anonymization [DFF^{+}23]. Answer [LS18a, SLWV17, WH19]. Answering [LLD^{+}20]. AnswerNet [WH19]. Answers [LS18a]. Apache [BTM16, HPWR20, LXLF23, MBG22, SZY^{+}18]. APIs [QHC^{+}22]. Application [ADF^{+}17, SZZ^{+}23, ZLZL22]. Applications [BLYM19, CCD19b, CCD19c, DQWQ18, Far23, FRS^{+}22, HKKC22, HWC20, LXLF23, LZH^{+}24, LB19, MTV21, NGM16, QHC^{+}22, SYv^{+}19, SZA^{+}22, WBS^{+}23, ZWV^{+}19, ZLC^{+}17]. Applied [LMPS21]. Approach [AI23, CGZ23, CKH16, DZS20, DLD^{+}20, GQP^{+}23, HLLZ22a, KJC^{+}18, LWZ^{+}22, LQHC21, LLO^{+}22, LWQ^{+}23, PWH16, RJS22, SAM^{+}20, SLV^{+}20, SLZ^{+}22, WJW^{+}18, WYLG21, WLF^{+}22, WXW^{+}22, XJQ^{+}21, XPC^{+}23, ZLZ^{+}22, ZGS^{+}22, ZS22, ZCSZ23, ZRY^{+}20]. Approaches [LLS^{+}20]. Approximate [MHR^{+}23, WYBH24].

Approximation [DDGU22, SWF16, TZCC22]. Apps [ZWH^{+}22]. Arc [HHX^{+}19]. Architecting [BVAW^{+}16]. Architectural [SBL^{+}22]. Architecture [BLYM19, CKH16]. Architecture-Centric [CKH16]. Architectures [Gou19, HCZ19]. Area [DZS20, MSC19]. Areas [HBLK17]. Argument [LGW^{+}23]. Array [MCM21]. Art [WYLD18]. Article [WWSB16, XLLC16]. Article-Level [WWSB16]. Artistic [MKJ^{+}24]. Aspect [CLL23, HLW^{+}23, WZL^{+}24b]. Aspect-Level [HLW^{+}23]. Assembling [FMD18]. Assessment [Tao15]. Asset [BSL^{+}22]. Assignment [LLF^{+}19]. Assisted [ZZL^{+}22]. Associating [LGW^{+}23]. Association [ICZ20, WWYH24]. ASTROIDE [BZY20]. Astronomical [BZY20]. Astronomy [ZBN^{+}20]. Asynchronous [CYXS23]. ATLAS [WCS^{+}23]. ATrie [STM22]. Attack [ZWD^{+}23]. Attacking [ZWD^{+}23]. Attacks [JLX^{+}23]. Attention [FWWC23, GCL^{+}23, HZF^{+}22, HLF^{+}22, WWYH24, YLZ^{+}22, ZWS^{+}22, ZAL^{+}23, ZWZ^{+}23]. Attentional [WWL^{+}23]. Attentive [WJW24]. Attribute [CDLW19, SLL^{+}21, SSAI21, YXZL23]. Attribute-Based [CDLW19]. Attribute-Fusing [YXZL23]. Attributed [HYLZ22]. Auditing [FYZ^{+}22, LLLZ21, LCY^{+}22, SZC^{+}21]. Augmentation [HLW^{+}23]. Augmented [FWWC23]. Authentication [SLL^{+}21, WYLG21]. Author [XLLC16]. Authorization [SLL^{+}21]. Auto [HZF^{+}22, JKF19]. Auto-Encoder [HZF^{+}22, JKF19]. AutoBDA [SPH23]. Autoencoder [CSZ^{+}21, GCL^{+}23, LWL^{+}23, XWG23a]. Automate [KAB^{+}21]. Automated [SLV^{+}20, XZGW21]. Autonomous [AMAT23, PLZL22, XZGW21]. Averaging [SF23]. Aware [CTC^{+}24, CJY20, DH18, FLYL21, FYZ^{+}22, HD19, HWTM20, HWC20, HGY^{+}23, JHT20, LYLW23, LLY^{+}18, NHL18, SLZ^{+}22, SGMB17, SGM21, SPH23, SBR22, SX21, WFZ^{+}23, WJW24]. Awareness [BHHDA19, MW22, PWW^{+}23, WOD^{+}18].

B4SDC [LDY^{+}22]. Background [IZ19]. Bag [DHQ^{+}23]. Bag-of-Features [DHQ^{+}23]. Balance [LHY^{+}22, QXZ^{+}18]. Balanced [GWLL24]. Bank [LZLJ22]. Bank-Level [LZLJ22]. Based [ADHS21, BSY^{+}24, BS21, BHO21, BTM16, BSL^{+}22, CKH16, CYC^{+}23, CKL^{+}23, CSL^{+}18, CDLW19, DHG21, DHQ^{+}23, FYD^{+}19, FYD^{+}21, GZS23, GRWL23, GCL^{+}23, HHZ^{+}24, HZF^{+}22, HWTM20, HX22, JRP^{+}22, JFG17, JDC^{+}22, JZS^{+}20, KFY^{+}23, LBZ19, LLLZ21, LCWS22, LYLW23, LZC^{+}23, LTG^{+}24, LZD^{+}23, LYC^{+}24, LLW^{+}24, LLC^{+}24, LLY^{+}17, MXL^{+}20, MKJ^{+}24, QXZ^{+}18, QYC^{+}23, QLW^{+}22, SAJP20, SF23, SZC^{+}21, SLWV17, SWC^{+}22, SLL^{+}22, SXS^{+}23, SPN^{+}22, Sun15, TLS24, WSW^{+}18, WWCK22, WXW^{+}22, WLD^{+}23, WWSB16, WTM18, WOD^{+}18, WWK^{+}21, XCV22, XSYW22, YAG^{+}22, YYD^{+}23, YLZ^{+}22, YDL^{+}22, YLZ^{+}19, YLL22, YLLC23, ZZR20, ZHL^{+}17, ZLZ^{+}20, ZZLW21, ZWZ^{+}22, ZS22, ZRL22, ZRLZ23, ZCSW23, ZHL^{+}24, ZGH^{+}22, ZSC^{+}22, ZZY^{+}23, ZRY^{+}20, ZWS21, ZMGIVO20, dSSN^{+}21, CQH^{+}18, WCS^{+}23]. Batch [Gou19]. Bayesian [GWLL24, HLLZ22a, HWW^{+}22]. BCI [HPESZ17]. Be [Cha16, SZSY24]. Behavior [ÇLM17, HHX^{+}19, ZSWZ19]. Behavioral [FZ18]. Behaviors [LZW^{+}23a, SLT^{+}22]. Benchmark [Wan16]. Benchmarking [ESC20]. Benefits [GZW23]. Best [WGG^{+}23]. Beyond [XDZ20]. Bi [YYD^{+}23, ZHL^{+}24]. Bi-Directional [YYD^{+}23]. Bi-Selection [ZHL^{+}24]. Big [Agg15a, Agg15b, Agg16, AI23, APK22, ADHS21, Ano15a, Ano15b, Ano17a, Ano18a, Ano19a, Ano20, Ano21, Ano22, BVAW^{+}16, BHO21, BTM16, BHHDA19, BLYM19, CZG^{+}19, Can16, CKH16, CWPW20, CGH^{+}22, CL23, CCD19b, CLF^{+}18, Cra15, CSL^{+}18, CSW23, DQWQ18, DLZ^{+}19, DFG^{+}19, DLL^{+}16, DH18, ECG^{+}19, FZ18, FYD^{+}19, FYD^{+}21, Fu16, GQZ21, Gil16, GDY24, Gou19, GXQ^{+}20, HD19, HCZ19, HVVP21, HPESZ17, HLC^{+}18, HLH^{+}20, JHT18, JWS^{+}22, KdRFA21, LLLZ21, LCWS22, LZC^{+}23, LZH^{+}24, LLSL19, LTTC16, LTTC17, LW18, LLY^{+}18, LYG^{+}21, LLF^{+}19, MLQ^{+}19, MCM21, PWW^{+}23, PWH16, PIMP17, PWS^{+}19, QXZ^{+}18, RJS22, SAJP20, SZC^{+}21, SSSB16, SCW18, SLL^{+}21, SVYY16, SRM17, SZA^{+}22, SX21, SWDX20, TvMvdH23, TAL19, TRB^{+}21, WTRK16, WQSA17, WYLD18, WZY^{+}18, WXZ^{+}19, WBD^{+}20, WYC^{+}20, WYLG21, WXW^{+}22, WTM18, WS17, WOD^{+}18, WZLS18a, WWK^{+}21, XWBL17, XCV22, XSYW22, YDY^{+}16, Yan15b, YLC^{+}22, Yao15]. Big [Yu15, YWL^{+}22, ZHL^{+}17, ZYYK18, ZWV^{+}19, ZGS^{+}22, ZQD^{+}22, ZYCL22, ZQZ^{+}17, ZCH^{+}18, ZOLP21, ZYX^{+}23, ZSL17, ZLC^{+}17, ZZZ^{+}18, ADOAH19, BZY20, CCD^{+}19a, CW18a, CW18b, CCD19c, CLG^{+}20, DLL^{+}20, DWXZ20, HFX^{+}20, HWB^{+}20, HWTM20, HWC20, LLZ^{+}19, LA20, MHR^{+}23, PA20, RDM^{+}20, XZGW21, XHZW21, YLW^{+}20, ZSK19, ZYCL20, ZBN^{+}20]. Big-Data [APK22, BVAW^{+}16, CL23, DFG^{+}19, RJS22, XHZW21]. Bilinear [YLZ^{+}19]. Billion [JDJ21]. Billion-Scale [JDJ21]. BiLSTM [PP22]. BiLSTM-SSVM [PP22]. BioCyBig [ICZ20]. Biological [ZLZ^{+}20, ZPMT^{+}21]. Biomarkers [ZSHS17]. Biomedical [PTL^{+}21, WBD^{+}20, ZLC^{+}17]. Bipartite [LHWL24]. Bit [WL22b]. BitAnalysis [SXYL23]. Bitcoin [SXYL23]. Bits [LMW21]. BL [YRL^{+}24]. Black [JLX^{+}23, ZWD^{+}23]. Black-Box [JLX^{+}23, ZWD^{+}23]. Blind [GXQ^{+}20]. Blob [WJS^{+}16]. Blob-Filaments [WJS^{+}16]. Block [CYXS23]. Block-Coordinate [CYXS23]. Blockchain [BSL^{+}22, HXW^{+}23, LDY^{+}22, SXS^{+}23]. Blockchain-Based [SXS^{+}23]. Blocking [ESC20]. Boosted [GSS23]. Boosting [HZF^{+}22]. Bounded [LZD^{+}23, ZGC^{+}24]. Box [JLX^{+}23, QWL^{+}23a, ZWD^{+}23]. Brain [YZW^{+}23a, ZYW^{+}23]. Branch [ZGS^{+}22]. Brokerage [QSC19]. Building [QHC^{+}22]. Bundling [CSW23]. Bus [ASYX21]. BusBeat [ASYX21].

Cache [BPP21]. Cache-Locality [BPP21]. Caching [CSL^{+}18, HKKC22]. CaL [LS18b]. Camera [YTY^{+}21]. Cameras [SCW18]. Campaign [YKL^{+}23]. Campaigns [WLZ^{+}20]. Can [Cha16]. Capacity [Mat16]. Care [WVD^{+}21]. Cascaded [MZF24]. Case [CZQ^{+}21, JFG17, LCS18, ZPMT^{+}21]. Catalog [YDM^{+}23]. Categories [CZ17a]. Causal [HGY^{+}23, LLZ^{+}23, MCL22, YWY^{+}22, YCYL24, ZZZ^{+}18]. Causality [ZSL17, ZZZ^{+}18]. CCA [WWCZ22]. Cell [LLCS17]. Cellular [FMD18, MCO^{+}22]. Center [GSS23, ZYW^{+}23]. Centers [JGS^{+}19, LLY^{+}18, SWTX18]. Centric [CKH16, KS22, PWC^{+}22]. cGAIL [ZLZL22]. Chain [YCYL24]. Challenge [NGM16]. Change [CCZ^{+}20, DLD^{+}17]. Change-Point [CCZ^{+}20]. Changes [Kit16, SD22]. Channel [GZL^{+}20, HWB^{+}20, LZMZ20, ZZX^{+}21]. Characterization [CW22]. Characterizing [CSW18, LCWL21]. Check [ZZLW21]. Check-in [ZZLW21]. Checkpoint [MCM21]. Checkpointing [LZLJ22]. Chest [PTL^{+}21, YWG^{+}21]. Chief [Tan20, Yan20]. China [CZQ^{+}21, LZS^{+}21]. Chinese [JZY22]. Choice [WCJI22, ZS22]. Choice-Based [ZS22]. Choices [LLY^{+}17]. Choking [Liu15]. CINTIA [MCM21]. CircRNA [WWYH24]. CircRNA-Disease [WWYH24]. Citation [GM20, WTX^{+}22, WWSB16]. Cities [HBLK17]. City [Kit16, MSRS21]. City-Scale [MSRS21]. CityLines [LLZ^{+}19]. Citywide [YTY^{+}21]. Class [LRL^{+}17, MTV21]. Class-Imbalance [MTV21]. Classification [CXW^{+}23, CYC^{+}23, Far21, GDY24, HXYZ20, HLW^{+}23, LLH^{+}23, LYC^{+}24, LZW^{+}23b, SLV^{+}20, WZZ^{+}19, WWCK22, XLQ^{+}22, XDZ20, XJQ^{+}21, XDH23, YMJ^{+}23, YLZ^{+}19, ZJZ^{+}19]. Classifier [AK19]. Classifiers [GWLL24]. Classify [CZ17a]. Cliques [QLZ^{+}23, QWL^{+}23b]. Closed [HS20]. Cloud [CCD^{+}19a, CWS^{+}21, CDLW19, DQWQ18, FYD^{+}19, FYZ^{+}22, GQZ21, GSS23, HHX^{+}19, HMW^{+}22, LRB^{+}20, LCY^{+}22, LXLF23, LB19, PLZL22, QSC19, SD18, SLV^{+}20, SLL^{+}21, SBL^{+}22, TF22, WLF^{+}22, WXW^{+}22, WTM18, XSYW22, XWZ^{+}23, YDY^{+}16, YZDZ19, YLC^{+}22, ZZR20, ZYYK18, ZYCL22, ZBB^{+}20]. Cloud-Based [WXW^{+}22]. Cloud/Fog [SBL^{+}22]. CloudFinder [RDM^{+}20]. Clouds [CWPW20, HLC^{+}18, JHT18, RDM^{+}20, ZZX^{+}21]. Clustering [BTM16, CTC^{+}24, GMZ^{+}22, GCA^{+}20, HD19, HFX^{+}20, HZY^{+}21, KZL^{+}22, LHWL24, LZMZ20, LZL^{+}22, LZC^{+}23, Liu15, MHR^{+}23, QLW^{+}22, SF23, SSSB16, WLD^{+}23, WWSB16, WZLS18b, XWG23a, YWW^{+}21b, YM24, ZYCL22, ZHKL24, dSSN^{+}21]. Clusters [ANP^{+}23, FY22, HKKC22, HYLZ22, JHT20, WJW^{+}18]. CNN [BHX^{+}23]. CNNs [XDH23]. Co [HZ22]. Co-Segmentation [HZ22]. Coalitional [SX21]. Coarse [GB23]. Coarse-Grained [GB23]. COBO [HCZ^{+}24]. Code [Sun15]. Codes [KGJØ18]. Coding [ZWS21]. Cohesive [LYLJ22]. Cohort [BED^{+}24].

Cold [YKL^{+}23]. Cold-Start [YKL^{+}23]. Collaboration [Can16, LTTC16, LTTC17]. Collaborative [ADOAH19, WTX^{+}22, YMJ^{+}23]. Collaborators [WYXL21]. Collection [CWC^{+}23, LDY^{+}22, PTL^{+}21, WLF^{+}22, YWL^{+}22]. Colocation [SAJP20]. Column [STM22]. Column-Stores [STM22]. Combined [WWYH24]. Comment [CM21]. Commentary [CLL23]. Commerce [QXZ^{+}18]. Commodity [GZL^{+}20, SWDX20, WJW^{+}18]. Common [XLLC16]. Communication [HWB^{+}20, ONP^{+}24, YWW^{+}21b, ZYB^{+}16]. Communication-Efficient [ONP^{+}24, YWW^{+}21b]. Communications [ZPT^{+}23]. Communities [ADF^{+}17, HWW^{+}22, LND22, QLW^{+}22]. Community [CCZ^{+}20, CZ17b, GCL^{+}23, HZF^{+}22, HLF^{+}22, LYW^{+}22, WYBH24, YJH21, ZSJ23, ZHKL24]. Commuting [PLZL22]. Comorbidity [GCA^{+}20]. Comparison [ADF^{+}17, LLS^{+}20]. Competitive [GZW23]. Completion [FZC20, LQZ23, ZWZ^{+}23]. Complex [SD18, SBR22, ZSK19, ZS22]. Complexity [SPH23, SZG^{+}23]. Complexity-Aware [SPH23]. Component [HLL^{+}23a]. Composable [KFY^{+}23, WZLS18a]. Composing [LZD^{+}23]. Comprehensive [APK22, WLF^{+}22]. Compress [LWL^{+}23]. Compression [DH18, LWL^{+}23, LWQ^{+}23]. Compressive [CLG^{+}20, JZS^{+}20, LZL^{+}22]. Compressors [LZD^{+}23]. Computation [HB16, WSW^{+}18, ZYCL20, ZZX^{+}21]. Computing [DWXZ20, GQZ21, HMW^{+}22, MLQ^{+}19, QSC19, QWL^{+}23b, SSSL22, SBL^{+}22, TJZ^{+}22, WYC^{+}20, WTM18, XCV22, YZDZ19, ZZR20, ZYCL22, ZSL^{+}22, ZYB^{+}16, ZWC^{+}16, ZMS17, ZBB^{+}20]. Concept [SLL^{+}22, TNX23]. Concepts [SZA^{+}22]. Concurrency [LS18b]. Conditional [HX22, ZLZL22]. Conductance [WYBH24]. Confidentiality [PWS^{+}19]. Configurable [TF22]. Congestion [NLC17]. Conjugate [Yan23]. Conjunctive [XWZ^{+}23]. Connected [HHZ^{+}24]. Connections [CSW18]. Connectivity [PIMP17]. Conquering [ZWH^{+}22]. Consider [LS18b]. Consistency [CDZ^{+}24, LRL^{+}17, LLN^{+}24, SGMB17, XYX^{+}23]. Consistent [KdRFA21, LZW^{+}23b, Sun15]. Constrained [JKF19, KZSW23, LWZ^{+}22, LJC^{+}22, MKJ^{+}24]. Constraint [GMZ^{+}22, SPH23]. Constraint-Driven [SPH23]. Consumption [FLYL21]. Contagion [LRC20]. Content [AR18, CLHH21, DH18, HLLZ22b, WTRK16]. Content-Aware [DH18]. Context [CJY20, KAB^{+}21, MW22, SYv^{+}19, WYLG21, YLL22, YLL23]. Context-Aware [CJY20]. Context-Awareness [MW22]. Context-Driven [SYv^{+}19]. Contexts [ZZLW21]. Continental [ZMGIVO20]. Continued [LTTC16]. Contrast [LLCS17]. Contrasting [XWG^{+}23b]. Contrastive [CTC^{+}24, YM24]. Control [BSL^{+}22, HLC^{+}18, Sak20, SYv^{+}19, XCV22]. Controversial [LCWL21]. Conversion [YKL^{+}23]. Convex [LLSL19]. Convolution [DBMS24, YXZL23, ZWZ^{+}23]. Convolutional [BSY^{+}24, CSZ^{+}21, CLHN22, STZL22, SLL^{+}22, STM^{+}20, XDZ20, YLLC23, ZAL^{+}23, ZZ23]. Coordinate [CYXS23]. Coordinates [CSW23]. Copernicus [SCPS20]. Coping [Mat16]. Copy [LCY^{+}22, Wan16]. Coresets [ZOLP21]. Cornac [PA20]. Corners [JZY22]. Correction [LY17]. Correlated [CMZL21, WWK^{+}21]. Correlating [ZMGIVO20]. Correlation [DHQ^{+}23]. Correlations [HVVP21, LLW^{+}24]. Cosine [HLL^{+}23a]. Cost [GQP^{+}23, HCZ^{+}24, KAB^{+}21, LXLF23, NHL18, SGM21, WW21, YAG^{+}22]. Cost-Driven [GQP^{+}23]. Cost-Effective [NHL18]. Cost-Effectively [KAB^{+}21].

Cost-Efficient [LXLF23, WW21].

Cost-Sensitive [HCZ^{+}24]. Counterfactual [ZPMT^{+}21]. Country [CZQ^{+}21]. Coupling [LLT^{+}23]. Courier [WHGW23]. Course [MZF24]. Covariance [CQH^{+}18, WCYL21]. Coverage [WW21]. COVID [CZQ^{+}21, DZ21, FRS^{+}22, LMPS21, LZS^{+}21, LCWL21, MSRS21, PTL^{+}21, SPYS23, WVD^{+}21, YWG^{+}21]. COVID-19 [CZQ^{+}21, DZ21, FRS^{+}22, LMPS21, LZS^{+}21, LCWL21, MSRS21, PTL^{+}21, SPYS23, WVD^{+}21, YWG^{+}21]. COVID-19-CT-CXR [PTL^{+}21]. CP [CZZ^{+}24]. CPU [MCA^{+}20]. Crime [LCC^{+}23, WYK^{+}19]. Criterion [Sun15]. Critical [BVAW^{+}16, LLCS17]. Crop [SLV^{+}20]. Cross [BSY^{+}24, HLL^{+}23b, LYD^{+}22, LLN^{+}24, SPYS23, SZSY24, WHGW23, YLC^{+}22, Zhe15]. Cross-Domain [HLL^{+}23b, SZSY24, YLC^{+}22, Zhe15]. Cross-Fertilization [SPYS23]. Cross-Language [SZSY24]. Cross-Modal [BSY^{+}24, LYD^{+}22, LLN^{+}24]. Cross-Region [WHGW23]. CrowdExpress [CYW^{+}22]. Crowdfunding [WLZ^{+}20]. Crowdsensing [CMZL21]. Crowdsourced [CYW^{+}22, LLD^{+}20]. Crowdsourcing [WW21]. CT [PTL^{+}21, PTL^{+}21, YWG^{+}21]. Cultural [LZS^{+}21]. Current [KAG^{+}22]. Curve [BPP21]. CXR [PTL^{+}21]. Cyber [BWW^{+}20, BHHDA19, CCD19b, CCD19c, FYD^{+}21, HLH^{+}20, HWC20, LYG^{+}21, Sak20, WYC^{+}20]. Cyber-Physical [BWW^{+}20, HLH^{+}20, HWC20, Sak20]. Cyber-Physical-Social [FYD^{+}21, LYG^{+}21, WYC^{+}20]. Cyberphysical [ICZ20]. Cybersecurity [TAL19]. Cyclic [LLT^{+}23].

D [CZLL19]. DAAC [BSL^{+}22]. Daily [MLK21]. Danmu [LZW^{+}22]. Data [Agg15a, Agg15b, Agg16, ADF^{+}17, AI23, APK22, ADHS21, AMAT23, BVAW^{+}16, BHO21, BTM16, BHHDA19, BED^{+}24, BLYM19, CZG^{+}19, Can16, CZZ^{+}24, ÇLM17, CW22, CKH16, CXT^{+}18, CW18a, CW18b, CWPW20, CGH^{+}22, CLW^{+}22, CWC^{+}23, CWS^{+}21, CL23, CCD19b, CQH^{+}18, CLF^{+}18, CSL^{+}18, CDLW19, CSW23, DQWQ18, DFF^{+}23, DLD^{+}20, DLZ^{+}19, DHG21, DHQ^{+}23, DFG^{+}19, DLL^{+}16, DH18, ECG^{+}19, FY22, FZ18, FRS^{+}22, FYD^{+}19, FYD^{+}21, FWWC23, FYZ^{+}22, GQZ21, Gil16, GM20, GDY24, Gou19, GZL^{+}20, GWLL24, GSS23, HD19, HHZ^{+}24, HXW^{+}23, HCZ19, HPWR20, HVVP21, HB16, HPESZ17, HLC^{+}18, HLH^{+}20, HLW^{+}23, JGS^{+}19, JFG17, JHT18, JFZ^{+}21, JWS^{+}22, KAG^{+}22, KdRFA21, KKKK16, KAB^{+}21, KZL^{+}22, KJC^{+}18, LMPS21, LTX16, LLLZ21, LCWS22, LCY^{+}22, LZC^{+}23, LWS^{+}24, LCC^{+}23, LZH^{+}24, LLSL19, LTTC16, LTTC17, LS18b, LW18, LLY^{+}18, LYG^{+}21, LDY^{+}22, LLO^{+}22, LWL^{+}23, LWQ^{+}23, LLF^{+}19, MLQ^{+}19, MCM21, MCH^{+}22, MCL22]. Data [NLC17, Ni15, NHL18, PWW^{+}23, PWH16, PRR22, PIMP17, PWS^{+}19, QXZ^{+}18, QHC^{+}22, QFW^{+}23, QLZ24, RJS22, RBM^{+}22, SAJP20, SPYS23, SJS^{+}21, SD18, SF23, SAM^{+}20, SD22, SZC^{+}21, SSSB16, SCW18, SLL^{+}21, SVYY16, SZY^{+}18, SRM17, SPH23, SZA^{+}22, SXS^{+}23, SZG^{+}23, STM^{+}20, SX21, SWTX18, SZZ^{+}23, SCPS20, TvMvdH23, TAL19, TF22, TRB^{+}21, TAL^{+}17, TCC21, VP21, WTRK16, WQSA17, WYLD18, WZY^{+}18, WYK^{+}19, WXZ^{+}19, WBD^{+}20, WYLG21, WLF^{+}22, WXW^{+}22, WCS^{+}23, WHX^{+}23, WWC^{+}23, WFZ^{+}23, WWYH24, WCYL21, WTM18, WS17, WOD^{+}18, WZLS18a, WWK^{+}21, XWBL17, XCV22, XSYW22, XWZ^{+}23, XZJT24, YDY^{+}16, YZDZ19, YWW^{+}21b, YWRY21, YLC^{+}22, Yao15, Yu15, YWL^{+}22, YLL22, YLLC23, YLL23, YZW^{+}23b, YWM^{+}21, ZZR20, ZHL^{+}17, ZYYK18, ZSK19, ZWV^{+}19, ZFLC19, ZSS^{+}21, ZGS^{+}22, ZQD^{+}22, ZYCL22, ZRL22, ZWZ^{+}23, ZBL23, ZQZ^{+}17, ZCH^{+}18, Zhe15, ZOLP21, ZYX^{+}23, ZSL17, ZLC^{+}17, ZZZ^{+}18, ZBB^{+}20, ADOAH19, BZY20, CCD^{+}19a, CCD19c]. Data [CLG^{+}20, DLL^{+}20, DWXZ20, HFX^{+}20, HWB^{+}20, HWTM20, HWC20, LLZ^{+}19, LA20, MHR^{+}23, PA20, RDM^{+}20, XZGW21, XHZW21, YLW^{+}20, ZSK19, ZYCL20, ZBN^{+}20, Ano15a, Ano15b, Ano17a, Ano18a, Ano19a, Ano20, Ano21, Ano22, Cra15, Yan15b]. Data-as-a-Service [WYLD18]. Data-Aware [HD19]. Data-Driven [CZZ^{+}24, GM20, JFZ^{+}21, KJC^{+}18, LLO^{+}22, PRR22, QHC^{+}22, ZSK19, LWQ^{+}23]. Data-Pattern [ECG^{+}19]. Data-Sharing [ZZR20]. Database [HXW^{+}23, WHX^{+}23]. Databases [KKF16, ZHW21]. Datacenter [YAG^{+}22]. Dataset [FMD18, LZW^{+}22, WZLS18b]. Datasets [ADHS21, DFF^{+}23, Far21, KHdMR20, Kot15, WZY^{+}18, XPC^{+}23]. Datastores [SGMB17]. DAWN [GZS23]. De-Identification [DWSJ19]. Deadline [SGM21]. Deadline-Aware [SGM21]. Debugging [LY17]. Decentralized [CYXS23, HRFS^{+}24]. Decision [GSS23, MSRS21, SZZW24, ZZLW21, ZRY^{+}20]. Decision-Making [MSRS21, SZZW24]. Decisions [JGS^{+}19]. Decline [XLL^{+}18]. Decoder [WJW24, ZGS^{+}22]. Decoding [ZZX^{+}21]. Decomposition [CZZ^{+}24, HHZ^{+}24, LCC^{+}23, LW18, LYG^{+}21]. Deduplication [CDLW19, HKKC22, YDY^{+}16, YZDZ19, YLC^{+}22]. Deep [ADHS21, CSZ^{+}21, CKL^{+}23, HLLZ22a, HPESZ17, HXYZ20, JRP^{+}22, JKF19, LQZ23, LTG^{+}24, LWZL24, MXL^{+}20, SLL^{+}22, SSY^{+}23, TNX23, TXH^{+}20, WLD^{+}23, WWCZ22, XWG23a, YDL^{+}22, ZLZ^{+}20, ZYCL20, ZSL^{+}22, Zhu15, ZZZ^{+}22, HCS^{+}22]. Deep-Learning [HLLZ22a]. Deepfake [YSM^{+}23]. Definite [WCYL21]. Definition [Far23]. Deformation [ZBB^{+}20]. Degree [ZHL^{+}24]. Delaunay [CGH^{+}22]. Deliveries [CYW^{+}22]. Delivery [HLLZ22b, LA20, WHGW23]. Demand [HSEY20, JFZ^{+}21, MXL^{+}20, MCA^{+}20, QSC19, WHGW23, ZHL^{+}17, ZZZ^{+}22]. Demographic [FWWC23]. Demographic-Augmented [FWWC23]. Dense [CL23, SDR^{+}21]. Densest [WS17]. Density [FY22, QLW^{+}22, ZDZC23, ZOLP21, dSSN^{+}21]. Density-Based [QLW^{+}22, dSSN^{+}21]. Density-Increasing [ZDZC23]. Dependencies [TAL19]. Deployment [CZLL19, LB19, ZLZ^{+}22]. Depression [YWM^{+}21]. Descent [CYXS23, QLZ24, SHL^{+}22]. Description [XML^{+}19, ZLY^{+}22]. Descriptors [WWCK22, YLLC18]. Design [JGS^{+}19]. Designing [LLZ^{+}19]. Desktop [TF22]. Destination [ZZZ^{+}22]. Detecting [DLD^{+}17, HHX^{+}19, XLL^{+}18]. Detection [AK19, ADOAH19, ASYX21, BLL^{+}20, BWW^{+}20, CLHH21, CCZ^{+}20, CZ17b, DZS20, DBMS24, GCL^{+}23, HZF^{+}22, HLF^{+}22, HCS^{+}22, HCZ^{+}24, HMW^{+}22, JPR^{+}20, JDC^{+}22, KZSW23, LCWS22, LBSL22, LYLW23, LQHC21, LLCS17, LYW^{+}22, LZZF23, LRC20, MSC19, NCS17, NLY22, SD22, SZSY24, TLS24, Wan16, WYBH24, WJS^{+}16, YDM^{+}23, YJH21, YWL^{+}22, ZLY^{+}22, ZDZC23, ZSJ23, ZHKL24]. Determination [WHT18]. Deterministic [GNPT21]. Development [HZZ^{+}21]. Device [Mat16]. Devices [GZL^{+}20, MKJ^{+}24]. Diagnosis [YZW^{+}23a, ZYW^{+}23]. Dialogue [WJW24]. Different [LLS^{+}20, XLL^{+}18]. Differential [BWW^{+}20, CMZL21, DWXZ20, TCC21, ZYX^{+}23]. Differentially [CWC^{+}23, WCS^{+}23, XJQ^{+}21, YWRY21, ZCSZ23]. Differentially-Private [XJQ^{+}21]. Differentiated [PWH16]. Difficulties [ZWH^{+}22]. Diffusion [GZW23]. DIGDUG [SDR^{+}21]. Digital [BSL^{+}22, BED^{+}24, JFZ^{+}21]. Dimension [WWL^{+}23, ZZL^{+}22]. Dimensional [CQH^{+}18, HHZ^{+}24, IFG^{+}18, KKKK16, LQZ23, QLZ24, WCYL21, XLP^{+}18, YHLS22]. Dimensionality [SZJ^{+}22]. DiNoDB [TAL^{+}17]. DInSAR [ZBB^{+}20]. DiP [SRM17]. DiP-SVM [SRM17]. Directed [ADF^{+}17, CHL23, HLF^{+}22]. Directional [CZLL19, YYD^{+}23]. Disability [TLL^{+}23]. Disaster [PWW^{+}23]. DISboards [LCS18]. Disclosure [SYv^{+}19, TAL19, TCC21]. Discovering [LYL^{+}21, NLC17, WS17, DSS16]. Discovery [Can16, HWC20, IZ19, LTTC16, LTTC17, LLZ^{+}23, TEOS17, VP21, ZWV^{+}19, ZCSZ23, ZLT20]. Discrete [Li15, NLG^{+}23, WL22b, XLC22]. Discrimination [LRL^{+}17]. Discriminator [ZYW^{+}23]. Discussion [LCS18]. Disease [WWYH24, YZW^{+}23a, ZYW^{+}23]. Disentangle [SZSY24]. Disjoint [GNPT21]. Disjoint-Parallel [GNPT21]. Displacement [WHGW23]. Distance [ZWS21]. Distance-Based [ZWS21]. Distillation [TLS24]. Distinctive [HBLK17]. Distributed [BLYM19, CWGA17, CWPW20, DFF^{+}23, DDGU22, DFG^{+}19, DLL^{+}16, LTX16, LHY^{+}22, LLH^{+}23, MLQ^{+}19, MBS^{+}19, Mat16, MTV21, Ni15, SF23, SHL^{+}22, SZG^{+}23, SWTX18, SWH^{+}22, WYC^{+}20, XJQ^{+}21, YWW^{+}21b, Yu15, ZSL^{+}22, ZYB^{+}16, ZCH^{+}18]. Distribution [LTG^{+}24, LLW^{+}24, SRM17, WZY^{+}18]. Diversified [LYLJ22, XLL^{+}22]. Diversity [XLL^{+}22]. DL [HKKC22]. DMP [HLLZ22b]. DNN [SWH^{+}22]. Document [HLW^{+}23]. Document-Level [HLW^{+}23]. Domain [GZS23, HLLZ22a, HLL^{+}23b, KS22, MLK21, RJS22, SZSY24, SZXL22, YLC^{+}22, Zhe15]. Domain-Shift [RJS22]. Domain-Specific [HLLZ22a]. Don’t [SZSY24]. Double [WL22b, YCYL24]. Double-Bit [WL22b].

Drift [LLH^{+}23, TNX23]. Drift-Sensitive [LLH^{+}23]. Driven [CZZ^{+}24, FWS^{+}23, GQP^{+}23, GM20, GZL^{+}20, ICZ20, JFZ^{+}21, KJC^{+}18, LLO^{+}22, PWW^{+}23, PRR22, QHC^{+}22, SCW18, SPH23, SYv^{+}19, ZSK19, LWQ^{+}23]. Driver [JPR^{+}20, ZRY^{+}20]. Drivers [ZLZL22]. Driving [CLW^{+}22]. Drought [TvMvdH23]. Drug [SZY^{+}22]. Drug-Target [SZY^{+}22]. DSTAGCN [ZZ23]. Dual [SF23, XYX^{+}23]. Duo [KZSW23]. Duplication [CWPW20]. During [LZS^{+}21, MSRS21]. DVFS [LLY^{+}18]. DVFS-Enabled [LLY^{+}18]. Dynamic [CWL^{+}21, CHL23, FY22, HLLZ22b, LMPS21, LBZ19, LLLZ21, SZC^{+}21, SPN^{+}22, WCW^{+}22, WLXF23, WZL^{+}24a, XPC^{+}23, ZGH^{+}22, ZZ23]. Dynamical [PRR22]. Dynamics [DDCHV21, GM20, Sak20, SCA^{+}17, SHH^{+}21, ZFLC19].

E-Commerce [QXZ^{+}18]. E-Health [BSL^{+}22]. E-Optimal [JZS^{+}20]. Early [ASYX21, TvMvdH23]. Early-Warning [TvMvdH23]. eBoF [DHQ^{+}23]. Economic [MSRS21, ZCH^{+}18]. Ecosystem [LZH^{+}24]. Edge [DWXZ20, EL22, MW22, TJZ^{+}22]. Editor [Tan20, Yan20]. Editor-in-Chief [Tan20, Yan20]. Editorial [Agg15a, Agg15b, Agg16, Can16, CW18a, CW18b, DZ21, JWS^{+}22, LTTC16, LTTC17, SVYY16, SSSL22, WQSA17, Yan17, Yan18, Yao15, ZMS17]. Editors [HCZ19]. EEG [HPESZ17, LRL^{+}17]. Effective [CLHH21, DBAM17, GKR17, KJC^{+}18, NHL18, SCA^{+}17, SZY^{+}18, ZZL^{+}22]. Effectively [KAB^{+}21]. Efficiency [WWC^{+}23, ZPT^{+}23, ZRY^{+}20]. Efficiency-Oriented [WWC^{+}23]. Efficient [APK22, CZZ^{+}24, CLHH21, CHL23, CMJ^{+}24, DDGU22, DWSJ19, FWS^{+}23, GLZ^{+}24, HS20, HXW^{+}23, HFX^{+}20, Hua15, JWLY23, KKF16, LOLL17, LRB^{+}20, LA20, LXLF23, LLSL19, LLZ^{+}23, MW22, MCO^{+}22, ONP^{+}24, SLC^{+}18, SLZ^{+}22, SWDX20, TRB^{+}21, WW21, WS17, WCJI22, XWZ^{+}23, YWW^{+}21b, YLC^{+}22, YWY^{+}22, YWL^{+}22, YRL^{+}24, ZRL22, ZRLZ23, ZQZ^{+}17, ZCH^{+}18, ZLG^{+}21, CM21]. Eigenvalues [HS20]. Eigenvectors [ZHKL24]. Elastic [CL23]. Electricity [HSEY20, WXZ^{+}19]. Email [LCWS22]. Embedded [CGZ23]. Embedding [CXW^{+}23, HLL^{+}23b, JGP^{+}22, LZZ^{+}23, LZZF23, SZXL22, SSY^{+}23, WGYW23, WBS^{+}23, YKL^{+}23]. Embeddings [FWS^{+}23, QWL^{+}23a]. Embracing [Liu15]. Emergency [SZA^{+}22, XLL^{+}18]. Emerging [HCZ19, YLB^{+}19]. Emotion [LRC20, SZSY24]. Emotions [SZSY24]. Empirical [ADF^{+}17, DBAM17, DEG^{+}22, ZYZ^{+}23]. Employment [TLL^{+}23]. Empowered [WJW^{+}18]. Emulate [BED^{+}24]. Enabled [ECG^{+}19, HWB^{+}20, LLY^{+}18]. Enabling [RBM^{+}22, WGYW23, ZSK19]. Encoder [HZF^{+}22, JKF19, WJW24, YLL22, YLL23, HCS^{+}22]. Encoder-Decoder [WJW24]. Encrypted [CWS^{+}21, CDLW19, XWZ^{+}23, YDY^{+}16, YLZ^{+}22]. Encryption [FLYL21, GQZ21, PWS^{+}19, WYLG21]. End [KAG^{+}22, KAB^{+}21, SBL^{+}22]. End-to-End [KAB^{+}21, SBL^{+}22]. End-User [KAG^{+}22]. Energy [FLYL21, SZG^{+}23]. Engine [BZY20, JRP^{+}22, KS22, TAL^{+}17]. Engineering [DSS16]. Engines [RBM^{+}22]. Enhance [BHO21]. Enhanced [BHX^{+}23, FZ18, SZY^{+}22, TLS24, ZHW21]. Enhancement [LTG^{+}24, YWG^{+}21, ZFQ^{+}23]. Enhancing [QYC^{+}23]. Enriching [MCA^{+}20]. Ensemble [DHQ^{+}23, MHR^{+}23, SJS^{+}21, WWYH24]. Ensure [PWS^{+}19]. Ensuring [WSW^{+}18]. Entanglements [LZS^{+}21]. Enterprise [MCL22]. Entities [ESC20]. Entity [BHO21, KS22, PP22, QWL^{+}23a, ZGH^{+}22, ZSC^{+}22]. Entity-Based [ZGH^{+}22].

Entity-Centric [KS22]. Entity-Resolution [BHO21]. Entropy [FWS^{+}23, HZY^{+}21, XPC^{+}23]. Entropy-Driven [FWS^{+}23]. Environment [FYD^{+}19, KdRFA21, XLC20, ZBB^{+}20]. Environments [HFX^{+}20, SD18, XLP^{+}18]. Epidemic [SPYS23]. Epidemiological [LMPS21]. Equations [SLC^{+}18, CM21]. Erasure [KGJØ18]. Error [LZD^{+}23]. Error-Bounded [LZD^{+}23]. Escaping [XHZW21]. Estimates [ZOLP21]. Estimating [HS20]. Estimation [CWC^{+}23, CLL23, EL22, HX22, JKF19, MSC19, STM^{+}20, TCC21, WCYL21, ZLZ^{+}22, ZSL17]. Euclid [STM^{+}20]. Euler [WZLS18b]. Evaluating [TCC21]. Evaluation [CQH^{+}18, GZW23, SD18, SWC^{+}22, Wan16]. Evaluation-Based [SWC^{+}22]. Event [ASYX21, BWW^{+}20, LGW^{+}23, LRC20, MW22, NCS17, SAJP20, SHH^{+}21, SD22, SBR22, ZSK19]. Events [NYL^{+}22, XLL^{+}18, XML^{+}19]. Evidence [CCD^{+}19a]. Evidential [GDY24]. Evolution [ANP^{+}23, ZZY^{+}23]. Evolutionary [HD19]. Evolving [CCZ^{+}20, CMJ^{+}24]. EvoSets [SZJ^{+}22]. Examples [QYC^{+}23]. Exchange [SXS^{+}23]. Execution [ZSK19]. Expansion [ZSC^{+}22]. Experience [TvMvdH23]. Experts [YMJ^{+}23]. Explainable [KZSW23]. Exploiting [LMPS21, LQHC21, LLN^{+}24, SWF16, TXH^{+}20, WS17, WCJI22, XLLC16]. Exploration [Kit16, MBS^{+}19, Tao15]. Exploratory [RBM^{+}22]. Exploring [HWTM20, LWL^{+}23, WWL^{+}23, ZZLW21, ZQK17, ZMGIVO20]. Explosion [Liu15]. Exposing [YSM^{+}23]. Extended [ZSL17]. Extending [LS18b]. Extracting [Gil16, LLD^{+}20]. Extraction [CLHN22, KFY^{+}23, LGW^{+}23].

Facial [CLGX23]. Factor [CLZ22, HLF^{+}22, LZL^{+}21, SHL^{+}22, YHLS22, ZJSL22].

Factorization [CJY20, FZC20, HZF^{+}22, WZL^{+}24b, ZYX^{+}23]. Factorizations [LZSL21]. Factors [MXL^{+}20]. Fair [JHT18]. Fake [DBMS24]. Fast [ADHS21, AR18, CJY20, HCS^{+}22, LZL^{+}22, MBG22, RBM^{+}22, WCYL21, YWW^{+}21b]. Faster [FWHM22]. Fatigue [JPR^{+}20]. FATS [LTG^{+}24]. Fault [LQHC21]. Fauxtography [KZSW23]. Feature [CDZ^{+}24, CGZ23, CSZ^{+}21, CLGX23, HHZ^{+}24, JDC^{+}22, LTG^{+}24, LLZ^{+}23, LLP^{+}23, MLK21, YYD^{+}23, ZYCL20, ZFQ^{+}23, ZCH^{+}18, ZWS21, ZZL^{+}24]. Feature-Based [HHZ^{+}24]. Feature-Weighted [CDZ^{+}24]. Features [DHQ^{+}23, JPR^{+}20, JZY22, TXH^{+}20, TLS24, WJS^{+}16, ZHL^{+}24]. FedCKE [HLL^{+}23b]. Federated [DBMS24, GQP^{+}23, HRFS^{+}24, HLL^{+}23b, HGY^{+}23, JWLY23, LPT^{+}23, LLL^{+}22, LLC^{+}24, ONP^{+}24, PB23, RDM^{+}20, XZJT24, YZW^{+}23a, ZPT^{+}23]. Feedback [CJY20]. Fertilization [SPYS23]. Few [SZZW24]. Few-Shot [SZZW24]. FG [DLL^{+}24]. FG-SBIR [DLL^{+}24]. FHE [LMW21]. Fi [KZL^{+}22, JPR^{+}20]. FIG [CDZ^{+}24]. Filaments [WJS^{+}16]. File [SWTX18]. Filtering [HH18, LLP^{+}23, WTX^{+}22]. Filtration [ZLG^{+}21]. Finance [ZYYK18, ZSK19]. Find [ADF^{+}17]. Finding [FY22]. Fine [CZ17a, CLW^{+}22, JPR^{+}20, LOLL17, LSH^{+}23, QWL^{+}23a, XWG^{+}23b, XZJT24, YDL^{+}22]. Fine-Grained [CZ17a, CLW^{+}22, JPR^{+}20, LOLL17, LSH^{+}23, QWL^{+}23a, XWG^{+}23b, YDL^{+}22]. Fine-Tuned [XZJT24]. Fingerprint [ZHW21]. Fingerprints [YSM^{+}23]. Fixation [YYD^{+}23]. Fixation-Based [YYD^{+}23]. FLAG [FWHM22]. Flash [YAG^{+}22]. Flexible [ZQZ^{+}17]. Flink [LXLF23]. Flow [RDD^{+}21, SCA^{+}17, XWG^{+}23b, ZZY^{+}23]. Flows [Kit16]. Fly [DLL^{+}24]. fMRI [MLQ^{+}19, PIMP17]. Fog [SBL^{+}22].

Forecast [HLLZ22a]. Forecasting [NYL^{+}22, WXZ^{+}19, WZL^{+}24a, YKL^{+}23, ZJG^{+}20, ZYZ^{+}23, ZZ23]. Forecasts [GSS23]. Forensics [ZHW21]. Forest [LLL^{+}22]. Forests [YWW^{+}21a]. Forgery [ZFQ^{+}23]. Formation [Ng16]. Formulation [HSEY20]. Forum [LCS18]. Foundation [SBL^{+}22]. Four [JZY22]. FPGAs [ZZL^{+}22]. Frame [MKJ^{+}24]. Framework [CYW^{+}22, CLGX23, CZ17b, DHG21, DLL^{+}16, GM20, KFY^{+}23, LBSL22, LZD^{+}23, SZY^{+}22, WTRK16, WYLD18, WZLS18a, WZL^{+}24b, XWI^{+}21]. Frameworks [NHL18]. Fraud [HCZ^{+}24, JDC^{+}22, LYLW23, WZH^{+}23]. FRC [MKJ^{+}24]. FRC-GIF [MKJ^{+}24]. Free [WGG^{+}23, ZLT20]. Freely [PTL^{+}21]. Frequency [CWC^{+}23]. Frequent [JY22]. Fully [HHZ^{+}24, XDZ20]. Fully-Connected [HHZ^{+}24]. Function [ZRLZ23]. Functional [PIMP17]. Fusing [YWM^{+}21, YXZL23]. Fusion [AK19, BHO21, CLGX23, MZF24, SZZW24, TLS24, WJS^{+}16, XPC^{+}23, YWG^{+}21, Zhe15]. Future [LSH^{+}23]. Fuzzy [BTM16, JRP^{+}22, KdRFA21].

Gain [YJH21]. Game [CZ17b, RJS22, SX21, WWK^{+}21, XJQ^{+}21]. Game-Theoretic [CZ17b]. Gamma [JRP^{+}22]. GAN [WCS^{+}23]. GAN-Based [WCS^{+}23]. GAT [HCZ^{+}24]. GAT-COBO [HCZ^{+}24]. Gate [ZGC^{+}24]. Gaussian [DEG^{+}22, LOLL17, LZMZ20]. GCN [YLLC23, YXZL23]. GCN-ST-MDIR [YLLC23]. GDPR [AMAT23]. General [LLY^{+}17]. Generalization [GZS23]. Generalized [WLD^{+}23]. Generation [MKJ^{+}24, WJW24, ZBB^{+}20]. Generative [FWWC23, HWW^{+}22, KFY^{+}23, WHT18, YZW^{+}23b, ZLZL22, ZLZ^{+}22]. Genetic [ZSHS17]. Genomic [ADHS21, ICZ20]. Geo [CWPW20, JKF19, WLXF23, ZCSW23]. Geo-Constrained [JKF19].

Geo-Distributed [CWPW20]. Geo-Human [WLXF23]. Geo-Indistinguishability [ZCSW23]. Geographical [LYL^{+}21, ZQK17]. Geographically [DFG^{+}19]. Geolocation [LLT^{+}23]. Geometric [MSC19]. Geospatial [KdRFA21, LRC20]. Geotagged [ÇLM17]. Geotagging [KKF16]. Gestational [LLS^{+}20]. GGNN [GRWL23]. GGraph [SLZ^{+}22]. GIF [MKJ^{+}24]. Global [CLHN22, DLL^{+}24, LYLW23, MCO^{+}22, SCPS20]. Global-Local [CLHN22]. GNN [JDC^{+}22, LYLW23]. GNN-Based [JDC^{+}22]. GNNs [WGYW23]. Good [KS22, TvMvdH23]. Goodman [JRP^{+}22]. GPS [ASYX21, CLW^{+}22, JKF19, LLY^{+}17, ZSS^{+}21]. GPU [CXT^{+}18, GRWL23, Mat16, ZSJ23]. GPU-Accelerated [CXT^{+}18, Mat16]. GPUs [JDJ21, SAJP20]. Gradient [QLZ24, SHL^{+}22, Yan23, ZWD^{+}23]. GradientFlow [SWH^{+}22]. Grained [CZ17a, CLW^{+}22, GB23, JPR^{+}20, LOLL17, LSH^{+}23, QWL^{+}23a, XWG^{+}23b, YDL^{+}22]. Granger [ZSL17]. Granular [SYv^{+}19]. Granularity [SZZW24]. Granules [CDZ^{+}24]. Graph [BSY^{+}24, CXW^{+}23, CTC^{+}24, CLHN22, FWS^{+}23, FWHM22, GRWL23, GCL^{+}23, HZF^{+}22, HLF^{+}22, HSX^{+}22, HCZ^{+}24, HLL^{+}23b, JY22, JDC^{+}22, LMPS21, LZL^{+}22, LHY^{+}22, LYLW23, LYC^{+}24, LRL^{+}17, LZZF23, LZW^{+}23b, PA20, SWF16, STZL22, SDR^{+}21, SLZ^{+}22, SWDX20, WWL^{+}23, WZH^{+}23, WBS^{+}23, XWG23a, XLQ^{+}22, XZJT24, YM24, YCYL24, YMJ^{+}23, YLLC23, YXZL23, ZZL^{+}22, ZWS^{+}22, ZWZ^{+}23, ZYB^{+}16, ZWLN22, ZWD^{+}23, ZLG^{+}21, ZSC^{+}22, ZZ23, ZSHS17, ZZZ^{+}22]. Graph-Based [GRWL23, LYC^{+}24]. Graph-Embedding [CXW^{+}23]. Graph-Parallel [ZYB^{+}16]. Graph-Regularized [ZSHS17]. Graph-Transaction [JY22]. Graphical [HLF^{+}22]. GraphMP [SWDX20]. Graphs [ADF^{+}17, CMJ^{+}24, GLZ^{+}24, HS20, HYLZ22, HH18, KS22, LHWL24, LWZ^{+}22, QLZ^{+}23, WGYW23, YRL^{+}24]. Grasping [SSAI21]. Green [NHL18, QSC19]. Green-Aware [NHL18]. Grid [CQH^{+}18, HSEY20, SAJP20, WXZ^{+}19, WOD^{+}18, SAM^{+}20]. Grid-Based [SAJP20]. Grids [LLF^{+}19]. Group [FYZ^{+}22, STM22, ZCSW23]. Group-Based [ZCSW23]. Grouping [LRC20]. Guarantee [SYS^{+}18]. Guarantees [Far23]. Guest [Agg15a, Agg15b, Agg16, Can16, CW18a, CW18b, DZ21, JWS^{+}22, LTTC16, LTTC17, SVYY16, SSSL22, WQSA17, Yao15, ZMS17, HCZ19]. Guided [HSX^{+}22, XYX^{+}23].

Habits [DBAM17]. Hadamard [TZCC22]. Hadoop [DH18]. Handle [WGYW23]. Handling [BTM16, HD19]. Harvesting [LWQ^{+}23]. Hashing [BSY^{+}24, Li15, LYD^{+}22, LLN^{+}24, NLG^{+}23, Sun15, TNX23, WL22b, XLC22, ZHW21]. HashTag [KGJØ18]. Having [Sak20]. HDFS [HD19]. HDM [WZLS18a]. Health [BSL^{+}22, FWWC23, GCA^{+}20]. Healthcare [DEG^{+}22, FRS^{+}22]. Heat [MXL^{+}20]. Heterogeneous [CZLL19, Gou19, JDC^{+}22, LZZ^{+}23, LZW^{+}23a, MCH^{+}22, MLK21, QFW^{+}23, WW21, WGYW23, WBS^{+}23, YZDZ19, ZSL^{+}22, ZZL^{+}24, ZSL17]. Hidden [HHX^{+}19]. Hierarchical [CLZ22, DLL^{+}16, HKKC22, KKKK16, LWZL24, LND22, SLL^{+}21, SCG^{+}23, WJW24, WWSB16, XDH23, YLL23, dSSN^{+}21]. Hierarchy [LWZ^{+}22]. Hierarchy-Constrained [LWZ^{+}22]. High [CDZ^{+}24, CXT^{+}18, CQH^{+}18, DLD^{+}17, FYD^{+}19, FYD^{+}21, HXYZ20, IFG^{+}18, KKKK16, LS18a, LWL^{+}23, QLZ24, WCYL21, XLP^{+}18, YHLS22, ZYCL22, ZDD^{+}23, ZHKL24]. High-Dimensional [CQH^{+}18, IFG^{+}18, QLZ24, WCYL21, XLP^{+}18, YHLS22]. High-Order [FYD^{+}19, ZYCL22, ZHKL24]. High-Order-Lanczos [FYD^{+}21]. High-Quality [LS18a]. High-Ratio [LWL^{+}23]. High-Resolution [HXYZ20]. High-Throughput [CXT^{+}18]. High-Utility [ZDD^{+}23]. Highly [CHL23, DBAM17]. Highly-Efficient [CHL23]. Hilbert [BPP21]. Hinge [PP22]. Historical [XLLC16]. Hoc [SWC^{+}22, TAL^{+}17]. Homogeneous [JGS^{+}19, WGYW23]. Horizontal [LLW^{+}24]. HOSVD [WYC^{+}20]. Hotspot [TEOS17]. Hub [LLZ^{+}19]. Hub-and-Spoke [LLZ^{+}19]. Huge [PA20]. Human [JFG17, KS22, LZC^{+}23, MSRS21, WLXF23, YLLC18]. Hurricane [NYL^{+}22]. Hybrid [CXW^{+}23, HFX^{+}20, LLZ^{+}19, LZLJ22, TF22, WXW^{+}22]. Hyperbolic [ZWS^{+}22]. Hyperedges [LZZ^{+}23]. Hypergraph [LLP^{+}23, WZL^{+}24a]. Hyperparameter [WHT18]. Hyperspectral [XDZ20].

I/O [SWDX20, YAG^{+}22]. I/O-Efficient [SWDX20]. Ideas [DSS16]. Identification [AI23, DWSJ19, EJCR22, HWW^{+}22, YLZ^{+}22, YLLC23]. Identify [HVVP21]. Identifying [CLW^{+}22, HYLZ22, WYXL21, ZWH^{+}22, ZSHS17, ZZZ^{+}18]. Identity [LLLZ21, SZC^{+}21, XCV22]. Identity-Based [LLLZ21, SZC^{+}21, XCV22]. Idiosyncratic [HBLK17]. IEEE [Ano15a, Ano15b, Ano17a, Ano18a, Ano19a, Ano20, Ano21, Ano22, Cra15, Yan15b]. IEMask [BHX^{+}23]. II [CW18b]. Image [CZQ^{+}21, CSZ^{+}21, CXW^{+}23, DHG21, DLL^{+}16, GXQ^{+}20, KKF16, Li15, LQZ23, LLCS17, NLG^{+}23, PTL^{+}21, Tao15, TNX23, TXH^{+}20, WWCK22, XDZ20, XYX^{+}23, YWG^{+}21, ZLZ^{+}20, ZWS21, Zhu15]. Imagery [HXYZ20, MSRS21, SLV^{+}20, ZBN^{+}20]. Images [CSW18, XDH23]. Imaging [CLG^{+}20, LRL^{+}17]. Imbalance [MTV21].

Imitation [ZLZL22]. Impact [Cha16, DLD^{+}17, GCA^{+}20]. Impacts [MXL^{+}20]. Implementation [CL23]. Implicit [CJY20]. Importance [ZHL^{+}24]. Improved [FYD^{+}21, QWL^{+}23a, YLL23]. Improves [WVD^{+}21]. Improving [JZY22, LZLJ22, XLL^{+}22, YWRY21]. Imputation [FWWC23, LWS^{+}24, YZW^{+}23b]. In-Class [LRL^{+}17]. In-Memory [AR18, CZG^{+}19, LZLJ22, STM22]. Incentive [XSYW22]. Incoming [AR18, Tan20]. Incomplete [CLZ22, QLZ24]. Incompletely [SWC^{+}22]. Incorporated [LCS18, ZJSL22, CLZ22]. Incorporating [CWC^{+}23, KAB^{+}21, LZW^{+}23a]. Increase [TLL^{+}23]. Increasing [ZDZC23]. Incremental [CYXS23, LYG^{+}21, TNX23, WYC^{+}20, WCW^{+}22, ZYCL20]. Index [Ano17a, Ano18a, Ano19a, Ano20, Ano21, Ano22, YRL^{+}24, ZRLZ23]. Indexing [AR18, Li15, Yan15a, ZQZ^{+}17]. Indistinguishability [ZCSW23]. Induced [ZJZ^{+}19]. Inductively [HYLZ22]. Industrial [TLS24]. Infants [LLS^{+}20]. Inference [HWTM20, MW22, MCH^{+}22, PB23, SCA^{+}17, WYK^{+}19, XWG^{+}23b, YTY^{+}21, ZCSZ23, ZPMT^{+}21]. Inferred [JFG17]. Inferring [BS21]. Influence [GZW23]. Influences [LZW^{+}23a]. Influential [LBL^{+}18, LBZ19]. Information [BHX^{+}23, CDZ^{+}24, EJCR22, HVVP21, KFY^{+}23, LWH18, LZZ^{+}23, Liu15, NCS17, Sun15, XDH23, XPC^{+}23, ZZL^{+}24]. Information-Enhanced [BHX^{+}23]. Information-Theoretic [Sun15]. Infrastructure [CW18a, CW18b, DQWQ18, FZ18, MG23, PA20, ZYYK18]. Infrastructures [WTM18]. Innovative [DLD^{+}20]. Inpainting [ZFQ^{+}23]. Input [SBR22]. Instances [GXQ^{+}20, ZHL^{+}24]. Integrating [HLF^{+}22, SJS^{+}21]. Integrative [ICZ20]. Integrity [LCY^{+}22, WSW^{+}18]. Intelligence [CCD19b, CCD19c]. Intelligent [SCW18, ZSL^{+}22]. Intensive [WVD^{+}21]. Inter [SSAI21]. Inter-Attribute [SSAI21]. Interaction [SZY^{+}22]. Interactions [JDC^{+}22, WWL^{+}23, WLXF23]. Interactive [DHQ^{+}23, MCO^{+}22, TAL^{+}17, ZS22]. Interactive-Speed [TAL^{+}17]. Internal [ZLT20]. Internet [IZ19, HPESZ17, MW22, SYv^{+}19]. Interoperable [HXW^{+}23]. Interpolation [ZRLZ23]. Interval [MCM21, SSSB16, XPC^{+}23, MCM21]. Interval-Valued [XPC^{+}23]. Interventions [TLL^{+}23]. Introduction [HCZ19, Yan15b]. Invalid [ZLG^{+}21]. Invariant [MLK21]. Invasive [SHH^{+}21]. Inversion [CW22]. Investigating [SHH^{+}21]. Investigation [SXYL23, TXH^{+}20]. IoT [CYC^{+}23]. IoTs [WWC^{+}23]. IRDA [WCW^{+}22]. ISP [CSL^{+}18]. Issue [CCD19b, CCD19c, DLL^{+}20, HCZ19, JWS^{+}22, LTTC16, SSSL22, YLW^{+}20, ZLC^{+}17]. Item [WWL^{+}23, WGG^{+}23]. Iterative [AK19, PWC^{+}22, SLZ^{+}22].

Java [CWGA17]. Join

[GNPT21, STM22, SDR^{+}21]. Joint [CGZ23, HLF^{+}22, LHWL24, SYS^{+}18]. JouleMR [NHL18]. Journal [Yan17, Yan18, Yan19].

Kernel

[SRM17, WWCK22, ZJZ^{+}19, ZOLP21]. Kernel-Induced [ZJZ^{+}19]. Keys [CWS^{+}21]. Keyword [LPT^{+}23, XWZ^{+}23]. Kira [ZBN^{+}20]. Knowledge [CWC^{+}23, HSX^{+}22, HLL^{+}23b, KS22, LCS18, LLD^{+}20, LWZ^{+}22, MZF24, SZY^{+}22, SZZW24, TLS24, WWL^{+}23, WZH^{+}23, YLZ^{+}19, ZSC^{+}22, ZPMT^{+}21]. Knowledge-Enhanced [SZY^{+}22]. Knowledge-Level [MZF24]. Koopman [PRR22]. KPIs [GSS23]. Kruskal [JRP^{+}22]. Kvasir [WTRK16].

Label [FWHM22, LYC^{+}24, LLW^{+}24, STZL22, YJH21, ZJZ^{+}19, Zhu15]. Label-Weighted [LYC^{+}24]. Labeled [GWLL24, PTL^{+}21, QWL^{+}23b]. Labeling [Liu15, XZGW21]. Lanczos [FYD^{+}19, FYD^{+}21]. Lanczos-Based [FYD^{+}19]. Landmark [JKF19]. Language [SZSY24]. Large [CM21, CLZ22, CHL23, CMJ^{+}24, CZ17b, CJY20, DFF^{+}23, HS20, HPWR20, HB16, HZZ^{+}21, JHT20, KKF16, KHdMR20, Kot15, Li15, LZL^{+}22, LLSL19, LQHC21, LZSL21, LJC^{+}22, Mat16, MCL22, MSC19, QWL^{+}23b, Sak20, SLC^{+}18, SHL^{+}22, SPN^{+}22, SWH^{+}22, WL22a, WCYL21, WZLS18b, Yan23, YYD^{+}23, YRL^{+}24, ZSS^{+}21, ZLT20, ZHW21, ZWS21]. Large-Scale [CM21, CLZ22, CJY20, HPWR20, HB16, HZZ^{+}21, JHT20, KHdMR20, Li15, LZL^{+}22, LLSL19, LJC^{+}22, Mat16, MSC19, Sak20, SLC^{+}18, SHL^{+}22, SWH^{+}22, WL22a, WZLS18b, Yan23, YYD^{+}23, ZSS^{+}21, ZWS21, CHL23, LZSL21]. Lasting [LYLJ22]. Latency [XLL^{+}18]. Latent [CLZ22, LZL^{+}21, LWQ^{+}23, SHL^{+}22, YHLS22, ZZLW21, ZJSL22]. Lattice [SD22]. Layer [YM24, ZJG^{+}20]. Leakage [LMW21]. Learned [ZRLZ23]. Learning [ADHS21, BLYM19, CZ17a, CSZ^{+}21, CLW^{+}22, CYXS23, CXW^{+}23, CYC^{+}23, CKL^{+}23, CLF^{+}18, DLL^{+}24, DHG21, FWHM22, GXQ^{+}20, GWLL24, HRFS^{+}24, HLLZ22a, HPESZ17, Hua15, HZ22, HLL^{+}23b, HLW^{+}23, HGY^{+}23, JRP^{+}22, JWS^{+}22, KFY^{+}23, LHWL24, LLS^{+}20, LBSL22, LZZ^{+}23, LTG^{+}24, LWZL24, LYC^{+}24, LLW^{+}24, LLC^{+}24, MTV21, MLK21, ONP^{+}24, PB23, QLZ24, SJS^{+}21, STZL22, SZZW24, SCG^{+}23, TJZ^{+}22, WH19, WTX^{+}22, WCW^{+}22, WWCK22, WHGW23, WLD^{+}23, WLXF23, WZL^{+}24a, WVD^{+}21, WWCZ22, XWI^{+}21, XZGW21, XZJT24, YWY^{+}22, Yan23, YM24, YCYL24, Yu15, ZLZ^{+}20, ZYCL20, ZYZZ20, ZLZL22, ZPT^{+}23, ZAL^{+}23, ZLC^{+}17, Zhu15, ZZZ^{+}22]. Learning-Based [ADHS21]. Least [TZCC22, WL22a]. Legal [ZGC^{+}24]. Lending [WZH^{+}23, ZWZ^{+}22]. Length [NLG^{+}23]. Level [HLW^{+}23, LZLJ22, MZF24, SBL^{+}22, WWSB16]. Leveled [LMW21]. Leveraging [ZPMT^{+}21]. LGM [LYLW23]. LGM-GNN [LYLW23]. LGRL [DLL^{+}24]. Libraries [ZWH^{+}22]. Library [CWGA17]. Life [LZC^{+}23]. Lifelong [SCG^{+}23]. Lifespans [DSS16]. Lifestyles [HBLK17]. Lightweight [CSW23, ZYB^{+}16, ZYX^{+}23]. Limited [ZLT20]. Line [GXQ^{+}20]. Linear [CM21, SLC^{+}18, TEOS17, YJH21]. Linearized [FZC20]. Link [LWZ^{+}22, LZZF23]. List [Ano19b]. List* [Ano17b]. Literature [LWH18, PTL^{+}21]. Litigation [WZL^{+}24b]. Live [CCD^{+}19a, DSS16]. Living [MLK21]. Load [HSEY20, SBR22]. Local [CWC^{+}23, CLHN22, DLL^{+}24, LOLL17, LYLW23, LND22, WYBH24, YWY^{+}22, ZCSZ23, ZDZC23, ZYX^{+}23]. Local-Global [DLL^{+}24]. Locality [BPP21, LS18b, XLC22, ZHW21]. Locality-Sensitive [ZHW21]. Localization [TLS24, ZFQ^{+}23]. Localness [HWTM20]. Location [HWTM20, LBL^{+}18, LQHC21, YWM^{+}21, ZZLW21]. Location-Based [HWTM20, ZZLW21]. Locations [ZQK17]. Logic [NLY22]. Logs [GB23]. Long [DSS16, YLZ^{+}22, YMJ^{+}23, YLL22, YLL23, ZZZ^{+}22]. Long-Short [YLL22, YLL23]. Long-Tailed [YMJ^{+}23]. Long-Term [ZZZ^{+}22]. LongArms [WZH^{+}23]. Longitudinal [FZ18]. Lookup [SWTX18]. Loops [BPP21]. Loss [PP22]. Lossy [LZD^{+}23, LWL^{+}23, LWQ^{+}23, ZZX^{+}21]. Low [ECG^{+}19, FLYL21, LQZ23, ZSHS17]. Low-Power [ECG^{+}19]. Low-Rank [ZSHS17]. LS [LW18, Yan15a]. LS-AMS [Yan15a]. LS-Decomposition [LW18].

LSTM [LCWS22, LLH^{+}23].

Machine [CYXS23, DHG21, Far21, LLS^{+}20, SRM17, SWDX20, SCG^{+}23, Yu15]. MAFI [JDC^{+}22]. Making [MSRS21, SZZW24]. Malicious [YWL^{+}22]. Malware [AK19, HZZ^{+}21]. Management [CZG^{+}19, HFX^{+}20, LZH^{+}24, SZA^{+}22, YZDZ19, YAG^{+}22, YWL^{+}22, ZWZ^{+}22]. Managing [MCM21]. MANETs [LDY^{+}22]. Manifold [WWCK22]. Manifold-Valued [WWCK22]. Many [LHWL24, LA20]. Many-to-Many [LA20]. Map [ZS22]. Mapping [SD22, TLS24, ZJZ^{+}19, ZMGIVO20]. MapReduce [DWSJ19, DFG^{+}19, Gou19, NGM16, SDR^{+}21, WJW^{+}18, WSW^{+}18, WCJI22, dSSN^{+}21]. Marketplaces [ZBL23]. Markov [HHX^{+}19, ZRY^{+}20]. Mask [BHX^{+}23]. Massive [AR18, CQH^{+}18, DHG21, LBL^{+}18, LRB^{+}20, LWS^{+}24, ZFLC19]. Matching [CWL^{+}21, KHdMR20, SX21, XLL^{+}22, ZS22, ZWS21]. Matching-Coalitional [SX21]. Mathematical [HSEY20, SZZ^{+}23]. Matrices [CLZ22, LQHC21, YHLS22]. Matrix [FZC20, HZF^{+}22, LQHC21, MBG22, WCYL21, ZZY^{+}23, ZYX^{+}23]. Max [JHT18]. Max-Min [JHT18]. Maximal [QLZ^{+}23]. MDIR [YLLC23]. Means [ZYCL22]. Measuring [MSRS21]. MEC [MCA^{+}20]. Mechanism [DEG^{+}22, FWS^{+}23, HLF^{+}22, WWC^{+}23, WWYH24, YLZ^{+}22]. Media [Agg15a, Agg15b, Agg16, ANP^{+}23, BED^{+}24, CSW18, Fu16, HBLK17, KJC^{+}18, LLT^{+}23, LRC20, LCWL21, MKJ^{+}24, PWW^{+}23, SSSL22, WQSA17, XML^{+}19, ZWV^{+}19]. Medical [CSZ^{+}21, LLD^{+}20, LWS^{+}24, SLL^{+}22, XYX^{+}23]. Meets [ZZX^{+}21]. Mega [Kit16]. Mega-City [Kit16]. Megabits [MCA^{+}20]. MemepiC [CZG^{+}19]. Memory [AR18, CZG^{+}19, EJCR22, IFG^{+}18, LYLW23, LZLJ22, Mat16, NGM16, STM22, SPN^{+}22, WXW^{+}22, YLZ^{+}22, ZLT20]. Memory-Based [LYLW23]. Merge [ZWLN22]. Merkle [WSW^{+}18]. Message [LMW21, Tan20, Yan20]. Meta [CYC^{+}23, ZSC^{+}22]. Meta-Learning [CYC^{+}23]. Metadata [SWTX18, ZQZ^{+}17]. MetaFlow [SWTX18]. Metagraph [LZC^{+}23]. Metagraph-Based [LZC^{+}23]. Metering [PWH16]. Method [CSW23, FZ18, HS20, HVVP21, LHY^{+}22, LLH^{+}23, MHR^{+}23, MKJ^{+}24, PWS^{+}19, YSM^{+}23, ZSWZ19, ZWD^{+}23, ZSC^{+}22]. Methodologies [SZA^{+}22, Zhe15]. Methodology [CCD^{+}19a, KdRFA21]. Methods [PRR22, Wan16, WBS^{+}23, WS17]. Metric [CXW^{+}23, GXQ^{+}20, WWCK22]. Metro [JFZ^{+}21, Kit16, ZZY^{+}23]. Metropolitan [HBLK17]. Microblogging [Liu15]. Microblogs [Yan15a]. Microfluidics [ICZ20]. Microfluidics-Driven [ICZ20]. Microscopy [LLCS17]. Millions [ZWH^{+}22]. Min [JHT18]. Minimization [ADHS21, DEG^{+}22, FZC20]. Mining [Agg15a, Agg15b, Agg16, BWW^{+}20, ÇLM17, DZS20, HB16, HXYZ20, JY22, LBL^{+}18, LBZ19, LYLJ22, LLW^{+}24, LZW^{+}22, QLZ^{+}23, QLW^{+}22, SAJP20, WQSA17, XDH23]. Misalignment [CZ17a]. Misinformation [EJCR22, LZS^{+}21]. Misled [SZSY24]. Missing [LCC^{+}23, SZY^{+}18, WFZ^{+}23, YLL22, YLLC23, YLL23]. MiSTR [LBSL22]. Mitosis [LLCS17]. Mitotic [NCS17]. Mix2SFL [ONP^{+}24]. Mixing [XYX^{+}23]. Mixture [LZMZ20]. Mixup [ONP^{+}24]. ML [HKKC22]. ML/DL [HKKC22]. MNL [CHL23]. Mobile [BS21, CMZL21, CYC^{+}23, FMD18, GQZ21, HX22, JFG17, SX21, XSYW22, ZFLC19, ZQK17]. Mobility [JFG17, KHdMR20, LZC^{+}23, SPYS23, XLC20]. MOD [ZZR20]. Modal [BSY^{+}24, LYD^{+}22, LLP^{+}23, LLN^{+}24, XML^{+}19]. Model [CHL23, DWXZ20, EJCR22, HHX^{+}19, HWW^{+}22, HLF^{+}22, HWB^{+}20, JFZ^{+}21, LBZ19, LZMZ20, PWC^{+}22, QFW^{+}23, SJS^{+}21, SZG^{+}23, SSY^{+}23, WHT18, WYK^{+}19, WGG^{+}23, WJW24, YSM^{+}23, YHLS22, ZLZ^{+}20, ZYCL20, ZRLZ23, ZYZ^{+}23, ZSL17]. Model-Agnostic [YSM^{+}23]. Modeling [ÇLM17, CGH^{+}22, CZ17b, GB23, JWS^{+}22, LTX16, NCS17, RDD^{+}21, SPYS23, ZHL^{+}17, ZZY^{+}23]. Modelling [AMAT23, CQH^{+}18, MCL22]. Models [HLF^{+}22, LWQ^{+}23, ZJSL22]. Modern [WYK^{+}19]. Modular [LZD^{+}23]. Modularity [YJH21]. Momentum [QYC^{+}23, ZJSL22]. Momentum-Incorporated [ZJSL22]. Monitoring [CZLL19, GZL^{+}20, SCW18, YLB^{+}19]. Mortality [WVD^{+}21]. Mosaicking [SCPS20]. Most [LBL^{+}18]. Movement [HLLZ22b]. Moving [CYC^{+}23]. MRMondrian [ZQD^{+}22]. MtMR [WSW^{+}18]. Multi [CW22, CLW^{+}22, CLGX23, CTC^{+}24, CLHN22, FY22, Fu16, GMZ^{+}22, HHZ^{+}24, HWW^{+}22, HVVP21, HCS^{+}22, LHWL24, LQZ23, LLP^{+}23, LLC^{+}24, SZY^{+}22, STZL22, SZZW24, TLS24, VP21, WYC^{+}20, WWCK22, WL22a, WCS^{+}23, WHGW23, WWL^{+}23, WLD^{+}23, WJW24, WWCZ22, WZL^{+}24b, XML^{+}19, XZGW21, YWG^{+}21, YZW^{+}23a, YM24, YKL^{+}23, ZLZ^{+}20, ZJG^{+}20, ZZX^{+}21, ZZL^{+}22, ZGS^{+}22, ZWZ^{+}22, ZAL^{+}23, ZZY^{+}23, ZHW21, ZYW^{+}23, Zhu15]. Multi-Agent [WHGW23]. Multi-Aspect [WZL^{+}24b]. Multi-Branch [ZGS^{+}22]. Multi-Center [ZYW^{+}23]. Multi-Clouds [ZZX^{+}21]. Multi-Density [FY22]. Multi-Dimension [WWL^{+}23]. Multi-Dimensional [HHZ^{+}24, LQZ23]. Multi-Discriminator [ZYW^{+}23]. Multi-Feature [CLGX23]. Multi-FPGAs [ZZL^{+}22]. Multi-Granularity [SZZW24]. Multi-Kernel [WWCK22]. Multi-Label [STZL22, Zhu15]. Multi-Layer [YM24, ZJG^{+}20]. Multi-Modal [LLP^{+}23, XML^{+}19]. Multi-Objective [ZWZ^{+}22]. Multi-Order [WYC^{+}20]. Multi-Party [LLC^{+}24, WCS^{+}23]. Multi-Rank [LQZ23]. Multi-Relation [CLHN22]. Multi-Relational [CW22]. Multi-Scale [HVVP21, TLS24, YWG^{+}21]. Multi-Semantics [HWW^{+}22]. Multi-Sensor [ZHW21]. Multi-Sensory [XZGW21]. Multi-Site [YZW^{+}23a]. Multi-Source [Fu16, VP21, ZAL^{+}23]. Multi-Task [CLW^{+}22, YKL^{+}23, ZLZ^{+}20]. Multi-Turn [WJW24]. Multi-View [CTC^{+}24, GMZ^{+}22, HCS^{+}22, LHWL24, SZY^{+}22, WL22a, WLD^{+}23, WWCZ22, YKL^{+}23, ZZY^{+}23]. Multidimensional [CSW23, MBS^{+}19, ZQD^{+}22]. Multilayered [YHLS22]. Multilayered-and-Randomized [YHLS22]. Multilevel [LWS^{+}24, ZZR20]. Multilinear [HLL^{+}23a]. Multimedia [AR18, ZSWZ19]. Multimodal [MXL^{+}20]. Multiobjective [CZLL19, GCL^{+}23]. Multiple [BLL^{+}20, CWS^{+}21, CQH^{+}18, JDC^{+}22, KZL^{+}22, LMW21, LCY^{+}22, LLY^{+}17, NLG^{+}23, SYS^{+}18, WWCK22, ZWS21]. Multiple-Length [NLG^{+}23]. Multiple-Perspective [KZL^{+}22]. Multiplication [MBG22]. Multivariable [SZY^{+}18]. Multivariate [SSAI21, WZL^{+}24a, ZYZ^{+}23]. Multiview [LBSL22]. Mutual [TLS24]. MWBS [LA20].

Named [PP22, ZGH^{+}22]. Named-Entity [PP22]. National [ZBB^{+}20]. Navigation [HSX^{+}22, XZGW21]. Nearest [GDY24, GRWL23, SWF16, XLP^{+}18, YWW^{+}21a]. Needs [NYL^{+}22]. Negative [LZL^{+}21, ZJSL22]. Neighbor [GDY24, GRWL23, SWF16, YWW^{+}21a]. Neighborhood [ZHL^{+}24]. Neighborhoods [LYL^{+}21]. Neighbors [XLP^{+}18]. Nested [HHX^{+}19]. Nested-Arc [HHX^{+}19]. Nesterov [QYC^{+}23, ZWD^{+}23]. Net [ZGC^{+}24, CZZ^{+}24]. Network [BHHDA19, CSZ^{+}21, CHL23, CCZ^{+}20, CLHN22, DDCHV21, EJCR22, FWWC23, FMD18, GZS23, HHZ^{+}24, HCZ^{+}24, HX22, JDC^{+}22, JGP^{+}22, JWS^{+}22, Kit16, LMPS21, LCS18, LYLW23, LZZ^{+}23, LLP^{+}23, MXL^{+}20, MZF24, NYL^{+}22, SLWV17, STZL22, SZXL22, SLL^{+}22, SWH^{+}22, SSY^{+}23, WTX^{+}22, WWSB16, XLL^{+}22, YWG^{+}21, YLLC23, YLL23, YZW^{+}23b, YXZL23, ZHL^{+}17, ZYZZ20, ZWS^{+}22, ZGS^{+}22, ZFQ^{+}23, ZWZ^{+}23, ZZ23, ZRY^{+}20, ZLT20, ZYW^{+}23]. Network-Based [MXL^{+}20, YLLC23, ZRY^{+}20]. Networks [CZLL19, CZ17b, CSL^{+}18, HLLZ22b, JFZ^{+}21, LBZ19, LYLJ22, LYW^{+}22, LZW^{+}23b, MSC19, QLW^{+}22, QWL^{+}23b, SWC^{+}22, SYS^{+}18, SPN^{+}22, STM^{+}20, WWL^{+}23, XLQ^{+}22, XDZ20, XLC20, YM24, YDL^{+}22, ZSWZ19, ZZLW21, ZLZ^{+}22, ZS22, ZSL^{+}22, ZAL^{+}23, ZWD^{+}23]. Neural [CSZ^{+}21, HCZ^{+}24, JRP^{+}22, LMPS21, LYLW23, LLP^{+}23, LZW^{+}23b, MXL^{+}20, SLL^{+}22, STM^{+}20, WWL^{+}23, WZL^{+}24b, XLQ^{+}22, YDL^{+}22, ZSL^{+}22, ZAL^{+}23, ZWD^{+}23]. Neurally [HSX^{+}22]. Neurally-Guided [HSX^{+}22]. News [DBMS24, EJCR22]. NGD [HH18]. NMF [WLD^{+}23]. NN [CWS^{+}21, ZWLN22]. Node [LZW^{+}23b, ZHKL24]. Nodes [LBZ19, LZZ^{+}23]. Nodes-Hyperedges [LZZ^{+}23]. Noise [LYC^{+}24, TCC21, WZZ^{+}19, ZCSW23]. Noise-Resistant [WZZ^{+}19]. Noiseless [Far23]. Non [JKF19, LZL^{+}21, WYK^{+}19, ZJSL22]. Non-Landmark [JKF19]. Non-Negative [LZL^{+}21, ZJSL22]. Non-Stationary [WYK^{+}19]. Nonlinear [KKKK16, LJC^{+}22]. Nonlinearly [LJC^{+}22]. Nonnegative [HZF^{+}22]. Nonparametric [BLYM19].

Normalization [SLL^{+}22]. NoSQL [SD18]. Novel [BPP21, CCD^{+}19a, CLGX23, LZW^{+}22, MSC19, XPC^{+}23, YJH21]. NPP [FYZ^{+}22].

O [YAG^{+}22]. O-Efficient [SWDX20]. O2O [XHZW21]. Object [CYC^{+}23, HZ22, LY17]. Objective [ZWZ^{+}22]. Objects [CZH^{+}17]. OD [ZZY^{+}23]. Off [GXQ^{+}20, XJQ^{+}21, ZLZ^{+}22]. Off-Deployment [ZLZ^{+}22]. Off-Line [GXQ^{+}20]. Offline [DDCHV21]. On-Demand [WHGW23]. On-Line [GXQ^{+}20]. On-the-Fly [DLL^{+}24]. On-Time [CYW^{+}22]. Online [CXT^{+}18, CLF^{+}18, DDCHV21, FY22, MZF24, SLWV17, WZH^{+}23, WLXF23, XWI^{+}21, ZZL^{+}24]. Onset [ZMGIVO20]. Operation [SAM^{+}20]. Operations [SDR^{+}21, YWG^{+}21]. OpinionRank [NLY22]. Opinions [DDCHV21]. Opportunistic [VP21, XLC20]. Optical [CLG^{+}20]. Optimal [CL23, GKR17, HSEY20, JZS^{+}20]. Optimization [CZLL19, CLZ22, FWHM22, GQP^{+}23, GCL^{+}23, JRP^{+}22, JWLY23, LWZL24, LWS^{+}24, LS18b, SD18, SGM21, SZZ^{+}23, WLZ^{+}20, YYD^{+}23]. Optimization-incorporated [CLZ22]. Optimized [HPESZ17]. Optimizing [LPT^{+}23, SWH^{+}22, TRB^{+}21, ZRY^{+}20]. Order [FYD^{+}19, FYD^{+}21, WYC^{+}20, ZYCL22, ZHKL24]. Ordered [XPC^{+}23]. ORE [ZRL22]. ORE-Based [ZRL22]. Organizational [ZZR20]. Oriented [CWPW20, JGP^{+}22, LB19, WWC^{+}23, ZSS^{+}21]. Origin [ZZZ^{+}22]. Origin-Destination [ZZZ^{+}22]. Orthogonal [FYD^{+}19, FYD^{+}21]. Ourselves [LYW^{+}22]. Out-Class [LRL^{+}17]. Outbreak [LMPS21, LZS^{+}21, XLL^{+}18]. Outcomes [GCA^{+}20]. Outgoing [Yan20]. Outlier [BLL^{+}20, HCS^{+}22]. Outsourced [FYD^{+}21, WHX^{+}23, XCV22, XLP^{+}18, ZRL22]. Outsourcing [CM21, LLSL19, LJC^{+}22, SLC^{+}18, ZZX^{+}21]. Overall [GZW23]. Overfitting [CLF^{+}18]. Overhead [ZYB^{+}16]. Overlapping [CCZ^{+}20, LND22, LYW^{+}22, ZSJ23]. Overview [FRS^{+}22, Zhe15, ZWC^{+}16]. Ownership [YAG^{+}22].

P [ZZR20]. P-MOD [ZZR20]. P2P [ZWZ^{+}22]. Package [BED^{+}24, CYW^{+}22]. Pandemic [CZQ^{+}21, FRS^{+}22, MSRS21]. Paradigm [Gou19]. Parallel [CYXS23, CSW23, FLYL21, GNPT21, LJC^{+}22, QLZ24, SAJP20, STM22, WJW24, ZYB^{+}16, ZSJ23]. Parallelism [LZLJ22]. Parametric [LLP^{+}23]. Part [Agg15b, CCD19b, CCD19c]. Partial [DH18, Wan16, WL22a]. Particle [CLZ22, GCL^{+}23, YYD^{+}23]. Partitioned [WHX^{+}23]. Partitioning [CZH^{+}17, ZZL^{+}22]. Party [LLC^{+}24, WCS^{+}23, ZWH^{+}22]. Passenger [Kit16, ZHL^{+}17, ZZY^{+}23]. Passive [KZL^{+}22]. Patchwise [XDZ20]. Patent [WZL^{+}24b]. Path [LRB^{+}20, ZDZC23, ZSC^{+}22]. Pathogenesis [ZPMT^{+}21]. Pathways [ZZZ^{+}18]. Pattern [BWW^{+}20, CWL^{+}21, ECG^{+}19, FZ18, HLLZ22b, IZ19, LZC^{+}23, MG23, XWI^{+}21, YLLC23, ZSWZ19]. Patterns [BS21, GCA^{+}20, JFG17, LSH^{+}23, LYL^{+}21, NLC17]. Peak [JHT20]. Peaks [ZDZC23]. People [NYL^{+}22, TLL^{+}23]. Perceptron [ZJG^{+}20]. Performance [CGH^{+}22, LTX16, LS18b, LB19, SWH^{+}22, ZAL^{+}23]. Performance-Oriented [LB19]. Personal [VP21]. Personality [XZJT24]. Personalize [XWI^{+}21]. Personalized [Fu16, MKJ^{+}24, TLL^{+}23, ZWZ^{+}22]. Perspective [AMAT23, BHHDA19, KZL^{+}22, ZWZ^{+}22]. Perspectives [WYLD18]. Perturbation [HZY^{+}21]. PES [EL22]. Petri [ZGC^{+}24].

Petuum [Yu15]. pg [ZZZ^{+}18]. pg-Causality [ZZZ^{+}18]. PHAED [WJW24]. Phase [LHY^{+}22, LLCS17, NGM16]. Phenomena [MBS^{+}19]. Phenomenon [HZZ^{+}21]. Phenometrics [ZMGIVO20]. Phenoregions [ZMGIVO20]. Phishing [LCWS22]. Phone [BS21, JFG17]. Physical [BWW^{+}20, FYD^{+}21, HLH^{+}20, HWTM20, HWC20, JLX^{+}23, LYG^{+}21, Sak20, WYC^{+}20]. Physical-Social-Aware [HWTM20]. Pick [GKR17]. Pipelines [PIMP17]. Pixel [YSM^{+}23]. Pixel-Wise [YSM^{+}23]. Places [VP21]. Planning [FMD18]. Plasma [WJS^{+}16]. Platform [ADOAH19, HZZ^{+}21, LZH^{+}24, MLQ^{+}19, Yu15]. PMM [DLD^{+}20]. PMU [CQH^{+}18]. POI [WLXF23]. Point [CCZ^{+}20, GKR17]. Policies [DWSJ19, QSC19]. Political [LZS^{+}21]. Pollutants [ZZZ^{+}18]. Pollution [HLLZ22a, HPWR20, YLL22, YLLC23, YLL23]. Polynomials [LQHC21]. Popular [Wan16]. Population [FWWC23]. Populations [LLCS17]. Positive [WCYL21]. Possibilistic [ZYCL22]. Posture [YLLC18]. Power [ECG^{+}19, JHT20, LLF^{+}19, WCJI22]. Power-Peak-Aware [JHT20]. Powerball [Yan23]. PPHOPCM [ZYCL22]. Practical [Ni15]. Practice [KGJØ18]. Pre [QFW^{+}23, WYLG21]. Pre-Authentication [WYLG21]. Pre-Training [QFW^{+}23]. Precondition [LWQ^{+}23]. Predict [LLS^{+}20]. Predictability [ZYZ^{+}23]. Predictable [SWC^{+}22]. Predicted [Cha16]. Predicting [GM20, LSH^{+}23, LLO^{+}22, WWYH24, YDL^{+}22, ZAL^{+}23]. Prediction [CGH^{+}22, HPESZ17, HX22, LOLL17, LWZ^{+}22, LCC^{+}23, LZD^{+}23, MXL^{+}20, SZY^{+}22, SZY^{+}18, SSY^{+}23, WZH^{+}23, WVD^{+}21, WZL^{+}24b, XWI^{+}21, YLZ^{+}19, YWM^{+}21, ZZLW21, ZLY^{+}22, ZQK17, ZZY^{+}23, ZZZ^{+}22]. Prediction-Based [LZD^{+}23]. Predictions [GSS23, SHH^{+}21]. Predictive [LTX16, LZW^{+}23a]. Predictor [FWHM22]. PredLife [LSH^{+}23]. Preference [XLL^{+}22]. Preferences [XLLC16]. Preprocessing [PIMP17]. Preservation [ZQD^{+}22]. Preserved [Ni15]. Preserving [ADOAH19, BPP21, DDGU22, DWXZ20, Far21, GQZ21, GLZ^{+}24, SRM17, WHX^{+}23, WWK^{+}21, XLC22, XWZ^{+}23, YLC^{+}22, ZYCL22]. Price [WXZ^{+}19]. Pricing [QSC19, ZBL23]. Principal [HLL^{+}23a]. Prior [CWC^{+}23, LQZ23]. Priority [EL22]. Privacy [ADOAH19, AMAT23, CMZL21, DDGU22, DWXZ20, Far21, Far23, FYZ^{+}22, GQZ21, GLZ^{+}24, HRFS^{+}24, HGY^{+}23, Ni15, TCC21, WBD^{+}20, WHX^{+}23, WWC^{+}23, WWK^{+}21, XJQ^{+}21, XWZ^{+}23, YLC^{+}22, ZQD^{+}22, ZYCL22, ZPT^{+}23, ZYX^{+}23]. Privacy-Accuracy [XJQ^{+}21]. Privacy-Aware [FYZ^{+}22, HGY^{+}23]. Privacy-Preserved [Ni15]. Privacy-Preserving [ADOAH19, Far21, GQZ21, GLZ^{+}24, WHX^{+}23, XWZ^{+}23, YLC^{+}22, ZYCL22]. Private [CWC^{+}23, DEG^{+}22, WCS^{+}23, XJQ^{+}21, YWRY21, ZCSZ23]. Privilege [ZZR20]. Privilege-Based [ZZR20]. Privileged [CZ17a]. PrivTDSI [ZCSZ23]. Probabilistic [CYW^{+}22, HLF^{+}22]. Probability [HX22]. Problems [LJC^{+}22]. Process [ZRY^{+}20]. Processes [LOLL17]. Processing [BZY20, CXT^{+}18, CWPW20, CL23, DLD^{+}20, DFG^{+}19, DLL^{+}16, Gou19, Kot15, LTX16, LRB^{+}20, LZH^{+}24, NHL18, PWC^{+}22, RDM^{+}20, SCW18, SLZ^{+}22, SBR22, SZG^{+}23, WZLS18a, ZSK19, ZBN^{+}20, ZZL^{+}22, ZLG^{+}21, ZBB^{+}20]. Processors [JGS^{+}19]. Product [WLZ^{+}20]. Profit [ZRY^{+}20]. Programming [HSEY20, LLSL19, LJC^{+}22]. Projection [YWW^{+}21a]. Promotion [WGG^{+}23]. Promotional [KJC^{+}18]. Prompt [KFY^{+}23]. Propagation [LY17, LZMZ20, NLC17, YJH21, ZJZ^{+}19].

Propensity [BS21]. Properties [SPN^{+}22]. Property [GLZ^{+}24]. Protect [LYW^{+}22]. Protecting [WTM18]. Protection [CMZL21]. Protocol [LLLZ21, SWC^{+}22, SBL^{+}22]. Protocol-Level [SBL^{+}22]. Provenance [HFX^{+}20]. Provision [WTRK16]. Proximal [FZC20]. Proximity [ZHKL24]. Proxy [BSY^{+}24, WYLG21]. Proxy-Based [BSY^{+}24]. Pruning [SDR^{+}21, YCYL24]. Public [DDCHV21, Far21, FYZ^{+}22, GLZ^{+}24, HZZ^{+}21, XML^{+}19]. Publishing [YWRY21]. Purpose [MCH^{+}22]. Purposes [JZS^{+}20]. Python [BED^{+}24].

QMSampler [SYS^{+}18]. Quality

[BHO21, GXQ^{+}20, LS18a, LLO^{+}22, PB23, SYS^{+}18, Tao15, YDL^{+}22, ZSL17]. QuantCloud [ZYYK18, ZSK19]. Quantitative [ZYYK18, ZSK19]. Quasi [QLZ^{+}23]. Quasi-Cliques [QLZ^{+}23]. Queries [DDGU22, GQP^{+}23, SD18, TAL^{+}17, YRL^{+}24]. Query [CWS^{+}21, DLZ^{+}19, GLZ^{+}24, LRB^{+}20, TAL^{+}17, XLP^{+}18, ZRL22]. Querying [ZDD^{+}23]. Question [LLD^{+}20, LS18a, SLWV17]. Questions [WH19]. Quorum [SGMB17]. Quorum-Replicated [SGMB17].

R [BHX^{+}23]. R-CNN [BHX^{+}23]. Radiation [IZ19]. Rainfall [ZJG^{+}20]. Random [LQHC21, YWW^{+}21a]. Randomized [TZCC22, YHLS22]. Range [ZRL22]. Rank [DZPM20, LQZ23, ZSHS17]. Ranking [MKJ^{+}24, YLZ^{+}19]. Ranking-Based [MKJ^{+}24]. Rate [CDZ^{+}24, WYK^{+}19, YKL^{+}23]. Rating [QXZ^{+}18, ZQK17]. Ratio [LWL^{+}23]. Ratios [ZHKL24]. Ratios-of-Eigenvectors [ZHKL24]. Ray [PTL^{+}21]. RCIVMM [ZS22]. RDF [BHO21, GQP^{+}23, Kot15, LPT^{+}23]. Re [WYLG21]. Re-Encryption [WYLG21].

Reachability [YRL^{+}24]. Real [ASYX21, CWPW20, SAM^{+}20, WJS^{+}16, YLB^{+}19]. Real-Time [ASYX21, CWPW20, SAM^{+}20, WJS^{+}16, YLB^{+}19]. Realize [FWS^{+}23]. Realtime [Yan15a]. Recognition [CKL^{+}23, FZ18, HLL^{+}23a, PP22, YLLC18, YMJ^{+}23, ZGC^{+}24, ZGH^{+}22]. Recommendation [CLL23, CJY20, DZPM20, DWSJ19, Fu16, LLC^{+}24, MZF24, QXZ^{+}18, QHC^{+}22, QYC^{+}23, WTX^{+}22, WWL^{+}23, WLXF23, WFZ^{+}23, WWSB16, XLLC16, XLL^{+}22, XHZW21, YXZL23, ZWZ^{+}22, ZYX^{+}23]. Recommender [LY17, LZL^{+}21, SHL^{+}22]. Reconstruction [FWWC23, ZGS^{+}22]. Records [LWS^{+}24, YTY^{+}21]. Recovery [ECG^{+}19, HHZ^{+}24, LQZ23, LW18, YLL22, YLLC23, YLL23]. Recruitment [WW21]. Recurrent [ZGC^{+}24]. Redshift [STM^{+}20]. Reduce [NGM16]. Reduced [LWQ^{+}23]. Reduction [FYD^{+}19, FYD^{+}21, SZJ^{+}22]. Redundancy [CLHH21]. Regarding [LCWL21]. Region [WHGW23]. Regions [DLD^{+}17]. Regression [GSS23, KKKK16, ZSHS17]. Regularization [LQZ23]. Regularized [HZY^{+}21, LCC^{+}23, LRL^{+}17, ZSHS17]. Reinforcement [LWZL24, TJZ^{+}22, WCW^{+}22, WHGW23]. Relation [CLHN22, WGYW23, YLZ^{+}19]. Relational [CW22, SD18, WVD^{+}21]. Relations [XLLC16]. Relationships [Hua15]. Relaxed [XLC22]. Release [Far21]. Relevant [WTRK16]. Reliable [TF22]. Remote [HXYZ20, TXH^{+}20, XDH23]. Repartitioning [LHY^{+}22]. Replicas [LCY^{+}22]. Replicated [SGMB17]. Report [TvMvdH23]. Representation [CHL23, DLL^{+}24, GMZ^{+}22, JZY22, QLZ24, STZL22, TJZ^{+}22, WTX^{+}22, ZYZZ20]. Representations [HXYZ20]. Requirement [WW21]. Research [WBD^{+}20, XHZW21]. Researcher [QFW^{+}23]. Researchers [DBAM17]. Residual [ZWZ^{+}23]. Resilient [GNPT21, LMW21]. Resistant [WZZ^{+}19]. Resolution [BHO21, HXYZ20]. Resource [DQWQ18, JHT18, LZH^{+}24, MKJ^{+}24, SX21, WCW^{+}22]. Resources [XLL^{+}18]. Resting [PIMP17]. Resting-State [PIMP17]. Result [LHY^{+}22]. Results [PRR22, SZJ^{+}22, Tao15]. Rethinking [CGZ23, WFZ^{+}23]. Retrieval [BSY^{+}24, EJCR22, Hua15, NLG^{+}23, TNX23, TXH^{+}20]. Retrieving [CZH^{+}17]. Return [ZWZ^{+}22]. Reverse [TLS24]. Review [LLCS17]. Reviewers [Ano17b, Ano18b, Ano19b]. Revocable [XCV22]. Rewiring [ZWD^{+}23]. RGSE [LZZF23]. Ride [GKR17, OVSF17]. Ridesharing [ZSS^{+}21]. Riemannian [CXW^{+}23, WWCK22]. Ring [YLB^{+}19]. Risk [DEG^{+}22, TCC21, ZWZ^{+}22]. Risk-Return [ZWZ^{+}22]. Road [ZS22]. Robotic [XZGW21]. Robust [Li15, LQZ23, LW18, LZZF23, LLN^{+}24, TRB^{+}21, WXZ^{+}19, ZWV^{+}19]. Robustness [ZZX^{+}21]. RoD [TCC21]. Role [JGP^{+}22]. Role-Oriented [JGP^{+}22]. Roles [LGW^{+}23]. Route [LLY^{+}17, ZS22]. Routing [SWC^{+}22]. Roving [ZHL^{+}17]. RPT [QFW^{+}23]. RTPM [CZZ^{+}24]. Rumor [LBSL22, SZSY24].

Safety [XML^{+}19]. Sampler [HS20].

Sampling [DLZ^{+}19, EL22, HX22, LLP^{+}23, SYS^{+}18, ZCSZ23]. Sanitization [BHHDA19]. Satellite [MSRS21, ZMGIVO20]. Satellite-Based [ZMGIVO20]. SBIR [DLL^{+}24]. Scalable [BTM16, CSW23, DFF^{+}23, FWS^{+}23, GDY24, HWC20, LHWL24, LLN^{+}24, LLF^{+}19, MG23, MBG22, ONP^{+}24, SWF16, SHL^{+}22, SDR^{+}21, SWTX18, TF22, WTRK16, WL22a, ZWV^{+}19, ZQD^{+}22, ZLT20, ZWS21]. Scale [CM21, CLZ22, CJY20, DHG21, HPWR20, HVVP21, HB16, HZZ^{+}21, JHT20, JDJ21, KHdMR20, Kot15, Li15, LZL^{+}22, LND22, LLSL19, LJC^{+}22, Mat16, MSRS21, MSC19, OVSF17, Sak20, SLC^{+}18, SHL^{+}22, SWH^{+}22, SCPS20, TLS24, WL22a, WZLS18b, YWG^{+}21, Yan23, YYD^{+}23, ZSS^{+}21, ZLT20, ZWS21, ZBB^{+}20, CHL23, LZSL21]. Scale-Free [ZLT20]. Scale-Out [Kot15]. ScaleJoin [GNPT21]. Scales [ZMGIVO20]. Scaling [WXW^{+}22]. Scanning [HZZ^{+}21]. SCCS [PLZL22]. Scene [HXYZ20, XDH23]. Scheduling [CWPW20, JHT20, LTX16, LXLF23, LLY^{+}18, TJZ^{+}22]. Schema [Ni15]. Scheme [FYZ^{+}22, HXW^{+}23, HLC^{+}18, LA20, TRB^{+}21, XSYW22]. Schemes [SXS^{+}23, ZGH^{+}22]. Scholar [Can16, LTTC16, LTTC17]. Scholarly [Gil16, XWBL17]. Science [SPH23]. Scientific [Cha16, LWH18, LWL^{+}23, LWQ^{+}23, XLLC16, ZGS^{+}22]. Score [CLL23]. SCOREH [ZHKL24]. SDN [CSL^{+}18]. SDN-Based [CSL^{+}18]. Search [Agg15a, Agg15b, Agg16, GRWL23, IFG^{+}18, JRP^{+}22, JDJ21, KS22, LPT^{+}23, Tao15, WQSA17, XWZ^{+}23, YWW^{+}21a, Yan15a, ZWS21]. Searching [Gil16, ZDZC23]. Seasonal [HX22]. SecFact [LZSL21]. Secret [TRB^{+}21]. Secure [ADHS21, CM21, CWS^{+}21, CDLW19, FYD^{+}19, FYD^{+}21, HXW^{+}23, HLC^{+}18, LLSL19, LZSL21, LJC^{+}22, PB23, SLC^{+}18, SLL^{+}21, XLP^{+}18, ZZR20, ZZX^{+}21, ZRL22]. Security [CZLL19, CCD19b, CCD19c, HRFS^{+}24, LLLZ21, LDY^{+}22, SX21, WTM18, WOD^{+}18]. Security-Aware [SX21]. SEEN [PWS^{+}19]. Segmentation [HZ22, XYX^{+}23, YWG^{+}21]. Seizure [HPESZ17, XWI^{+}21]. Selection [CGZ23, GKR17, JZS^{+}20, KJC^{+}18, LTG^{+}24, LLZ^{+}23, WGG^{+}23, YYD^{+}23, ZGS^{+}22, ZHL^{+}24, ZCH^{+}18, ZZL^{+}24]. Selective [PWS^{+}19]. Self [ECG^{+}19, GMZ^{+}22, LZZ^{+}23, LQHC21, LZW^{+}23b, YZW^{+}23a, ZWZ^{+}23]. Self-Adjoint [LQHC21]. Self-Attention [ZWZ^{+}23]. Self-Consistent [LZW^{+}23b]. Self-Recovery [ECG^{+}19]. Self-Representation [GMZ^{+}22]. Self-Supervised [LZZ^{+}23, YZW^{+}23a]. Semantic [BHO21, CZ17a, HSX^{+}22, HXYZ20, Hua15, Li15, TNX23]. Semantic-Based [BHO21]. Semantically [WTRK16]. Semantics [HWW^{+}22, SZSY24]. Semi [HHX^{+}19, LYC^{+}24, LZW^{+}23b, LLC^{+}24, XLQ^{+}22, XZGW21, XYX^{+}23, ZJZ^{+}19, Zhu15, CLW^{+}22]. Semi-Markov [HHX^{+}19]. Semi-Supervised [LYC^{+}24, LZW^{+}23b, LLC^{+}24, XLQ^{+}22, XZGW21, XYX^{+}23, ZJZ^{+}19, Zhu15]. semi-Traj2Graph [CLW^{+}22]. Sense [LCWL21]. Sensibility [LCWL21]. Sensing [HXYZ20, HWC20, JZS^{+}20, KZL^{+}22, PWS^{+}19, TXH^{+}20, XDH23, ZWV^{+}19]. Sensing-Based [JZS^{+}20]. Sensitive [HCZ^{+}24, LLH^{+}23, ZHW21]. Sensitivity [SZJ^{+}22]. Sensor [CZLL19, JZS^{+}20, VP21, WLF^{+}22, ZHL^{+}17, ZSS^{+}21, ZHW21]. Sensor-Cloud [WLF^{+}22]. Sensory [LW18, XZGW21]. Sentiment [HLW^{+}23]. Sentinel [SLV^{+}20, SCPS20, ZBB^{+}20]. Sentinel-1 [SLV^{+}20, SCPS20, ZBB^{+}20]. Separable [LLSL19, SDR^{+}21]. Seq2CASE [CLL23]. Sequence [CLL23, Fu16, ZGC^{+}24, ZDD^{+}23]. Sequences [LLCS17]. Sequential [MG23, YWRY21]. Serendipitous [WYXL21]. Series [DZS20, HVVP21, HMW^{+}22, SZY^{+}18, SSAI21, WZL^{+}24a, ZYZ^{+}23, ZBB^{+}20]. Servers [HHX^{+}19]. Service [QSC19, SWTX18, WYLD18, XHZW21, ZZX^{+}21, ZQK17]. Services [HWTM20, HMW^{+}22, PWH16, WYC^{+}20, WZH^{+}23]. Set [CXW^{+}23, LBL^{+}18, WWCK22, ZSC^{+}22]. Setting [HGY^{+}23]. Settings [JY22]. SGAE [XWG23a]. Shadow [WCJI22]. Share [TRB^{+}21]. Shared [CSW18, JHT18, JHT20, PLZL22, SPN^{+}22]. Shared-Memory [SPN^{+}22]. Sharing [FYZ^{+}22, GKR17, HXW^{+}23, LCS18, Ni15, OVSF17, SXS^{+}23, TRB^{+}21, WCS^{+}23, WWC^{+}23, ZZR20]. Shedding [SBR22]. Shift [RJS22]. Shifting [HSEY20]. Short [LLH^{+}23, YLZ^{+}22, YLL22, YLL23, ZJG^{+}20]. Short-Term [ZJG^{+}20]. Shot [SZZW24]. Shuffling [WCJI22]. SIESTA [MG23]. Signal [JPR^{+}20, MCO^{+}22]. Significant [QWL^{+}23b, TEOS17]. Similarity [CLF^{+}18, IFG^{+}18, JDJ21]. SIMiner [LBZ19]. Simple [CGZ23]. Simplexed [CGH^{+}22]. Simulating [OVSF17]. Simulation [ZGS^{+}22]. Singapore [JFG17]. Single [LCY^{+}22, SWDX20]. Site [YZW^{+}23a]. Situation [PWW^{+}23]. Situational [BHHDA19, WOD^{+}18, ZSWZ19]. Size [TRB^{+}21]. Sketching [CZZ^{+}24]. Skew [GNPT21]. Skew-Resilient [GNPT21]. SLA [SGMB17]. SLA-Aware [SGMB17]. Sleep [GZL^{+}20]. Sleepy [GZL^{+}20]. Small [LLS^{+}20]. Smart [CQH^{+}18, HSEY20, PWH16, PLZL22, SCW18, WXZ^{+}19, WOD^{+}18, SAM^{+}20]. Smartphone [GB23]. SmartQ [LS18a]. SMC [Ni15]. Snippets [YLLC18]. Social [ANP^{+}23, BED^{+}24, ÇLM17, CSW18, DDCHV21, FYD^{+}21, Fu16, HBLK17, HWTM20, HWC20, KS22, KJC^{+}18, LZS^{+}21, LCS18, LBZ19, LYG^{+}21, LYW^{+}22, LZW^{+}23a, LLT^{+}23, LRC20, LCWL21, NYL^{+}22, PWW^{+}23, SLWV17, SSSL22, SX21, TvMvdH23, WYC^{+}20, XML^{+}19, XSYW22, ZSWZ19, ZWV^{+}19, ZZLW21, ZQK17]. SocialQ&A [SLWV17]. Software [DSS16]. Solution [SX21]. Solved [FZC20]. SOREL [ZRL22]. Sorting [Mat16]. Source [Fu16, LRL^{+}17, VP21, ZAL^{+}23]. Sources [WBS^{+}23]. Space [BLL^{+}20, CLG^{+}20, DLL^{+}20, SJS^{+}21]. Spaces [IFG^{+}18]. Spam [ADOAH19]. Spamdoop [ADOAH19]. Spark [BTM16, BZY20, HPWR20, MBG22, SZY^{+}18, SGM21]. Sparse [CM21, CL23, FZC20, HB16, LZL^{+}22, MTV21, SLC^{+}18, SWF16, TZCC22, WHT18, WZH^{+}23, YHLS22, ZZX^{+}21, ZZLW21, ZAL^{+}23, ZSHS17]. Spatial [FWWC23, GB23, SAJP20, WW21, XDZ20, XWG^{+}23b, YSM^{+}23, YLLC23, ZRLZ23, ZOLP21, ZZ23, ZRY^{+}20]. Spatial-Attention [FWWC23]. Spatial-Temporal [XWG^{+}23b, YLLC23, ZZ23]. Spatio [LZH^{+}24, NLC17, SHH^{+}21, WJS^{+}16, YZW^{+}23b, ZSL17]. Spatio-Temporal [LZH^{+}24, NLC17, SHH^{+}21, WJS^{+}16, YZW^{+}23b, ZSL17]. Spatiotemporal [LCC^{+}23, RBM^{+}22, SAM^{+}20, SD22, SSSB16, ZZZ^{+}18]. Spatiotemporal-Interval [SSSB16]. Speaker [WJW24]. Speaker-Aware [WJW24]. Special [CCD19b, CCD19c, DLL^{+}20, HCZ19, JWS^{+}22, LTTC16, SSSL22, YLW^{+}20, ZLC^{+}17]. Species [SHH^{+}21]. Specific [HLLZ22a]. Spectral [HZY^{+}21, Li15, LZL^{+}22, XDZ20, YWW^{+}21b, ZHKL24]. Spectral-Spatial [XDZ20]. Spectroscopic [STM^{+}20]. Speed [LOLL17, TAL^{+}17, WZY^{+}18]. Split [ONP^{+}24, ZPT^{+}23]. SPMM [CL23]. Spoke [LLZ^{+}19]. Spouses [LLZ^{+}23]. Spread [SPYS23, SHH^{+}21]. Spring [ZMGIVO20]. Squares [TZCC22, WL22a]. SSD [WJW^{+}18]. SSD-Empowered [WJW^{+}18]. SSVM [PP22]. ST [YLLC23]. Stable [QLW^{+}22]. Stack [PWC^{+}22]. Stack-Centric [PWC^{+}22]. Stacked [XWG23a]. Star [STM22]. Stark [MBG22]. STaRS [OVSF17]. Start [YKL^{+}23]. State [CQH^{+}18, HHX^{+}19, PIMP17, SBR22, WYLD18, Yan17, Yan18, Yan19]. State-Aware [SBR22]. State-of-the-Art [WYLD18]. States [XLL^{+}18]. Static [JY22]. Stationary [WYK^{+}19]. Statistical [DDGU22, WZZ^{+}19]. Stealing [ZZL^{+}22].

Stem [LLCS17]. STGAN [YZW^{+}23b]. Stigmergy [LBZ19]. Stigmergy-Based [LBZ19]. Stochastic [LWS^{+}24, QLZ24, SHL^{+}22, Yan23]. Storage [CMJ^{+}24, CDLW19, ECG^{+}19, HKKC22, HLC^{+}18, HFX^{+}20, LLLZ21, LCY^{+}22, MBS^{+}19, SZC^{+}21, TF22, WZY^{+}18, YZDZ19, YAG^{+}22]. Stores [STM22, SD18]. Storing [KdRFA21]. Strassen [MBG22]. Strategies [JWLY23, KJC^{+}18]. Strategy [GQZ21, LWZL24, WWYH24, XHZW21, ZLZL22]. Stream [AR18, CWGA17, CXT^{+}18, GNPT21, JHT18, LTX16, LLH^{+}23]. Streaming [CQH^{+}18, EL22, LXLF23, LB19, SD22, WYBH24, ZZL^{+}24]. Streams [ANP^{+}23, FY22, FLYL21, Ni15, PWS^{+}19, SBR22, WLXF23, YDM^{+}23, YLB^{+}19]. Strength [MCO^{+}22]. Strong [XWZ^{+}23]. Structural [GMZ^{+}22, JWS^{+}22, LBSL22, QXZ^{+}18]. Structural-Temporal [LBSL22]. Structure [CTC^{+}24, HGY^{+}23, LZZF23, SCA^{+}17, SWF16, SLL^{+}21, SLZ^{+}22, WZL^{+}24a, Yan15a, YWY^{+}22, YCYL24]. Structure-Aware [SLZ^{+}22]. Structured [AI23, LMPS21, PP22, ZSHS17, ZPMT^{+}21]. Student [RDD^{+}21, ZAL^{+}23]. Studies [ICZ20]. Study [BED^{+}24, CZQ^{+}21, DBAM17, HZZ^{+}21, JFG17, LCS18, ZYZ^{+}23, ZPMT^{+}21]. Style [CLW^{+}22]. Sub [WZY^{+}18]. Sub-Datasets [WZY^{+}18]. Subgraph [JY22, WS17]. Subgraphs [LYLJ22]. Subjective [NLY22]. Sublinear [PWH16]. Submission [HZZ^{+}21]. Subsampled [TZCC22]. Subspace [MLK21]. Subspaces [SZJ^{+}22]. Substation [LLF^{+}19]. Success [TLL^{+}23]. Summarization [HHX^{+}19, SAM^{+}20]. Supervised [CLL23, LZZ^{+}23, LYC^{+}24, LYD^{+}22, LZW^{+}23b, LLC^{+}24, NLG^{+}23, XLQ^{+}22, XZGW21, XLC22, XYX^{+}23, YZW^{+}23a, ZJZ^{+}19, Zhu15]. Supervision [HZ22, KZSW23]. Supply [WLZ^{+}20].

Supplying [LS18a]. Support

[Far21, KAG^{+}22, MSRS21, SRM17].

Supporting [CDLW19]. Surface [ZBB^{+}20]. Surprise [WYXL21]. Surveillance [SCW18, XML^{+}19, YTY^{+}21]. Survey [APK22, AMAT23, CKL^{+}23, DFG^{+}19, HRFS^{+}24, JY22, JWLY23, JGP^{+}22, KAG^{+}22, LCY^{+}22, LZH^{+}24, PRR22, SLT^{+}22, SXS^{+}23, SZZ^{+}23, WBS^{+}23, XWBL17, YWRY21, ZYZZ20, ZBL23]. SVD [FYD^{+}19, FYD^{+}21]. SVM [SRM17, YLZ^{+}19]. Swarm [CLZ22, GCL^{+}23, YYD^{+}23]. Symmetric [ZJSL22]. Synchronous [JWLY23]. System [AK19, BSL^{+}22, CZG^{+}19, CMJ^{+}24, DEG^{+}22, ECG^{+}19, Gil16, ICZ20, JWLY23, LY17, LLZ^{+}19, LS18a, LDY^{+}22, MZF24, Ng16, PLZL22, RDM^{+}20, SLWV17, SXYL23, WWSB16, YLB^{+}19, ZSS^{+}21, ZYB^{+}16]. Systematic [TXH^{+}20]. Systems [AMAT23, BVAW^{+}16, BWW^{+}20, CM21, CKH16, GQP^{+}23, HCZ19, HLH^{+}20, HWB^{+}20, HWC20, HMW^{+}22, LPT^{+}23, LZLJ22, LZL^{+}21, MW22, PRR22, QYC^{+}23, Sak20, SLC^{+}18, SHL^{+}22, SWTX18, TF22, WYC^{+}20, WLF^{+}22, WXW^{+}22, YAG^{+}22, ZQZ^{+}17]. SZ3 [LZD^{+}23].

Tackling [PA20]. Tagging [ZGH^{+}22].

Tailed [YMJ^{+}23]. Tales [HBLK17]. Target [SZY^{+}22]. Targeted [ZDD^{+}23]. Task [CWPW20, CLW^{+}22, LLY^{+}18, TJZ^{+}22, YKL^{+}23, ZLZ^{+}20, ZCSW23]. Task-Duplication [CWPW20]. Tasks [KFY^{+}23, XSYW22]. Taxi [OVSF17, XLC20, ZHL^{+}17, ZSS^{+}21, ZLZL22, ZRY^{+}20]. Taxi-Passenger-Demand [ZHL^{+}17]. Taxonomy [APK22]. TBD [Ano18b]. Team [Ng16]. Technique [HD19, MSC19]. Techniques [WBS^{+}23]. Technologies [CLG^{+}20]. Technology [ZBN^{+}20]. Telecom [HCZ^{+}24]. Telemetries [BLL^{+}20]. Telescope [BHHDA19].

Temperature [ZMGIVO20]. Temporal [DHQ^{+}23, HVVP21, JY22, LBSL22, LZH^{+}24, LYLJ22, MCL22, NLC17, NCS17, QLW^{+}22, SHH^{+}21, SSAI21, WJS^{+}16, XWG^{+}23b, YSM^{+}23, YLLC23, YZW^{+}23b, ZGS^{+}22, ZZ23, ZSL17]. Temporary [TAL^{+}17]. Tendency [YLLC18]. Tensor [CZZ^{+}24, CW22, CJY20, FYD^{+}19, FYD^{+}21, HHZ^{+}24, LQZ23, LCC^{+}23, LYG^{+}21, WWCZ22, WZL^{+}24b]. Tensor-Train [LYG^{+}21]. Tensors [CW22]. Term [YLZ^{+}22, YLL22, YLL23, ZJG^{+}20, ZZZ^{+}22]. Terms [LCWL21]. Terrain [CZLL19]. Test [CQH^{+}18, LTG^{+}24]. Testing [BLL^{+}20]. Text [LLH^{+}23, YLB^{+}19]. Textual [DH18]. TgStore [CMJ^{+}24]. Their [ADF^{+}17]. Theoretic [CZ17b, HVVP21, RJS22, Sun15]. Theoretical [XJQ^{+}21]. Theory [KGJØ18, QXZ^{+}18, WWK^{+}21]. Theory-Based [QXZ^{+}18]. Thermal [LLY^{+}18]. Thermal-Aware [LLY^{+}18]. Things [HPESZ17, MW22, SYv^{+}19]. Third [ZWH^{+}22]. Third-Party [ZWH^{+}22]. Threat [AMAT23, CCD19b, CCD19c]. Three [NLY22]. Throughput [CXT^{+}18]. Ticks [MCA^{+}20]. Time [ASYX21, BVAW^{+}16, CWPW20, CYW^{+}22, CMJ^{+}24, DZS20, HVVP21, HMW^{+}22, SAM^{+}20, SZY^{+}18, SSAI21, WZL^{+}24a, WJS^{+}16, YJH21, YLB^{+}19, ZYZ^{+}23, ZBB^{+}20]. Time-Critical [BVAW^{+}16]. Time-Evolving [CMJ^{+}24]. Time-Series [ZYZ^{+}23]. Tool [KAG^{+}22]. Top [DZPM20, GQP^{+}23, LYLJ22]. Top- [LYLJ22, DZPM20, GQP^{+}23]. Topological [PIMP17]. Topology [XLC20]. Total [YAG^{+}22]. Trace [MCA^{+}20]. Traces [KHdMR20]. Tracking [ANP^{+}23, FY22, LY17, SZJ^{+}22, WJS^{+}16]. Trade [XJQ^{+}21]. Trade-Off [XJQ^{+}21]. Traffic [HX22, LOLL17, NLC17, SCA^{+}17, WZZ^{+}19, YLZ^{+}22, YTY^{+}21, YZW^{+}23b, ZFLC19, ZLZ^{+}22, ZWZ^{+}23, ZZ23]. Trafficking [KS22]. Train [LYG^{+}21].

Training [DWXZ20, GXQ^{+}20, JWLY23, PP22, QFW^{+}23, SWH^{+}22]. Traj2Graph [CLW^{+}22]. Trajectories [ASYX21, CYC^{+}23, GB23, LBL^{+}18, LRB^{+}20, LLY^{+}17]. Trajectory [CLW^{+}22, DLZ^{+}19, FZ18, KZL^{+}22]. Transaction [JY22]. Transactions [Ano15a, Ano15b, Ano17a, Ano18a, Ano19a, Ano20, Ano21, Ano22, Cra15, Yan15b]. Transfer [ADHS21, DZPM20, HLW^{+}23, ZLZ^{+}20]. Transferability [QYC^{+}23]. Transferable [QFW^{+}23]. Transform [TZCC22]. Transformations [JLX^{+}23]. Transformed [LQZ23]. Transformer [DBMS24]. Transit [LLZ^{+}19]. Transition [ZGC^{+}24]. Transport [DLD^{+}17]. Trapezoidal [MSC19]. Traps [XHZW21]. Travel [Fu16, MCH^{+}22]. Traversal [DLZ^{+}19]. Tree [GSS23, SPN^{+}22, WSW^{+}18]. Tree-Based [SPN^{+}22, WSW^{+}18]. Trees [KKKK16]. Trend [ZZY^{+}23]. Trends [TAL19]. Trial [WGG^{+}23]. Triangle [EL22, ZLT20]. Triangulation [CGH^{+}22]. Trident [MBS^{+}19]. Trust [XSYW22]. Trustworthiness [SWC^{+}22, SBL^{+}22, YWL^{+}22]. Trustworthy [HMW^{+}22, LS18a, NLY22, WLF^{+}22]. Truth [HWC20, ZWV^{+}19, ZCSZ23]. TS [CZZ^{+}24]. TS-RTPM-Net [CZZ^{+}24]. Tuned [XZJT24]. Tuning [SGMB17]. Turn [WJW24]. TUSQ [ZDD^{+}23]. Tutorial [PWW^{+}23]. Tweets [Kit16]. Twitter [SD22]. Two [GCA^{+}20, HBLK17, LHY^{+}22, ONP^{+}24]. Two-Phase [LHY^{+}22]. Two-Way [GCA^{+}20, ONP^{+}24]. Type [ZZL^{+}24]. Types [LGW^{+}23]. Typing [QWL^{+}23a].

Ubiquitous [LLO^{+}22, ZSL^{+}22]. UGCC [LLT^{+}23]. Uncertain [JFZ^{+}21, QLZ^{+}23]. Uncertainty [HWC20, WFZ^{+}23, XYX^{+}23]. Uncertainty-Aware [HWC20, WFZ^{+}23].

Uncertainty-Guided [XYX^{+}23]. Unconstrained [LZL^{+}21, ZGH^{+}22]. Uncovering [LND22]. Understand [HBLK17]. Understanding [Agg15a, Agg15b, Agg16, JGS^{+}19, LYL^{+}21, LZW^{+}22, WQSA17, ZFLC19, ZWH^{+}22, ZAL^{+}23, ZLC^{+}17]. Undirected [HLF^{+}22]. Unexpected [SD22]. Unified [BSL^{+}22, BZY20, CZG^{+}19, GM20, GXQ^{+}20, Ng16]. Unit [WVD^{+}21, ZGC^{+}24]. Universal [KKKK16]. Unlabeled [GWLL24]. Unsupervised [CGZ23, LLN^{+}24, WL22a, ZDZC23]. Unveiling [WZY^{+}18]. Update [ZLG^{+}21]. Updating [SPN^{+}22]. Urban [ÇLM17, DLD^{+}17, LLZ^{+}19, LYL^{+}21, LLO^{+}22, MCH^{+}22, SCA^{+}17, WYK^{+}19, XLC20, XWG^{+}23b, ZFLC19, ZSS^{+}21, ZS22, ZLY^{+}22, ZWC^{+}16, ZMS17, ZSL17, ZZZ^{+}18]. Use [BS21, LCWL21]. User [CSW18, KAG^{+}22, KHdMR20, LZW^{+}23a, LLT^{+}23, LLF^{+}19, PWH16, SLT^{+}22, WWL^{+}23, YWL^{+}22, ZSWZ19, ZZLW21, ZSS^{+}21, WGG^{+}23]. User-Item [WWL^{+}23]. User-Oriented [ZSS^{+}21]. User-Shared [CSW18]. User-Substation [LLF^{+}19]. Users [FYZ^{+}22, LZW^{+}22, LCWL21, ZQK17]. Using [BTM16, BED^{+}24, CZ17b, DWSJ19, DFG^{+}19, FMD18, GB23, GSS23, HD19, HBLK17, JZY22, KKKK16, KKF16, MBG22, MSC19, NLY22, RBM^{+}22, SPYS23, SLV^{+}20, TCC21, VP21, WYK^{+}19, WZH^{+}23, XLL^{+}18, XML^{+}19, XPC^{+}23, YSM^{+}23, ZGC^{+}24, ZBN^{+}20, ZJG^{+}20, ZAL^{+}23, ZOLP21, dSSN^{+}21, CWPW20]. USTF [Ng16]. Utility [YWRY21, ZDD^{+}23].

Validation [DHG21]. Valued [NLY22, WWCK22, XPC^{+}23]. Variability [SSAI21]. Variational [GCL^{+}23, LZMZ20]. Various [KFY^{+}23, Sak20]. Vector [Far21, SRM17]. Vectors [IFG^{+}18, XLP^{+}18]. Vehicles [PLZL22].

Vehicular [HLLZ22b, SWC^{+}22]. Velocity [FLYL21]. Velocity-Aware [FLYL21]. Venue [ZZLW21]. Verifiable [CWL^{+}21, HLC^{+}18, XCV22]. Verification [DHG21]. Verifications [WSW^{+}18]. Vertical [LLW^{+}24]. Vertically [WHX^{+}23]. VHR [XDH23]. Via [LCC^{+}23, CLW^{+}22, CLHN22, FZC20, FWS^{+}23, GXQ^{+}20, GZL^{+}20, HHZ^{+}24, HWW^{+}22, HLF^{+}22, HPESZ17, HCS^{+}22, HFX^{+}20, LHWL24, LLP^{+}23, LLT^{+}23, LLN^{+}24, QFW^{+}23, SHL^{+}22, SSY^{+}23, TZCC22, WGYW23, YCYL24, YMJ^{+}23, YLL23, ZCSZ23, ZCSW23, ZWD^{+}23]. Video [ECG^{+}19, LYL^{+}21, SCW18]. Videos [CLHH21, LZW^{+}22, Wan16]. View [CTC^{+}24, GMZ^{+}22, HCS^{+}22, LHWL24, SZY^{+}22, WL22a, WLD^{+}23, WWCZ22, YKL^{+}23, ZZY^{+}23]. Viewing [LYL^{+}21]. Viral [ZPMT^{+}21]. Virtual [BED^{+}24]. Virtualized [WTM18]. Visual [CZ17a, CKL^{+}23, CSW23, Hua15, HH18, Kit16, LWH18, LLY^{+}17, SJS^{+}21, SHH^{+}21, SLT^{+}22, SZZ^{+}23, ZWC^{+}16]. Visual-Semantic [CZ17a]. Visualization [FZ18, MCO^{+}22, PRR22, PA20, SZZ^{+}23, SXYL23, ZOLP21]. Visualizing [RDD^{+}21]. Viziometrics [LWH18]. Vol [Ano17a, Ano18a, Ano19a, Ano20, Ano21, Ano22]. Volume [MCL22, YTY^{+}21]. Voluminous [RBM^{+}22]. Volunteered [RDM^{+}20, TF22]. Voting [ZS22]. Vulnerability [TAL19]. VWAP [LWZL24].

Walk [HS20]. Wallet [SXYL23]. Warning [TvMvdH23]. Watchdog [SCG^{+}23]. Water [LLO^{+}22]. Way [GCA^{+}20, ONP^{+}24]. Weak [HZ22]. Weakly [CLL23, LYD^{+}22, PTL^{+}21, Zhu15]. Weather [DLD^{+}17]. Web [QHC^{+}22, ADF^{+}17, CKH16, CLHH21, ESC20, QHC^{+}22, SVYY16, WTRK16, XLL^{+}18, Yao15, ZWS21]. Web-Based [CKH16]. Website [LLD^{+}20, NLY22].

Weighted [CDZ^{+}24, CHL23, LYC^{+}24]. Welcome [Cra15]. Who [WGG^{+}23]. Wi [KZL^{+}22, JPR^{+}20]. Wi-Fi [KZL^{+}22, JPR^{+}20]. WiFi [GZL^{+}20]. WiFind [JPR^{+}20]. Will [DSS16]. Windows [HX22]. Wireless [DWXZ20, GZL^{+}20, HWB^{+}20, LZMZ20, LA20, YLW^{+}20, ZYCL20]. Wise [YSM^{+}23]. within [ZBB^{+}20]. Without [ZZL^{+}24, Liu15]. Word [JZY22]. Work [ZZL^{+}22]. Worker [WW21]. Workflow [SPH23]. Workflows [CWPW20]. Workload [YAG^{+}22]. Workloads [RDM^{+}20]. Wrangling [KAB^{+}21]. WukaStore [TF22].

X [PTL^{+}21]. X-Ray [PTL^{+}21].

References

Agreste:2017:ECA

[ADF^{+}17] S. Agreste, P. De Meo, G. Fiumara, G. Piccione, S. Piccolo, D. Rosaci, G. M. L. Sarné, and A. V. Vasilakos. An empirical comparison of algorithms to find communities in directed graphs and their application in Web data analytics. IEEE Transactions on Big Data, 3(3): 289–306, September 2017. CO-DEN ???? ISSN 2332-7790.

Aledhari:2021:DLB

[ADHS21] Mohammed Aledhari, Marianne Di Pierro, Mohamed Hefeida, and Fahad Saeed. A deep learning-based data minimization algorithm for fast and secure transfer of big genomic datasets. IEEE Transactions on Big Data, 7(2):271–284, June 2021. CODEN ???? ISSN 2332-7790.

AlMahmoud:2019:SPP

[ADOAH19] A. AlMahmoud, E. Damiani, H. Otrok, and Y. Al-Hammadi. Spamdoop: a privacy-preserving Big Data platform for collaborative spam detection. IEEE Transactions on Big Data, 5(3):293–304, September 2019. CODEN ???? ISSN 2332-7790.

Aggarwal:2015:GEBa

[Agg15a] Charu C. Aggarwal. Guest editorial: Big media data: Understanding, search, and mining. IEEE Transactions on Big Data, 1(3):82– 83, September 2015. CO-DEN ???? ISSN 2332-7790.

URL http://www.computer.
org/csdl/trans/bd/2015/03/
07355488.pdf.

Aggarwal:2015:GEBb

[Agg15b] Charu C. Aggarwal. Guest editorial: Big media data: Understanding, search, and mining (Part 2). IEEE Transactions on Big Data, 1(4): 151, December 2015. CO-DEN ???? ISSN 2332-7790.

URL http://www.computer.
org/csdl/trans/bd/2015/04/
07395031.pdf.

Aggarwal:2016:GEB

[Agg16] Charu C. Aggarwal. Guest editorial: Big media data: Understanding, search, and mining. IEEE Transactions on Big Data, 2(1):31, March 2016. CO-DEN ???? ISSN 2332-7790.

URL http://www.computer.
org/csdl/trans/bd/2016/01/
07473987.pdf.

Ahmed:2023:SAT

[AI23] Hameeza Ahmed and Muhammad Ali Ismail. A structured approach towards big data identification. IEEE Transactions on Big Data, 9(1):147–159, February 2023. ISSN 2332-7790.

Abawajy:2019:ICF

[AK19] J. H. Abawajy and A. Kelarev.

Iterative classifier fusion system for the detection of Android malware. IEEE Transactions on Big Data, 5(3):282– 292, September 2019. CODEN ???? ISSN 2332-7790.

Azam:2023:DPT

[AMAT23] Naila Azam, Lito Michala, Shuja Ansari, and Nguyen Binh Truong. Data privacy threat modelling for autonomous systems: a survey from the GDPR’s perspective. IEEE Transactions on Big Data, 9 (2):388–414, April 2023. ISSN 2332-7790.

Anonymous:2015:ITBa

[Ano15a] Anonymous. IEEE Transactions on Big Data. IEEE Transactions on Big Data, 1 (1):47, March 2015. CO-DEN ???? ISSN 2332-7790.

URL http://www.computer.
org/csdl/trans/bd/2015/01/
07152999.pdf.

Anonymous:2015:ITBb

[Ano15b] Anonymous. IEEE Transactions on Big Data. IEEE Transactions on Big Data, 1 (1):48, March 2015. CO-DEN ???? ISSN 2332-7790.

                                        URL http://www.computer.
                                        org/csdl/trans/bd/2015/01/
                                        07153000.pdf.

Anonymous:2017:IIT

[Ano17a] Anonymous. 2016 index IEEE Transactions on Big Data vol.

2. IEEE Transactions on Big Data, 3(1):1–6, March 2017. CODEN ???? ISSN 2332-7790.

Anonymous:2017:RL

[Ano17b] Anonymous. 2016 reviewers list*. IEEE Transactions on Big Data, 3(1):118–123, March 2017. CODEN ???? ISSN 2332-7790.

Anonymous:2018:IIT

[Ano18a] Anonymous. 2017 index IEEE Transactions on Big Data vol.

3. IEEE Transactions on Big Data, 4(1):1–7, March 2018. CODEN ???? ISSN 2332-7790.

Anonymous:2018:TR

[Ano18b] Anonymous. TBD 2017 reviewers. IEEE Transactions on Big Data, 4(1):138–140, March 2018. CODEN ???? ISSN 2332-7790.

Anonymous:2019:IIT

[Ano19a] Anonymous. 2018 index IEEE Transactions on Big Data vol.

4. IEEE Transactions on Big Data, 5(1):1–9, March 2019. CODEN ???? ISSN 2332-7790.

Anonymous:2019:RL

[Ano19b] Anonymous. 2018 reviewers list. IEEE Transactions on Big Data, 5(1):106–118, March 2019. CODEN ???? ISSN 2332-7790.

Anonymous:2020:IIT

[Ano20] Anonymous. 2019 index IEEE Transactions on Big Data vol.

5. IEEE Transactions on Big Data, 6(1):1–9, March 2020. CODEN ???? ISSN 2332-7790.

Anonymous:2021:IIT

[Ano21] Anonymous. 2020 index IEEE Transactions on Big Data vol.

6. IEEE Transactions on Big Data, 7(1):1–13, March 2021. CODEN ???? ISSN 2332-7790.

Anonymous:2022:IIT

[Ano22] Anonymous. 2021 index IEEE Transactions on Big Data vol.

7. IEEE Transactions on Big Data, 8(1):1–16, February 2022. CODEN ???? ISSN 2332-7790.

Anwar:2023:TEC

[ANP^{+}23] Tarique Anwar, Surya Nepal, Cecile Paris, Jian Yang, Jia Wu, and Quan Z. Sheng. Tracking the evolution of clusters in social media streams. IEEE Transactions on Big Data, 9 (2):701–715, April 2023. ISSN 2332-7790.

Alazzawe:2022:EBD

[APK22] Anis Alazzawe, Amitangshu Pal, and Krishna Kant. Efficient big-data access: Taxonomy and a comprehensive survey. IEEE Transactions on Big Data, 8(2):356–376, February 2022. CODEN ???? ISSN 2332-7790.

Antaris:2018:MSI

[AR18] S. Antaris and D. Rafailidis. In-memory stream indexing of massive and fast incoming multimedia content. IEEE Transactions on Big Data, 4(1):40– 54, March 2018. CODEN ???? ISSN 2332-7790.

Aoki:2021:BEE

[ASYX21] Shunsuke Aoki, Kaoru Sezaki, Nicholas Jing Yuan, and Xing Xie. BusBeat: Early event detection with real-time bus GPS trajectories. IEEE Transactions on Big Data, 7(2):371– 382, June 2021. CODEN ???? ISSN 2332-7790.

Bour:2024:APP

[BED^{+}24] Charline Bour, Abir Elbeji, Luigi De Giovanni, Adrian Ahne, and Guy Fagherazzi. AL-TRUIST: a Python package to emulate a virtual digital cohort study using social media data. IEEE Transactions on Big Data, 10(4):568–575, August 2024. ISSN 2332-7790.

Bou-Harb:2019:BDS

[BHHDA19] Elias Bou-Harb, Martin Husák, Mourad Debbabi, and Chadi Assi. Big data sanitization and cyber situational awareness: a network telescope perspective. IEEE Transactions on Big Data, 5(4):439–453, December 2019. CODEN ???? ISSN 2332-7790.

Benbernou:2021:SBE

[BHO21] Salima Benbernou, Xin Huang, and Mourad Ouziri. Semanticbased and entity-resolution fusion to enhance quality of big RDF data. IEEE Transactions on Big Data, 7(2):436– 450, June 2021. CODEN ???? ISSN 2332-7790.

Bi:2023:IRC

[BHX^{+}23] Xiuli Bi, Jinwu Hu, Bin Xiao, Weisheng Li, and Xinbo Gao. IEMask R-CNN: Informationenhanced mask R-CNN. IEEE Transactions on Big Data, 9 (2):688–700, April 2023. ISSN 2332-7790.

Barreyre:2020:MTO

[BLL^{+}20] Clémentine Barreyre, Béatrice Laurent, Jean-Michel Loubes, Loic Boussouf, and Bertrand Cabon. Multiple testing for outlier detection in space telemetries. IEEE Transactions on Big Data, 6(3):443–451, September 2020. CODEN ???? ISSN 2332-7790.

Bruce:2019:NDL

[BLYM19] S. Bruce, Z. Li, H. Yang, and S. Mukhopadhyay. Nonparametric distributed learning architecture for big data: Algorithm and applications. IEEE Transactions on Big Data, 5(2): 166–179, June 2019. CODEN ???? ISSN 2332-7790.

Bohm:2021:NHC

[BPP21] Christian Böhm, Martin Perdacher, and Claudia Plant. A novel Hilbert curve for cachelocality preserving loops. IEEE Transactions on Big Data, 7(2): 241–254, June 2021. CODEN ???? ISSN 2332-7790.

Bati:2021:AIA

[BS21] Ghassan F. Bati and Vivek K. Singh. Altrumetrics: Inferring altruism propensity based on mobile phone use patterns. IEEE Transactions on Big Data, 7(2):397–406, June 2021. CODEN ???? ISSN 2332-7790.

Biswas:2022:DDA

[BSL^{+}22] Sujit Biswas, Kashif Sharif, Fan Li, Iqbal Alam, and Saraju P. Mohanty. DAAC: Digital asset access control in a unified blockchain based e-health system. IEEE Transactions on Big Data, 8(5):1273–1287, October 2022. ISSN 2332-7790.

Bai:2024:PBG

[BSY^{+}24] Yibing Bai, Zhenqiu Shu, Jun Yu, Zhengtao Yu, and Xiao-Jun Wu. Proxy-based graph convolutional hashing for cross-modal retrieval. IEEE Transactions on Big Data, 10(4):371–385, August 2024. ISSN 2332-7790.

Bharill:2016:FBS

[BTM16] N. Bharill, A. Tiwari, and A. Malviya. Fuzzy based scalable clustering algorithms for handling big data using Apache Spark. IEEE Transactions on Big Data, 2(4):339–352, December 2016. CODEN ???? ISSN 2332-7790.

Basanta-Val:2016:ATC

[BVAW^{+}16] P. Basanta-Val, N. C. Audsley, A. J. Wellings, I. Gray, and N. Fernández-García. Architecting time-critical big-data systems. IEEE Transactions on Big Data, 2(4):310–324, December 2016. CODEN ???? ISSN 2332-7790.

Bhuiyan:2020:EDT

[BWW^{+}20] Md Zakirul Alam Bhuiyan, Jie Wu, Gary M. Weiss, Thaier Hayajneh, Tian Wang, and Guojun Wang. Event detection through differential pattern mining in cyber-physical systems. IEEE Transactions on Big Data, 6(4):652–665, December 2020. CODEN ???? ISSN 2332-7790.

Brahem:2020:AUA

[BZY20] Mariem Brahem, Karine Zeitouni, and Laurent Yeh. ASTROIDE: a unified astronomical Big Data processing engine over Spark. IEEE Transactions on Big Data, 6(3):477–491, September 2020. CODEN ???? ISSN 2332-7790.

Candan:2016:GEB

[Can16] K. Selcuk Candan. Guest editorial: Big scholar data discovery and collaboration. IEEE Transactions on Big Data, 2 (1):1–2, March 2016. CO-DEN ???? ISSN 2332-7790.

URL http://www.computer.
org/csdl/trans/bd/2016/01/
07473985.pdf.

Castiglione:2019:NMA

[CCD^{+}19a] Aniello Castiglione, Giuseppe Cattaneo, Giancarlo De Maio, Alfredo De Santis, and Gianluca Roscigno. A novel methodology to acquire live Big Data evidence from the cloud. IEEE Transactions on Big Data, 5(4): 425–438, December 2019. CO-DEN ???? ISSN 2332-7790.

Choo:2019:SIBa

[CCD19b] K. R. Choo, M. Conti, and A. Dehghantanha. Special issue on big data applications in cyber security and threat intelligence: Part 1. IEEE Transactions on Big Data, 5(3):279– 281, September 2019. CODEN ???? ISSN 2332-7790.

Choo:2019:SIBb

[CCD19c] Kim-Kwang Raymond Choo, Mauro Conti, and Ali Dehghantanha. Special issue on Big Data applications in cyber security and threat intelligence Part 2. IEEE Transactions on Big Data, 5(4):423–424, December 2019. CODEN ???? ISSN 2332-7790.

Cheng:2020:OCC

[CCZ^{+}20] Jiujun Cheng, Minjun Chen, MengChu Zhou, Shangce Gao, Chunmei Liu, and Cong Liu. Overlapping community change-point detection in an evolving network. IEEE Transactions on Big Data, 6(1):189– 200, March 2020. CODEN ???? ISSN 2332-7790.

Cui:2019:ABS

[CDLW19] H. Cui, R. H. Deng, Y. Li, and G. Wu. Attribute-based storage supporting secure deduplication of encrypted data in cloud. IEEE Transactions on Big Data, 5(3):330–342, September 2019. CODEN ???? ISSN 2332-7790.

Cai:2024:FFW

[CDZ^{+}24] Jianghe Cai, Yuhui Deng, Yi Zhou, Jiande Huang, and Geyong Min. FIG: Featureweighted information granules with high consistency rate. IEEE Transactions on Big Data, 10(4):400–414, August 2024. ISSN 2332-7790.

Chen:2022:SAD

[CGH^{+}22] Yuxing Chen, Peter Goetsch, Mohammad A. Hoque, Jiaheng Lu, and Sasu Tarkoma. dsimplexed: Adaptive Delaunay triangulation for performance modeling and prediction on big data analytics. IEEE Transactions on Big Data, 8(2):458– 469, February 2022. CODEN ???? ISSN 2332-7790.

Chang:2023:REU

[CGZ23] Heng Chang, Jun Guo, and Wenwu Zhu. Rethinking embedded unsupervised feature selection: a simple joint approach. IEEE Transactions on Big Data, 9(1):380–387, February 2023. ISSN 2332-7790.

Chawla:2016:CSI

[Cha16] Nitesh V. Chawla. Can scientific impact be predicted? IEEE Transactions on Big Data, 2(1): 18–30, March 2016. CODEN ???? ISSN 2332-7790.

Chen:2023:MHE

[CHL23] Minzhi Chen, Chunlin He, and Xin Luo. MNL: a highlyefficient model for large-scale dynamic weighted directed network representation. IEEE Transactions on Big Data, 9(3): 889–903, June 2023. ISSN 2332-7790.

Chou:2020:FTF

[CJY20] Szu-Yu Chou, Jyh-Shing Roger Jang, and Yi-Hsuan Yang. Fast tensor factorization for largescale context-aware recommendation from implicit feedback. IEEE Transactions on Big Data, 6(1):201–208, March 2020. CODEN ???? ISSN 2332-7790.

Chen:2016:ABD

[CKH16] H. Chen, R. Kazman, and S. Haziyev. Agile big data analytics for Web-based systems: An architecture-centric approach. IEEE Transactions on Big Data, 2(3):234–248, September 2016. CODEN ???? ISSN 2332-7790.

Chen:2023:SVA

[CKL^{+}23] Dongpan Chen, Dehui Kong, Jinghua Li, Shaofan Wang, and Baocai Yin. A survey of visual affordance recognition based on deep learning. IEEE Transactions on Big Data, 9(6):1458– 1476, December 2023. ISSN 2332-7790.

Choi:2023:DSE

[CL23] Unho Choi and Kyungyong Lee.

Dense or sparse : Elastic SPMM implementation for optimal bigdata processing. IEEE Transactions on Big Data, 9(2):637– 652, April 2023. ISSN 2332-7790.

Cong:2018:OSL

[CLF^{+}18] Y. Cong, J. Liu, B. Fan, P. Zeng, H. Yu, and J. Luo. Online similarity learning for big data with overfitting. IEEE Transactions on Big Data, 4(1): 78–89, March 2018. CODEN ???? ISSN 2332-7790.

Coluccia:2020:OCI

[CLG^{+}20] Giulio Coluccia, Cinzia Lastri, Donatella Guzzi, Enrico Magli, Vanni Nardino, Lorenzo Palombi, Ivan Pippi, Valentina Raimondi, Chiara Ravazzi, Florin Garoi, Daniela Coltuc, Raffaele Vitulli, and Alessandro Zuccaro Marchi. Optical compressive imaging technologies for space Big Data. IEEE Transactions on Big Data, 6(3): 430–442, September 2020. CO-DEN ???? ISSN 2332-7790.

Chen:2023:NMF

[CLGX23] Huanyu Chen, Weisheng Li, Xinbo Gao, and Bin Xiao. Novel multi-feature fusion facial aesthetic analysis framework. IEEE Transactions on Big Data, 9(5):1302–1320, October 2023. ISSN 2332-7790.

Chen:2021:EEC

[CLHH21] Yixin Chen, Dongsheng Li, Yu Hua, and Wenbo He. Effective and efficient content redundancy detection of Web videos. IEEE Transactions on Big Data, 7(1):187–198, March 2021. CODEN ???? ISSN 2332-7790.

Cheng:2022:MRE

[CLHN22] Harry Cheng, Lizi Liao, Linmei Hu, and Liqiang Nie. Multirelation extraction via a globallocal graph convolutional network. IEEE Transactions on Big Data, 8(6):1716–1728, December 2022. ISSN 2332-7790.

Cheng:2023:SWS

[CLL23] Chien-Tse Cheng, Yu-Hsun Lin, and Chung-Shou Liao. Seq2CASE: Weakly supervised sequence to commentary aspect score estimation for recommendation. IEEE Transactions on Big Data, 9(6):1670–1682, December 2023. ISSN 2332-7790.

Celikten:2017:MUB

[ÇLM17] E. Çelikten, G. Le Falher, and M. Mathioudakis. Modeling urban behavior by mining geotagged social data. IEEE Transactions on Big Data, 3(2):220– 233, June 2017. CODEN ???? ISSN 2332-7790.

Chen:2022:STI

[CLW^{+}22] Chao Chen, Qiang Liu, Xingchen Wang, Chengwu Liao, and Daqing Zhang. semi-Traj2Graph identifying fine-grained driving style with GPS trajectory data via multi-task learning. IEEE Transactions on Big Data, 8 (6):1550–1565, December 2022. ISSN 2332-7790.

Chen:2022:HPS

[CLZ22] Jia Chen, Xin Luo, and MengChu Zhou. Hierarchical particle swarm optimizationincorporated latent factor analysis for large-scale incomplete matrices. IEEE Transactions on Big Data, 8(6):1524–1536, December 2022. ISSN 2332-7790.

Cao:2021:CES

[CM21] Zhengjun Cao and Olivier Markowitch. Comment on Efficient Secure Outsourcing of Large-Scale Sparse Linear Systems of Equations. IEEE Transactions on Big Data, 7(6):973– 974, December 2021. CODEN ???? ISSN 2332-7790. See [SLC^{+}18].

Cheng:2024:TES

[CMJ^{+}24] Yongli Cheng, Yan Ma, Hong Jiang, Lingfang Zeng, Fang Wang, Xianghao Xu, and Yuhang Wu. TgStore: an efficient storage system for large time-evolving graphs. IEEE Transactions on Big Data, 10 (2):158–173, April 2024. ISSN 2332-7790.

Chen:2021:CDP

[CMZL21] Jianwei Chen, Huadong Ma, Dong Zhao, and Liang Liu. Correlated differential privacy protection for mobile crowdsensing. IEEE Transactions on Big Data, 7(4):784–795, April 2021. CODEN ???? ISSN 2332-7790.

Chu:2018:MSP

[CQH^{+}18] L. Chu, R. Qiu, X. He, Z. Ling, and Y. Liu. Massive streaming PMU data modelling and analytics in smart grid state evaluation based on multiple high-dimensional covariance test. IEEE Transactions on Big Data, 4(1):55–64, March 2018. CODEN ???? ISSN 2332-7790.

Crago:2015:WIT

[Cra15] Stephen Crago. Welcome to the IEEE Transactions on Big Data. IEEE Transactions on Big Data, 1(1):1, March 2015. CODEN ???? ISSN 2332-7790.

                                        URL http://www.computer.
                                        org/csdl/trans/bd/2015/01/
                                        07265157.pdf.

Cui:2018:SBB

[CSL^{+}18] Y. Cui, J. Song, M. Li, Q. Ren, Y. Zhang, and X. Cai. SDNbased big data caching in ISP networks. IEEE Transactions on Big Data, 4(3):356–367, September 2018. CODEN ???? ISSN 2332-7790.

Cheung:2018:CUC

[CSW18] M. Cheung, J. She, and N. Wang. Characterizing user connections in social media through user-shared images. IEEE Transactions on Big Data, 4(4):447–458, December 2018. CODEN ???? ISSN 2332-7790.

Cui:2023:VAM

[CSW23] Wenqiang Cui, Girts Strazdins, and Hao Wang. Visual analysis of multidimensional big data: a scalable lightweight bundling method for parallel coordinates. IEEE Transactions on Big Data, 9(1):106– 117, February 2023. ISSN 2332-7790.

Chen:2021:DFL

[CSZ^{+}21] Min Chen, Xiaobo Shi, Yin Zhang, Di Wu, and Mohsen Guizani. Deep feature learning for medical image analysis with convolutional autoencoder neural network. IEEE Transactions on Big Data, 7(4):750– 758, April 2021. CODEN ???? ISSN 2332-7790.

Chen:2024:GSA

[CTC^{+}24] Rui Chen, Yongqiang Tang, Xiangrui Cai, Xiaojie Yuan, Wenlong Feng, and Wensheng Zhang. Graph structure aware contrastive multi-view clustering. IEEE Transactions on Big Data, 10(3):260–274, June 2024. ISSN 2332-7790.

Chen:2018:GEBa

[CW18a] J. Chen and H. Wang. Guest editorial: Big Data Infrastructure I. IEEE Transactions on Big Data, 4(2):148–149, June 2018. CODEN ???? ISSN 2332-7790.

Chen:2018:GEBb

[CW18b] J. Chen and H. Wang. Guest editorial: Big Data Infrastructure II. IEEE Transactions on Big Data, 4(3):299–300, September 2018. CODEN ???? ISSN 2332-7790.

Chang:2022:MRD

[CW22] Shih Yu Chang and Hsiao-Chun Wu. Multi-relational data characterization by tensors: Tensor inversion. IEEE Transactions on Big Data, 8(6):1650– 1663, December 2022. ISSN 2332-7790.

Chen:2023:IPK

[CWC^{+}23] Xue Chen, Cheng Wang, Jipeng Cui, Qing Yang, Teng Hu, and Changjun Jiang. Incorporating prior knowledge in local differentially private data collection for frequency estimation. IEEE Transactions on Big Data, 9 (2):499–511, April 2023. ISSN 2332-7790.

Chan:2017:DSL

[CWGA17] Y. Chan, A. Wellings, I. Gray, and N. Audsley. A distributed stream library for Java 8. IEEE Transactions on Big Data, 3(3): 262–275, September 2017. CO-DEN ???? ISSN 2332-7790.

Chen:2021:TDV

[CWL^{+}21] Fei Chen, Donghong Wang, Qiuzhen Lin, Jianyong Chen, Zhong Ming, Wei Yu, and Jing Qin. Towards dynamic verifiable pattern matching. IEEE Transactions on Big Data, 7(2):

421–435, June 2021. CODEN ???? ISSN 2332-7790.

Chen:2020:BDP

[CWPW20] Huangke Chen, Jinming Wen, Witold Pedrycz, and Guohua Wu. Big data processing workflows oriented realtime scheduling algorithm using task-duplication in geodistributed clouds. IEEE Transactions on Big Data, 6(1): 131–144, March 2020. CODEN ???? ISSN 2332-7790.

Cheng:2021:SNQ

[CWS^{+}21] Ke Cheng, Liangmin Wang, Yulong Shen, Hua Wang, Yongzhi Wang, Xiaohong Jiang, and Hong Zhong. Secure k-NN query on encrypted cloud data with multiple keys. IEEE Transactions on Big Data, 7(4): 689–702, April 2021. CODEN ???? ISSN 2332-7790.

Chen:2018:GAH

[CXT^{+}18] Z. Chen, J. Xu, J. Tang, K. A.

Kwiat, C. A. Kamhoua, and C. Wang. GPU-accelerated high-throughput online stream data processing. IEEE Transactions on Big Data, 4(2):191– 202, June 2018. CODEN ???? ISSN 2332-7790.

Chen:2023:HRG

[CXW^{+}23] Ziheng Chen, Tianyang Xu, Xiao-Jun Wu, Rui Wang, and Josef Kittler. Hybrid Riemannian graph-embedding metric learning for image set classification. IEEE Transactions on Big Data, 9(1):75–92, February 2023. ISSN 2332-7790.

Chen:2023:MLB

[CYC^{+}23] Yuanyi Chen, Peng Yu, Wenwang Chen, Zengwei Zheng, and Minyi Guo. Meta-learning based classification for moving object trajectories in mobile IoT. IEEE Transactions on Big Data, 9(2):584–596, April 2023. ISSN 2332-7790.

Chen:2022:CPF

[CYW^{+}22] Chao Chen, Sen Yang, Yasha Wang, Bin Guo, and Daqing Zhang. CrowdExpress: a probabilistic framework for on-time crowdsourced package deliveries. IEEE Transactions on Big Data, 8(3):827–842, June 2022. ISSN 2332-7790.

Chen:2023:API

[CYXS23] Hao Chen, Yu Ye, Ming Xiao, and Mikael Skoglund. Asynchronous parallel incremental block-coordinate descent for decentralized machine learning. IEEE Transactions on Big Data, 9(4):1252–1259, August 2023. ISSN 2332-7790.

Chen:2017:LCF

[CZ17a] K. Chen and Z. Zhang. Learning to classify fine-grained categories with privileged visualsemantic misalignment. IEEE Transactions on Big Data, 3(1): 37–43, March 2017. CODEN ???? ISSN 2332-7790.

Chopade:2017:FCD

[CZ17b] P. Chopade and J. Zhan. A framework for community detection in large networks using game-theoretic modeling. IEEE Transactions on Big Data, 3(3): 276–288, September 2017. CO-DEN ???? ISSN 2332-7790.

Cai:2019:MTU

[CZG^{+}19] Q. Cai, H. Zhang, W. Guo, G. Chen, B. C. Ooi, K. Tan, and W. Wong. MemepiC: Towards a unified in-memory big data management system. IEEE Transactions on Big Data, 5(1): 4–17, March 2019. CODEN ???? ISSN 2332-7790.

Chen:2017:ROP

[CZH^{+}17] Z. Chen, W. Zhang, B. Hu, X. Cao, S. Liu, and D. Meng. Retrieving objects by partitioning. IEEE Transactions on Big Data, 3(1):44–54, March 2017. CODEN ???? ISSN 2332-7790.

Cao:2019:TMD

[CZLL19] Bin Cao, Jianwei Zhao, Zhihan Lv, and Xin Liu. 3D terrain multiobjective deployment optimization of heterogeneous directional sensor networks in security monitoring. IEEE Transactions on Big Data, 5(4):495– 505, December 2019. CODEN ???? ISSN 2332-7790.

Chen:2021:CIC

[CZQ^{+}21] Huimin Chen, Zeyu Zhu, Fanchao Qi, Yining Ye, Zhiyuan Liu, Maosong Sun, and Jianbin Jin. Country image in COVID-19 pandemic: a case study of China. IEEE Transactions on Big Data, 7(1):81–92, March 2021. CODEN ???? ISSN 2332-7790.

Cao:2024:TRN

[CZZ^{+}24] Xingyu Cao, Xiangtao Zhang, Ce Zhu, Jiani Liu, and Yipeng Liu. TS-RTPM-Net: Datadriven tensor sketching for efficient CP decomposition. IEEE Transactions on Big Data, 10: 1–11, February 2024. ISSN 2332-7790.

Datta:2017:HHE

[DBAM17] S. Datta, P. Basuchowdhuri, S. Acharya, and S. Majumder. The habits of highly effective researchers: An empirical study. IEEE Transactions on Big Data, 3(1):3–17, March 2017. CODEN ???? ISSN 2332-7790.

Djenouri:2024:FCT

[DBMS24] Youcef Djenouri, Ahmed Nabil Belbachir, Tomasz Michalak, and Gautam Srivastava. A federated convolution transformer for fake news detection. IEEE Transactions on Big Data, 10 (3):214–225, June 2024. ISSN 2332-7790.

Dong:2021:DPO

[DDCHV21] Yucheng Dong, Zhaogang Ding, Francisco Chiclana, and Enrique Herrera-Viedma. Dynamics of public opinions in an online and offline social network. IEEE Transactions on Big Data, 7(4):610–618, April 2021. CODEN ???? ISSN 2332-7790.

Derbeko:2022:EPP

[DDGU22] Philip Derbeko, Shlomi Dolev, Ehud Gudes, and Jeffrey D. Ullman. Efficient and privacy preserving approximation of distributed statistical queries. IEEE Transactions on Big Data, 8(5):1399–1413, October 2022. ISSN 2332-7790.

Ding:2022:PER

[DEG^{+}22] Jiahao Ding, Sai Mounika Errapotu, Yuanxiong Guo, Haixia Zhang, Dongfeng Yuan, and Miao Pan. Private empirical risk minimization with analytic Gaussian mechanism for healthcare system. IEEE Transactions on Big Data, 8(4):1107– 1117, August 2022. ISSN 2332-7790.

DeCapitanidiVimercati:2023:SDD

[DFF^{+}23] Sabrina De Capitani di Vimercati, Dario Facchinetti, Sara Foresti, Giovanni Livraga, Gianluca Oldani, Stefano Paraboschi, Matthew Rossi, and Pierangela Samarati. Scalable distributed data anonymization for large datasets. IEEE Transactions on Big Data, 9(3):818– 831, June 2023. ISSN 2332-7790.

Dolev:2019:SGD

[DFG^{+}19] S. Dolev, P. Florissi, E. Gudes, S. Sharma, and I. Singer. A survey on geographically distributed big-data processing using MapReduce. IEEE Transactions on Big Data, 5(1):60– 80, March 2019. CODEN ???? ISSN 2332-7790.

Dong:2018:CAP

[DH18] D. Dong and J. Herbert. Content-aware partial compression for textual big data analysis in Hadoop. IEEE Transactions on Big Data, 4(4):459– 472, December 2018. CODEN ???? ISSN 2332-7790.

Ding:2021:MLB

[DHG21] Junhua Ding, Xin-Hua Hu, and Venkat Gudivada. A machine learning based framework for verification and validation of massive scale image data. IEEE Transactions on Big Data, 7(2): 451–467, June 2021. CODEN ???? ISSN 2332-7790.

Ding:2023:EIT

[DHQ^{+}23] Zhifei Ding, Jiahao Han, Rongtao Qian, Liming Shen, Siru Chen, Lingxin Yu, Yu Zhu, and Richen Liu. eBoF: Interactive temporal correlation analysis for ensemble data based on bag-of-features. IEEE Transactions on Big Data, 9(6):1726– 1737, December 2023. ISSN 2332-7790.

Ding:2017:DAU

[DLD^{+}17] Y. Ding, Y. Li, K. Deng, H. Tan, M. Yuan, and L. M. Ni. Detecting and analyzing urban regions with high impact of weather change on transport. IEEE Transactions on Big Data, 3(2):126–139, June 2017. CODEN ???? ISSN 2332-7790.

DeMarch:2020:IAP

[DLD^{+}20] R. De March, C. Leuzzi, M. Deffacis, F. Caronte, A. F. Mulone, and R. Messineo. Innovative approach for PMM data processing and analytics. IEEE Transactions on Big Data, 6(3):452– 459, September 2020. CODEN ???? ISSN 2332-7790.

Dong:2016:HDP

[DLL^{+}16] L. Dong, Z. Lin, Y. Liang, L. He, N. Zhang, Q. Chen, X. Cao, and E. Izquierdo. A hierarchical distributed processing framework for big image data. IEEE Transactions on Big Data, 2(4):297–309, December 2016. CODEN ???? ISSN 2332-7790.

Datcu:2020:SIB

[DLL^{+}20] Mihai Datcu, Jacqueline Le Moigne, Sveinung Loekken, Pierre Soille, and Gui-Song Xia. Special issue on Big Data from space. IEEE Transactions on Big Data, 6(3):427– 429, September 2020. CODEN ???? ISSN 2332-7790.

Dai:2024:LLG

[DLL^{+}24] Dawei Dai, Yingge Liu, Yutang Li, Shiyu Fu, Shuyin Xia, and Guoyin Wang. LGRL: Localglobal representation learning for on-the-fly FG-SBIR. IEEE Transactions on Big Data, 10 (4):543–555, August 2024. ISSN 2332-7790.

Ding:2019:SBT

[DLZ^{+}19] Yichen Ding, Yanhua Li, Xun Zhou, Zhuojie Huang, Simin You, and Jun Luo. Sampling big trajectory data for traversal trajectory aggregate query. IEEE Transactions on Big Data, 5(4): 550–563, December 2019. CO-DEN ???? ISSN 2332-7790.

Dai:2018:CIR

[DQWQ18] W. Dai, L. Qiu, A. Wu, and M. Qiu. Cloud infrastructure resource allocation for big data applications. IEEE Transactions on Big Data, 4(3):313– 324, September 2018. CODEN ???? ISSN 2332-7790.

Datta:2016:HLW

[DSS16] Subhajit Datta, Santonu Sarkar, and A. S. M. Sajeev. How long will this live? Discovering the lifespans of software engineering ideas. IEEE Transactions on Big Data, 2(2):124–137, June 2016. CODEN ???? ISSN 2332-7790.

dosSantos:2021:HDB

[dSSN^{+}21] Joelson Antônio dos Santos, Talat Iqbal Syed, Murilo C. Naldi, Ricardo J. G. B. Campello, and Joerg Sander. Hierarchical density-based clustering using MapReduce. IEEE Transactions on Big Data, 7(1):102– 114, March 2021. CODEN ???? ISSN 2332-7790.

Ding:2019:ERI

[DWSJ19] X. Ding, L. Wang, Z. Shao, and H. Jin. Efficient recommendation of de-identification policies using MapReduce. IEEE Transactions on Big Data, 5(3):343– 354, September 2019. CODEN ???? ISSN 2332-7790.

Du:2020:DPP

[DWXZ20] Miao Du, Kun Wang, Zhuoqun Xia, and Yan Zhang. Differential privacy preserving of training model in wireless Big Data with edge computing. IEEE Transactions on Big Data, 6(2): 283–295, June 2020. CODEN ???? ISSN 2332-7790.

Dong:2021:GEA

[DZ21] Yuxiao Dong and Marinka Zitnik. Guest editorial: AI for COVID-19. IEEE Transactions on Big Data, 7(1):1–2, March 2021. CODEN ???? ISSN 2332-7790.

Dai:2020:TRT

[DZPM20] Wei Dai, Qing Zhang, Weike Pan, and Zhong Ming. Transfer to rank for top-N recommendation. IEEE Transactions on Big Data, 6(4):770–779, December 2020. CODEN ???? ISSN 2332-7790.

DaSilva:2020:TSM

[DZS20] João Paulo Da Silva, Jurandir Zullo, and Luciana Alvim Santos Romani. A time series mining approach for agricultural area detection. IEEE Transactions on Big Data, 6(3):537– 546, September 2020. CODEN ???? ISSN 2332-7790.

Edstrom:2019:DPE

[ECG^{+}19] J. Edstrom, D. Chen, Y. Gong, J. Wang, and N. Gong. Datapattern enabled self-recovery low-power storage system for big video data. IEEE Transactions on Big Data, 5(1):95– 105, March 2019. CODEN ???? ISSN 2332-7790.

Ebadi:2022:MNI

[EJCR22] Nima Ebadi, Mohsen Jozani, Kim-Kwang Raymond Choo, and Paul Rad. A memory network information retrieval model for identification of news misinformation. IEEE Transactions on Big Data, 8(5):1358– 1370, October 2022. ISSN 2332-7790.

Etemadi:2022:PPE

[EL22] Roohollah Etemadi and Jianguo Lu. PES: Priority edge sampling in streaming triangle estimation. IEEE Transactions on Big Data, 8(2):470– 481, February 2022. CODEN ???? ISSN 2332-7790.

Efthymiou:2020:BBA

[ESC20] Vasilis Efthymiou, Kostas Stefanidis, and Vassilis Christophides. Benchmarking blocking algorithms for Web entities. IEEE Transactions on Big Data, 6(2): 382–395, June 2020. CODEN ???? ISSN 2332-7790.

Farokhi:2021:PPP

[Far21]  Farhad Farokhi.  Privacy-

preserving public release of datasets for support vector machine classification. IEEE Transactions on Big Data, 7(5): 893–899, November 2021. CO-DEN ???? ISSN 2332-7790.

Farokhi:2023:NPD

[Far23] Farhad Farokhi. Noiseless privacy: Definition, guarantees, and applications. IEEE Transactions on Big Data, 9(1):51– 62, February 2023. ISSN 2332-7790.

Fei:2021:VAP

[FLYL21] Xiongwei Fei, Kenli Li, Wangdong Yang, and Keqin Li. Velocity-aware parallel encryption algorithm with low energy consumption for streams. IEEE Transactions on Big Data, 7(4): 619–631, April 2021. CODEN ???? ISSN 2332-7790.

Francesco:2018:AUC

[FMD18] P. D. Francesco, F. Malandrino, and L. A. DaSilva. Assembling and using a cellular dataset for mobile network analysis and planning. IEEE Transactions on Big Data, 4(4):614–620, December 2018. CODEN ???? ISSN 2332-7790.

Fei:2022:OHD

[FRS^{+}22] Zhe Fei, Yevgen Ryeznik, Oleksandr Sverdlov, Chee Wei Tan, and Weng Kee Wong. An overview of healthcare data analytics with applications to the COVID-19 pandemic. IEEE Transactions on Big Data, 8 (6):1463–1480, December 2022. ISSN 2332-7790.

Fu:2016:PTS

[Fu16] Yun Fu. Personalized travel sequence recommendation on multi-source big social media. IEEE Transactions on Big Data, 2(1):43–56, March 2016. CODEN ???? ISSN 2332-7790.

Fu:2022:FFL

[FWHM22] Weijie Fu, Meng Wang, Shijie Hao, and Tingting Mu. FLAG: Faster learning on anchor graph with label predictor optimization. IEEE Transactions on Big Data, 8(3):579–591, June 2022. ISSN 2332-7790.

Fang:2023:HRE

[FWS^{+}23] Peng Fang, Fang Wang, Zhan Shi, Hong Jiang, Dan Feng, Xianghao Xu, and Wei Yin. How to realize efficient and scalable graph embeddings via an entropy-driven mechanism. IEEE Transactions on Big Data, 9(1):358–371, February 2023. ISSN 2332-7790.

Feng:2023:SAD

[FWWC23] Yujie Feng, Jiangtao Wang, Yasha Wang, and Xu Chu. Spatial-attention and demographicaugmented generative adversarial imputation network for population health data reconstruction. IEEE Transactions on Big Data, 9(4):1057–1070, August 2023. ISSN 2332-7790.

Fahy:2022:FTM

[FY22] Conor Fahy and Shengxiang Yang. Finding and tracking multi-density clusters in online dynamic data streams. IEEE Transactions on Big Data, 8(1): 178–192, February 2022. CO-DEN ???? ISSN 2332-7790.

Feng:2019:SHO

[FYD^{+}19] J. Feng, L. T. Yang, G. Dai, W. Wang, and D. Zou. A secure high-order Lanczos-based orthogonal tensor SVD for big data reduction in cloud environment. IEEE Transactions on Big Data, 5(3):355– 367, September 2019. CODEN ???? ISSN 2332-7790.

Feng:2021:ISH

[FYD^{+}21] Jun Feng, Laurence T. Yang, Guohui Dai, Jinjun Chen, and Zheng Yan. An improved secure high-order-Lanczos based orthogonal tensor SVD for outsourced cyber-physical-social big data reduction. IEEE Transactions on Big Data, 7(4): 808–818, April 2021. CODEN ???? ISSN 2332-7790.

Fu:2022:NNP

[FYZ^{+}22] Anmin Fu, Shui Yu, Yuqing Zhang, Huaqun Wang, and Chanying Huang. NPP: a new privacy-aware public auditing scheme for cloud data sharing with group users. IEEE Transactions on Big Data, 8(1):14– 24, February 2022. CODEN ???? ISSN 2332-7790.

Fang:2018:EVM

[FZ18] H. Fang and Z. Zhang. An enhanced visualization method to aid behavioral trajectory pattern recognition infrastructure for big longitudinal data. IEEE Transactions on Big Data, 4(2): 289–298, June 2018. CODEN ???? ISSN 2332-7790.

Fan:2020:MCS

[FZC20] Jicong Fan, Mingbo Zhao, and Tommy W. S. Chow. Matrix completion via sparse factorization solved by accelerated proximal alternating linearized minimization. IEEE Transactions on Big Data, 6(1):119– 130, March 2020. CODEN ???? ISSN 2332-7790.

Gupta:2023:MST

[GB23] Vinayak Gupta and Srikanta Bedathur. Modeling spatial trajectories using coarse-grained smartphone logs. IEEE Transactions on Big Data, 9(2):608– 620, April 2023. ISSN 2332-7790.

Ghosh:2020:CPI

[GCA^{+}20] Debopriya Ghosh, Javier Cabrera, Tarek N. Adam, Petros Levounis, and Nabil R. Adam. Comorbidity patterns and its impact on health outcomes: Two-way clustering analysis. IEEE Transactions on Big Data, 6(2):359–368, June 2020. CODEN ???? ISSN 2332-7790.

Guo:2023:CDB

[GCL^{+}23] Kun Guo, Zhanhong Chen, Xu Lin, Ling Wu, Zhi-Hui Zhan, Yuzhong Chen, and Wenzhong Guo. Community detection based on multiobjective particle swarm optimization and graph attention variational autoencoder. IEEE Transactions on Big Data, 9 (2):569–583, April 2023. ISSN 2332-7790.

Gong:2024:SEN

[GDY24] Chaoyu Gong, Jim Demmel, and Yang You. Scalable evidential K-nearest neighbor classification on big data. IEEE Transactions on Big Data, 10 (3):226–237, June 2024. ISSN 2332-7790.

Giles:2016:ASE

[Gil16]  C. Lee Giles.  Algorithm-

Seer: A system for extracting and searching for algorithms in scholarly big data. IEEE Transactions on Big Data, 2(1):3– 17, March 2016. CODEN ???? ISSN 2332-7790.

Goel:2017:OPP

[GKR17] P. Goel, L. Kulik, and K. Ramamohanarao. Optimal pick up point selection for effective ride sharing. IEEE Transactions on Big Data, 3(2):154–168, June 2017. CODEN ???? ISSN 2332-7790.

Guan:2024:EPP

[GLZ^{+}24] Yunguo Guan, Rongxing Lu, Songnian Zhang, Yandong Zheng, Jun Shao, and Guiyi Wei. Efficient and privacypreserving aggregate query over public property graphs. IEEE Transactions on Big Data, 10 (2):146–157, April 2024. ISSN 2332-7790.

Gogoglou:2020:DDU

[GM20] Antonia Gogoglou and Yannis Manolopoulos. A data-driven unified framework for predicting citation dynamics. IEEE Transactions on Big Data, 6(4): 727–740, December 2020. CO-DEN ???? ISSN 2332-7790.

Gao:2022:MVC

[GMZ^{+}22] Xiaowei Gao, Xiaoke Ma, Wensheng Zhang, Jianbin Huang, He Li, Yanni Li, and Jiangtao Cui. Multi-view clustering with self-representation and structural constraint. IEEE Transactions on Big Data, 8(4): 882–893, August 2022. ISSN 2332-7790.

Gulisano:2021:SDD

[GNPT21] Vincenzo Gulisano, Yiannis Nikolakopoulos, Marina Papatriantafilou, and Philippas Tsigas. ScaleJoin: A deterministic, disjoint-parallel and skewresilient stream join. IEEE Transactions on Big Data, 7(2): 299–312, June 2021. CODEN ???? ISSN 2332-7790.

Goudarzi:2019:HAB

[Gou19] M. Goudarzi. Heterogeneous architectures for big data batch processing in MapReduce paradigm. IEEE Transactions on Big Data, 5(1):18–33, March 2019. CODEN ???? ISSN 2332-7790.

Ge:2023:CDT

[GQP^{+}23] Ningchao Ge, Zheng Qin, Peng Peng, Mingdao Li, Lei Zou, and Keqin Li. A cost-driven top-K queries optimization approach on federated RDF systems. IEEE Transactions on Big Data, 9(2):665–676, April 2023. ISSN 2332-7790.

Gai:2021:PPD

[GQZ21] Keke Gai, Meikang Qiu, and Hui Zhao. Privacy-preserving data encryption strategy for big data in mobile cloud computing. IEEE Transactions on Big Data, 7(4):678–688, April 2021. CODEN ???? ISSN 2332-7790.

Groh:2023:GGB

[GRWL23] Fabian Groh, Lukas Ruppert, Patrick Wieschollek, and Hendrik P. A. Lensch. GGNN: Graph-based GPU nearest neighbor search. IEEE Transactions on Big Data, 9(1):267– 279, February 2023. ISSN 2332-7790.

Gyeera:2023:RAP

[GSS23] Thomas Weripuo Gyeera, Anthony J. H. Simons, and Mike Stannett. Regression analysis of predictions and forecasts of cloud data center KPIs using the boosted decision tree algorithm. IEEE Transactions on Big Data, 9(4):1071–1085, August 2023. ISSN 2332-7790.

Guo:2024:LBB

[GWLL24] Lu Guo, Limin Wang, Qilong Li, and Kuo Li. Learning balanced Bayesian classifiers from labeled and unlabeled data. IEEE Transactions on Big Data, 10(4):330–342, August 2024. ISSN 2332-7790.

Gu:2020:LUB

[GXQ^{+}20] Ke Gu, Xin Xu, Junfei Qiao, Qiuping Jiang, Weisi Lin, and Daniel Thalmann. Learning a unified blind image quality metric via on-line and off-line big training instances. IEEE Transactions on Big Data, 6(4):780– 791, December 2020. CODEN ???? ISSN 2332-7790.

Gu:2020:SWC

[GZL^{+}20] Yu Gu, Yifan Zhang, Jie Li, Yusheng Ji, Xin An, and Fuji Ren. Sleepy: Wireless channel data driven sleep monitoring via commodity WiFi devices. IEEE Transactions on Big Data, 6(2): 258–268, June 2020. CODEN ???? ISSN 2332-7790.

Gao:2023:DDG

[GZS23] Shuai Gao, Zhongbao Zhang, and Sen Su. DAWN: Domain generalization based network alignment. IEEE Transactions on Big Data, 9(3):878– 888, June 2023. ISSN 2332-7790.

Guo:2023:OEB

[GZW23] Jianxiong Guo, Yapu Zhang, and Weili Wu. An overall evaluation on benefits of competitive influence diffusion. IEEE Transactions on Big Data, 9 (2):653–664, April 2023. ISSN 2332-7790.

Hochbaum:2016:SCL

[HB16] Dorit S. Hochbaum and Philipp Baumann. Sparse computation for large-scale data mining. IEEE Transactions on Big Data, 2(2):151–174, June 2016. CODEN ???? ISSN 2332-7790.

Hu:2017:TTC

[HBLK17] T. Hu, E. Bigelow, J. Luo, and H. Kautz. Tales of two cities: Using social media to understand idiosyncratic lifestyles in distinctive metropolitan areas. IEEE Transactions on Big Data, 3(1):55–66, March 2017. CODEN ???? ISSN 2332-7790.

Hou:2022:FMV

[HCS^{+}22] Dongdong Hou, Yang Cong, Gan Sun, Jiahua Dong, Jun Li, and Kai Li. Fast multi-view outlier detection via Deep Encoder. IEEE Transactions on Big Data, 8(4):1047–1058, August 2022. ISSN 2332-7790.

He:2019:GEI

[HCZ19] B. He, Y. Chen, and J. Zhou.

Guest Editors introduction: Special issue on big data systems on emerging architectures. IEEE Transactions on Big Data, 5(1):2–3, March 2019. CODEN ???? ISSN 2332-7790.

Hu:2024:GCC

[HCZ^{+}24] Xinxin Hu, Haotian Chen, Junjie Zhang, Hongchang Chen, Shuxin Liu, Xing Li, Yahui Wang, and Xiangyang Xue. GAT-COBO: Cost-sensitive graph neural network for telecom fraud detection. IEEE Transactions on Big Data, 10 (4):528–542, August 2024. ISSN 2332-7790.

Hajeer:2019:HBD

[HD19] M. Hajeer and D. Dasgupta. Handling big data using a dataaware HDFS and evolutionary clustering technique. IEEE Transactions on Big Data, 5(2): 134–147, June 2019. CODEN ???? ISSN 2332-7790.

Hu:2020:EPM

[HFX^{+}20] Die Hu, Dan Feng, Yulai Xie, Gongming Xu, Xinrui Gu, and Darrell Long. Efficient provenance management via clustering and hybrid storage in Big Data environments. IEEE Transactions on Big Data, 6(4): 792–803, December 2020. CO-DEN ???? ISSN 2332-7790.

Huang:2023:TPA

[HGY^{+}23] Jianli Huang, Xianjie Guo, Kui Yu, Fuyuan Cao, and Jiye Liang. Towards privacy-aware causal structure learning in federated setting. IEEE Transactions on Big Data, 9(6):1525– 1535, December 2023. ISSN 2332-7790.

Huang:2018:NFG

[HH18] X. Huang and C. Huang. NGD: Filtering graphs for visual analysis. IEEE Transactions on Big Data, 4(3):381–395, September 2018. CODEN ???? ISSN 2332-7790.

Haider:2019:DAB

[HHX^{+}19] W. Haider, J. Hu, Y. Xie, X. Yu, and Q. Wu. Detecting anomalous behavior in cloud servers by nested-arc hidden semi-Markov model with state summarization. IEEE Transactions on Big Data, 5(3):305– 316, September 2019. CODEN ???? ISSN 2332-7790.

Han:2024:MDD

[HHZ^{+}24] Zhi-Long Han, Ting-Zhu Huang, Xi-Le Zhao, Hao Zhang, and Yun-Yang Liu. Multidimensional data recovery via feature-based fully-connected tensor network decomposition. IEEE Transactions on Big Data, 10(4):386–399, August 2024. ISSN 2332-7790.

Hamandawana:2022:AMD

[HKKC22] Prince Hamandawana, Awais Khan, Jongik Kim, and Tae-Sun Chung. Accelerating ML/ DL applications with hierarchical caching on deduplication storage clusters. IEEE Transactions on Big Data, 8(6):1622– 1636, December 2022. ISSN 2332-7790.

Hu:2018:SVA

[HLC^{+}18] C. Hu, W. Li, X. Cheng, J. Yu, S. Wang, and R. Bie. A secure and verifiable access control scheme for big data storage in clouds. IEEE Transactions on Big Data, 4(3):341– 355, September 2018. CODEN ???? ISSN 2332-7790.

He:2022:JCD

[HLF^{+}22] Dongxiao He, Huixin Liu, Zhiyong Feng, Xiaobao Wang, Di Jin, Wenze Song, and Yuxiao Huang. A joint community detection model: Integrating directed and undirected probabilistic graphical models via factor graph with attention mechanism. IEEE Transactions on Big Data, 8(4):994–1006, August 2022. ISSN 2332-7790.

Hu:2020:BDC

[HLH^{+}20] Shiyan Hu, Xin Li, Haibo He, Shuguang Cui, and Manish Parashar. Big data for cyberphysical systems. IEEE Transactions on Big Data, 6(4):606– 608, December 2020. CODEN ???? ISSN 2332-7790.

Han:2023:CMP

[HLL^{+}23a] Feng Han, Chengcai Leng, Bing Li, Anup Basu, and Licheng Jiao. Cosine multilinear principal component analysis for recognition. IEEE Transactions on Big Data, 9(6):1620–1630, December 2023. ISSN 2332-7790.

Huang:2023:FCD

[HLL^{+}23b] Wei Huang, Jia Liu, Tianrui Li, Shenggong Ji, Dexian Wang, and Tianqiang Huang. Fed-CKE: Cross-domain knowledge graph embedding in federated learning. IEEE Transactions on Big Data, 9(3):792–804, June 2023. ISSN 2332-7790.

Han:2022:DSB

[HLLZ22a] Yang Han, Jacqueline C. K. Lam, Victor O. K. Li, and Qi Zhang. A domain-specific Bayesian deep-learning approach for air pollution forecast. IEEE Transactions on Big Data, 8(4):1034–1046, August 2022. ISSN 2332-7790.

Huang:2022:DCD

[HLLZ22b] Weiyi Huang, Peng Li, Bo Li, and Tao Zhang. DMP: Content delivery with dynamic movement pattern in vehicular networks. IEEE Transactions on Big Data, 8(5):1371–1386, October 2022. ISSN 2332-7790.

Huang:2023:TLD

[HLW^{+}23] Xiaosai Huang, Jing Li, Jia Wu, Jun Chang, and Donghua Liu. Transfer learning with document-level data augmentation for aspect-level sentiment classification. IEEE Transactions on Big Data, 9(6):1643– 1657, December 2023. ISSN 2332-7790.

Huang:2022:TSA

[HMW^{+}22] Chengqiang Huang, Geyong Min, Yulei Wu, Yiming Ying, Ke Pei, and Zuochang Xiang. Time series anomaly detection for trustworthy services in cloud computing systems. IEEE Transactions on Big Data, 8(1): 60–72, February 2022. CODEN ???? ISSN 2332-7790.

Hosseini:2017:ODL

[HPESZ17] M. Hosseini, D. Pompili, K. Elisevich, and H. Soltanian-Zadeh. Optimized deep learning for EEG big data and seizure prediction BCI via Internet of Things. IEEE Transactions on Big Data, 3(4):392–404, December 2017. CODEN ???? ISSN 2332-7790.

Hildebrandt:2020:LSD

[HPWR20] Kai Hildebrandt, Fabian Panse, Niklas Wilcke, and Norbert Ritter. Large-scale data pollution with Apache Spark. IEEE Transactions on Big Data, 6(2): 396–411, June 2020. CODEN ???? ISSN 2332-7790.

Hallaji:2024:DFL

[HRFS^{+}24] Ehsan Hallaji, Roozbeh Razavi-Far, Mehrdad Saif, Boyu Wang, and Qiang Yang. Decentralized federated learning: a survey on security and privacy. IEEE Transactions on Big Data, 10 (2):194–213, April 2024. ISSN 2332-7790.

Han:2020:CWS

[HS20] Guyue Han and Harish Sethu. Closed walk sampler: an efficient method for estimating eigenvalues of large graphs. IEEE Transactions on Big Data, 6(1):29–42, March 2020. CODEN ???? ISSN 2332-7790.

Hu:2020:MPF

[HSEY20] R. Lily Hu, Ryan Skorupski, Robert Entriken, and Yinyu Ye. A mathematical programming formulation for optimal load shifting of electricity demand for the smart grid. IEEE Transactions on Big Data, 6(4): 638–651, December 2020. CO-DEN ???? ISSN 2332-7790.

He:2022:NGS

[HSX^{+}22] Liang He, Bin Shao, Yanghua Xiao, Yatao Li, Tie-Yan Liu, Enhong Chen, and Huanhuan Xia. Neurally-guided semantic navigation in knowledge graph. IEEE Transactions on Big Data, 8(3):607–615, June 2022. ISSN 2332-7790.

Hua:2015:LVS

[Hua15] Xian-Sheng Hua. Learning visual semantic relationships for efficient visual retrieval. IEEE Transactions on Big Data, 1(4): 152–161, December 2015. CO-DEN ???? ISSN 2332-7790.

Ho:2021:AAI

[HVVP21] Nguyen Ho, Huy Vo, Mai Vu, and Torben Bach Pedersen. AMIC: an adaptive information theoretic method to identify multi-scale temporal correlations in big time series data. IEEE Transactions on Big Data, 7(1):128–146, March 2021. CODEN ???? ISSN 2332-7790.

Huang:2020:BDE

[HWB^{+}20] Jie Huang, Cheng-Xiang Wang, Lu Bai, Jian Sun, Yang Yang, Jie Li, Olav Tirkkonen, and Ming-Tuo Zhou. A Big Data enabled channel model for 5G wireless communication systems. IEEE Transactions on Big Data, 6(2):211–222, June 2020. CODEN ???? ISSN 2332-7790.

Huang:2020:SUA

[HWC20] Chao Huang, Dong Wang, and Nitesh V. Chawla. Scalable uncertainty-aware truth discovery in Big Data social sensing applications for cyber-physical systems. IEEE Transactions on Big Data, 6(4):702–713, December 2020. CODEN ???? ISSN 2332-7790.

Huang:2020:PSA

[HWTM20] Chao Huang, Dong Wang, Jun Tao, and Brian Mann. On physical-social-aware localness inference by exploring Big Data from location-based services. IEEE Transactions on Big Data, 6(4):679–690, December 2020. CODEN ???? ISSN 2332-7790.

He:2022:ICM

[HWW^{+}22] Dongxiao He, Yanli Wu, Youyou Wang, Zhizhi Yu, Zhiyong Feng, Xiaobao Wang, and Yuxiao Huang. Identification of communities with multisemantics via Bayesian generative model. IEEE Transactions on Big Data, 8(4):869–881, August 2022. ISSN 2332-7790.

Huang:2022:MNT

[HX22] Jin Huang and Ming Xiao. Mobile network traffic prediction based on seasonal adjacent windows sampling and conditional probability estimation. IEEE Transactions on Big Data, 8(5): 1155–1168, October 2022. ISSN 2332-7790.

Hao:2023:ESD

[HXW^{+}23] Kun Hao, Junchang Xin, Zhiqiong Wang, Zhongming Yao, and Guoren Wang. Efficient and secure data sharing scheme on interoperable blockchain database. IEEE Transactions on Big Data, 9(4): 1171–1185, August 2023. ISSN 2332-7790.

Hu:2020:MDS

[HXYZ20] Fan Hu, Gui-Song Xia, Wen Yang, and Liangpei Zhang. Mining deep semantic representations for scene classification of high-resolution remote sensing imagery. IEEE Transactions on Big Data, 6(3):522– 536, September 2020. CODEN ???? ISSN 2332-7790.

Hu:2022:AII

[HYLZ22] Lun Hu, Shicheng Yang, Xin Luo, and MengChu Zhou. An algorithm of inductively identifying clusters from attributed graphs. IEEE Transactions on Big Data, 8(2):523–534, February 2022. CODEN ???? ISSN 2332-7790.

Huang:2022:WSL

[HZ22] Aiping Huang and Tiesong Zhao. Weak supervision learning for object co-segmentation. IEEE Transactions on Big Data, 8(4):1129–1140, August 2022. ISSN 2332-7790.

He:2022:BNM

[HZF^{+}22] Chaobo He, Yulong Zheng, Xiang Fei, Hanchao Li, Zeng Hu, and Yong Tang. Boosting nonnegative matrix factorization based community detection with graph attention autoencoder. IEEE Transactions on Big Data, 8(4):968–981, August 2022. ISSN 2332-7790.

Hu:2021:RSC

[HZY^{+}21] Xiao Hu, Hui Zhang, Chun-Ming Yang, XuJian Zhao, and Bo Li. Regularized spectral clustering with entropy perturbation. IEEE Transactions on Big Data, 7(6):967–972, December 2021. CODEN ???? ISSN 2332-7790.

Huang:2021:LSS

[HZZ^{+}21] Heqing Huang, Cong Zheng, Junyuan Zeng, Wu Zhou, Sencun Zhu, Peng Liu, Ian Molloy, Suresh Chari, Ce Zhang, and Quanlong Guan. A largescale study of Android malware development phenomenon on public malware submission and scanning platform. IEEE Transactions on Big Data, 7(2): 255–270, June 2021. CODEN ???? ISSN 2332-7790.

Ibrahim:2020:BCS

[ICZ20] Mohamed Ibrahim, Krishnendu Chakrabarty, and Jun Zeng. BioCyBig: a cyberphysical system for integrative microfluidics-driven analysis of genomic association studies. IEEE Transactions on Big Data, 6(4):609–623, December 2020. CODEN ???? ISSN 2332-7790.

Iscen:2018:MVS

[IFG^{+}18] A. Iscen, T. Furon, V. Gripon, M. Rabbat, and H. Jégou. Memory vectors for similarity search in high-dimensional spaces. IEEE Transactions on Big Data, 4(1):65–77, March 2018. CODEN ???? ISSN 2332-7790.

Iglesias:2019:PDI

[IZ19] Félix Iglesias and Tanja Zseby.

Pattern discovery in internet background radiation. IEEE Transactions on Big Data, 5(4): 467–480, December 2019. CO-DEN ???? ISSN 2332-7790.

Jiang:2022:MGB

[JDC^{+}22] Nan Jiang, Fuxian Duan, Honglong Chen, Wei Huang, and Ximeng Liu. MAFI: GNNbased multiple aggregators and feature interactions network for fraud detection over heterogeneous graph. IEEE Transactions on Big Data, 8(4):905– 919, August 2022. ISSN 2332-7790.

Johnson:2021:BSS

[JDJ21] Jeff Johnson, Matthijs Douze, and Hervé Jégou. Billionscale similarity search with GPUs. IEEE Transactions on Big Data, 7(3):535–547, July 2021. CODEN ???? ISSN 2332-7790.

Jiang:2017:ABH

[JFG17] S. Jiang, J. Ferreira, and M. C.

Gonzalez. Activity-based human mobility patterns inferred from mobile phone data: A case study of Singapore. IEEE Transactions on Big Data, 3(2): 208–219, June 2017. CODEN ???? ISSN 2332-7790.

Jiang:2021:DDD

[JFZ^{+}21] Ruobing Jiang, Zhenni Feng, Desheng Zhang, Shuai Wang, Yanmin Zhu, Fan Zhang, and Tian He. Data-driven digital advertising with uncertain demand model in metro networks. IEEE Transactions on Big Data, 7(2):313–326, June 2021. CODEN ???? ISSN 2332-7790.

Jiao:2022:SRO

[JGP^{+}22] Pengfei Jiao, Xuan Guo, Ting Pan, Wang Zhang, Yulong Pei, and Lin Pan. A survey on role-oriented network embedding. IEEE Transactions on Big Data, 8(4):933–952, August 2022. ISSN 2332-7790.

Jia:2019:UPD

[JGS^{+}19] Z. Jia, W. Gao, Y. Shi, S. A. McKee, Z. Ji, J. Zhan, L. Wang, and L. Zhang. Understanding processors design decisions for data analytics in homogeneous data centers. IEEE Transactions on Big Data, 5(1):81–94, March 2019. CODEN ???? ISSN 2332-7790.

Jiang:2018:TMM

[JHT18] Y. Jiang, Z. Huang, and D. H. K. Tsang. Towards maxmin fair resource allocation for stream big data analytics in shared clouds. IEEE Transactions on Big Data, 4(1):130– 137, March 2018. CODEN ???? ISSN 2332-7790.

Jiang:2020:PPA

[JHT20] Yuxuan Jiang, Zhe Huang, and Danny H. K. Tsang. On powerpeak-aware scheduling for largescale shared clusters. IEEE Transactions on Big Data, 6(2): 412–426, June 2020. CODEN ???? ISSN 2332-7790.

Jiang:2019:DGC

[JKF19] S. Jiang, Y. Kong, and Y. Fu.

Deep geo-constrained autoencoder for non-landmark GPS estimation. IEEE Transactions on Big Data, 5(2):120– 133, June 2019. CODEN ???? ISSN 2332-7790.

Jiang:2023:PBB

[JLX^{+}23] Wenbo Jiang, Hongwei Li, Guowen Xu, Tianwei Zhang, and Rongxing Lu. Physical black-box adversarial attacks through transformations. IEEE Transactions on Big Data, 9(3): 964–974, June 2023. ISSN 2332-7790.

Jia:2020:WDF

[JPR^{+}20] Weijia Jia, Hongjian Peng, Na Ruan, Zhiqing Tang, and Wei Zhao. WiFind: Driver fatigue detection with finegrained Wi-Fi signal features. IEEE Transactions on Big Data, 6(2):269–282, June 2020. CODEN ???? ISSN 2332-7790.

Jayaraman:2022:FDN

[JRP^{+}22] Sethuraman Jayaraman, Manikandan Ramachandran, Rizwan Patan, Mahmoud Daneshmand, and Amir H. Gandomi. Fuzzy deep neural learning based on Goodman and Kruskal’s gamma for search engine optimization. IEEE Transactions on Big Data, 8(1):268– 277, February 2022. CODEN ???? ISSN 2332-7790.

Jiang:2023:TES

[JWLY23] Zhifeng Jiang, Wei Wang, Bo Li, and Qiang Yang. Towards efficient synchronous federated training: a survey on system optimization strategies. IEEE Transactions on Big Data, 9(2):437–454, April 2023. ISSN 2332-7790.

Jin:2022:GES

[JWS^{+}22] Di Jin, Wenjun Wang, Guojie Song, Philip S. Yu, and Jiawei Han. Guest editorial: Special issue on network structural modeling and learning in big data. IEEE Transactions on Big Data, 8(4):867–868, August 2022. ISSN 2332-7790.

Jazayeri:2022:FSM

[JY22] Ali Jazayeri and Christopher C. Yang. Frequent subgraph mining algorithms in static and temporal graph-transaction settings: a survey. IEEE Transactions on Big Data, 8(6):1443– 1462, December 2022. ISSN 2332-7790.

Joneidi:2020:OSS

[JZS^{+}20] Mohsen Joneidi, Alireza Zaeemzadeh, Behzad Shahrasbi, Guo-Jun Qi, and Nazanin Rahnavard. E-optimal sensor selection for compressive sensingbased purposes. IEEE Transactions on Big Data, 6(1):51– 65, March 2020. CODEN ???? ISSN 2332-7790.

Jin:2022:ICW

[JZY22] Hai Jin, Zhaobo Zhang, and Pingpeng Yuan. Improving Chinese word representation using four corners features. IEEE Transactions on Big Data, 8(4): 982–993, August 2022. ISSN 2332-7790.

Koehler:2021:IDC

[KAB^{+}21] Martin Koehler, Edward Abel, Alex Bogatu, Cristina Civili, Lacramioara Mazilu, Nikolaos Konstantinou, Alvaro A. A. Fernandes, John Keane, Leonid Libkin, and Norman W. Paton. Incorporating data context to cost-effectively automate endto-end data wrangling. IEEE Transactions on Big Data, 7(1): 169–186, March 2021. CODEN ???? ISSN 2332-7790.

Khalajzadeh:2022:SAC

[KAG^{+}22] Hourieh Khalajzadeh, Mohamed Abdelrazek, John Grundy, John Hosking, and Qiang He. Survey and analysis of current end-user data analytics tool support. IEEE Transactions on Big Data, 8(1):152–165, February 2022. CODEN ???? ISSN 2332-7790.

Khalfi:2021:NMS

[KdRFA21] Besma Khalfi, Cyril de Runz, Sami Faiz, and Herman Akdag. A new methodology for storing consistent fuzzy geospatial data in big data environment. IEEE Transactions on Big Data, 7(2): 468–482, June 2021. CODEN ???? ISSN 2332-7790.

Kan:2023:CGF

[KFY^{+}23] Zhigang Kan, Linhui Feng, Zhangyue Yin, Linbo Qiao, Xipeng Qiu, and Dongsheng Li. A composable generative framework based on prompt learning for various information extraction tasks. IEEE Transactions on Big Data, 9(4):1238– 1251, August 2023. ISSN 2332-7790.

Kralevska:2018:HEC

[KGJØ18] K. Kralevska, D. Gligoroski, R. E. Jensen, and H. Øverby. HashTag erasure codes: From theory to practice. IEEE Transactions on Big Data, 4(4):516– 529, December 2018. CODEN ???? ISSN 2332-7790.

Kondor:2020:TMU

[KHdMR20] Dániel Kondor, Behrooz Hashemian, Yves-Alexandre de Montjoye, and Carlo Ratti. Towards matching user mobility traces in large-scale datasets. IEEE Transactions on Big Data, 6(4): 714–726, December 2020. CO-DEN ???? ISSN 2332-7790.

Kitsuregawa:2016:VEC

[Kit16] Masaru Kitsuregawa. Visual exploration of changes in passenger flows and tweets on megacity metro network. IEEE Transactions on Big Data, 2(1): 85–99, March 2016. CODEN ???? ISSN 2332-7790.

Kuang:2018:EPS

[KJC^{+}18] K. Kuang, M. Jiang, P. Cui, H. Luo, and S. Yang. Effective promotional strategies selection in social media: A datadriven approach. IEEE Transactions on Big Data, 4(4):487– 501, December 2018. CODEN ???? ISSN 2332-7790.

Kit:2016:EIG

[KKF16] D. Kit, Y. Kong, and Y. Fu. Efficient image geotagging using large databases. IEEE Transactions on Big Data, 2(4):325– 338, December 2016. CODEN ???? ISSN 2332-7790.

Khan:2016:UNR

[KKKK16] Farhan Khan, Dariush Kari, Ilyas Alper Karatepe, and Suleyman S. Kozat. Universal nonlinear regression on high dimensional data using adaptive hierarchical trees. IEEE Transactions on Big Data, 2(2):175– 188, June 2016. CODEN ???? ISSN 2332-7790.

Kotoulas:2015:SPL

[Kot15] Spyros Kotoulas. Scale-out processing of large RDF datasets. IEEE Transactions on Big Data, 1(4):138–150, December 2015. CODEN ???? ISSN 2332-7790.

Kejriwal:2022:KGS

[KS22] Mayank Kejriwal and Pedro Szekely. Knowledge graphs for social good: an entitycentric search engine for the human trafficking domain. IEEE Transactions on Big Data, 8(3): 592–606, June 2022. ISSN 2332-7790.

Koh:2022:MPC

[KZL^{+}22] Zann Koh, Yuren Zhou, Billy Pik Lik Lau, Chau Yuen, Bige Tunçer, and Keng Hua Chong. Multiple-perspective clustering of passive wi-fi sensing trajectory data. IEEE Transactions on Big Data, 8(5):1312–1325, October 2022. ISSN 2332-7790.

Kou:2023:WWT

[KZSW23] Ziyi Kou, Daniel Zhang, Lanyu Shang, and Dong Wang. What and why? Towards duo explainable fauxtography detection under constrained supervision. IEEE Transactions on Big Data, 9(1):133–146, February 2023. ISSN 2332-7790.

Li:2020:MEM

[LA20] Ruidong Li and Hitoshi Asaeda. MWBS: an efficient many-to-many wireless Big Data delivery scheme. IEEE Transactions on Big Data, 6(2):233–247, June 2020. CODEN ???? ISSN 2332-7790.

Liu:2019:POD

[LB19] X. Liu and R. Buyya. Performanceoriented deployment of streaming applications on cloud. IEEE Transactions on Big Data, 5(1): 46–59, March 2019. CODEN ???? ISSN 2332-7790.

Li:2018:MMI

[LBL^{+}18] Y. Li, J. Bao, Y. Li, Y. Wu, Z. Gong, and Y. Zheng. Mining the most influential klocation set from massive trajectories. IEEE Transactions on Big Data, 4(4):556–570, December 2018. CODEN ???? ISSN 2332-7790.

Li:2022:MMS

[LBSL22] Jianian Li, Peng Bao, Huawei Shen, and Xuanya Li. MiSTR: a multiview structural-temporal learning framework for rumor detection. IEEE Transactions on Big Data, 8(4):1007–1019, August 2022. ISSN 2332-7790.

Li:2019:SSB

[LBZ19] W. Li, Q. Bai, and M. Zhang.

SIMiner: A stigmergy-based model for mining influential nodes in dynamic social networks. IEEE Transactions on Big Data, 5(2):223–237, June 2019. CODEN ???? ISSN 2332-7790.

Liang:2023:CPM

[LCC^{+}23] Weichao Liang, Jie Cao, Lei Chen, Youquan Wang, Jia Wu, Amin Beheshti, and Jiangnan Tang. Crime prediction with missing data via spatiotemporal regularized tensor decomposition. IEEE Transactions on Big Data, 9(5):1392–1407, October 2023. ISSN 2332-7790.

Li:2018:AKS

[LCS18] Z. Li, H. Chandler, and H. Shen. Analysis of knowledge sharing activities on a social network incorporated discussion forum: A case study of DISboards. IEEE Transactions on Big Data, 4(4):432– 446, December 2018. CODEN ???? ISSN 2332-7790.

Lyu:2021:SSC

[LCWL21] Hanjia Lyu, Long Chen, Yu Wang, and Jiebo Luo. Sense and sensibility: Characterizing social media users regarding the use of controversial terms for COVID-19. IEEE Transactions on Big Data, 7(6):952–960, December 2021. CODEN ???? ISSN 2332-7790.

Li:2022:LBP

[LCWS22] Qi Li, Mingyu Cheng, Junfeng Wang, and Bowen Sun. LSTM based phishing detection for big email data. IEEE Transactions on Big Data, 8(1):278– 288, February 2022. CODEN ???? ISSN 2332-7790.

Li:2022:SIA

[LCY^{+}22] Angtai Li, Yu Chen, Zheng Yan, Xiaokang Zhou, and Shohei Shimizu. A survey on integrity auditing for data storage in the cloud: From single copy to multiple replicas. IEEE Transactions on Big Data, 8(5):1428– 1442, October 2022. ISSN 2332-7790.

Liu:2022:BBS

[LDY^{+}22] Gao Liu, Huidong Dong, Zheng Yan, Xiaokang Zhou, and Shohei Shimizu. B4SDC: a blockchain system for security data collection in MANETs. IEEE Transactions on Big Data, 8(3):739–752, June 2022. ISSN 2332-7790.

Li:2023:EEA

[LGW^{+}23] Qian Li, Shu Guo, Jia Wu, Jianxin Li, Jiawei Sheng, Hao Peng, and Lihong Wang. Event extraction by associating event types and argument roles. IEEE Transactions on Big Data, 9 (6):1549–1560, December 2023. ISSN 2332-7790.

Lao:2024:TSM

[LHWL24] Jinghuan Lao, Dong Huang, Chang-Dong Wang, and Jian-Huang Lai. Towards scalable multi-view clustering via joint learning of many bipartite graphs. IEEE Transactions on Big Data, 10(1):77–91, February 2024. ISSN 2332-7790.

Li:2022:TPM

[LHY^{+}22] He Li, Jianbin Huang, Hang Yuan, Jiangtao Cui, Xiaoke Ma, Shaojie Qiao, and Xindong Wu. A two-phase method to balance the result of distributed graph repartitioning. IEEE Transactions on Big Data, 8(6):1580– 1591, December 2022. ISSN 2332-7790.

Li:2015:RDS

[Li15] Xuelong Li. Robust discrete spectral hashing for large-scale image semantic indexing. IEEE Transactions on Big Data, 1(4): 162–171, December 2015. CO-DEN ???? ISSN 2332-7790.

Liu:2015:EIE

[Liu15] Huan Liu. Embracing information explosion without choking: Clustering and labeling in microblogging. IEEE Transactions on Big Data, 1(1):35– 46, March 2015. CODEN ???? ISSN 2332-7790.

Luo:2022:PSO

[LJC^{+}22] Changqing Luo, Jinlong Ji, Xuhui Chen, Ming Li, Laurence T. Yang, and Pan Li. Parallel secure outsourcing of largescale nonlinearly constrained nonlinear programming problems. IEEE Transactions on Big Data, 8(2):346–355, February 2022. CODEN ???? ISSN 2332-7790.

Liu:2024:MPF

[LLC^{+}24] Xin Liu, Jiuluan Lv, Feng Chen, Qingjie Wei, Hangxuan He, and Ying Qian. Multi-party federated recommendation based on semi-supervised learning. IEEE Transactions on Big Data, 10 (4):356–370, August 2024. ISSN 2332-7790.

Liu:2017:MDP

[LLCS17] A. Liu, Y. Lu, M. Chen, and Y. Su. Mitosis detection in phase contrast microscopy image sequences of stem cell populations: A critical review. IEEE Transactions on Big Data, 3(4): 443–457, December 2017. CO-DEN ???? ISSN 2332-7790.

Li:2020:EMK

[LLD^{+}20] Yaliang Li, Chaochun Liu, Nan Du, Wei Fan, Qi Li, Jing Gao, Chenwei Zhang, and Hao Wu. Extracting medical knowledge from crowdsourced question answering website. IEEE Transactions on Big Data, 6(2):309– 321, June 2020. CODEN ???? ISSN 2332-7790.

Lyu:2019:SUS

[LLF^{+}19] B. Lyu, Y. Li, J. Fu, A. C.

Trapp, H. Xie, and Y. Liao. Scalable user-substation assignment with big data from power grids. IEEE Transactions on Big Data, 5(2):209–222, June 2019. CODEN ???? ISSN 2332-7790.

Li:2023:DSD

[LLH^{+}23] Peipei Li, Yingying Liu, Yang Hu, Yuhong Zhang, Xuegang Hu, and Kui Yu. A driftsensitive distributed LSTM method for short text stream classification. IEEE Transactions on Big Data, 9(1):341– 357, February 2023. ISSN 2332-7790.

Liu:2022:FF

[LLL^{+}22] Yang Liu, Yingting Liu, Zhijie Liu, Yuxuan Liang, Chuishi Meng, Junbo Zhang, and Yu Zheng. Federated forest. IEEE Transactions on Big Data, 8(3):843–854, June 2022. ISSN 2332-7790.

Li:2021:SIB

[LLLZ21] Xiong Li, Shanpeng Liu, Rongxing Lu, and Xiaosong Zhang. On security of an identity-based dynamic data auditing protocol for big data storage. IEEE Transactions on Big Data, 7(6):975–977, December 2021. CODEN ???? ISSN 2332-7790.

Liu:2024:SUH

[LLN^{+}24] Xingbo Liu, Jiamin Li, Xiushan Nie, Xuening Zhang, Shaohua Wang, and Yilong Yin. Scalable unsupervised hashing via exploiting robust cross-modal consistency. IEEE Transactions on Big Data, 10(4):514– 527, August 2024. ISSN 2332-7790.

Liu:2022:PUW

[LLO^{+}22] Ye Liu, Yuxuan Liang, Kun Ouyang, Shuming Liu, David S. Rosenblum, and Yu Zheng. Predicting urban water quality with ubiquitous data — a datadriven approach. IEEE Transactions on Big Data, 8(2):564– 578, February 2022. CODEN ???? ISSN 2332-7790.

Liu:2023:MMH

[LLP^{+}23] Zijian Liu, Yang Luo, Xitong Pu, Geyong Min, and Chunbo Luo. A multi-modal hypergraph neural network via parametric filtering and feature sampling. IEEE Transactions on Big Data, 9(5):1365–1379, October 2023. ISSN 2332-7790.

Li:2020:CDM

[LLS^{+}20] Jianqiang Li, Lu Liu, Jingchao Sun, Haowen Mo, Ji-Jiang Yang, Shi Chen, Huiting Liu, Qing Wang, and Hui Pan. Comparison of different machine learning approaches to predict small for gestational age infants. IEEE Transactions on Big Data, 6(2):334–346, June 2020. CODEN ???? ISSN 2332-7790.

Liao:2019:ESO

[LLSL19] W. Liao, C. Luo, S. Salinas, and P. Li. Efficient secure outsourcing of large-scale convex separable programming for big data. IEEE Transactions on Big Data, 5(3):368– 378, September 2019. CODEN ???? ISSN 2332-7790.

Liu:2023:USM

[LLT^{+}23] Yimin Liu, Xiangyang Luo, Zhiyuan Tao, Meng Zhang, and Shaoyong Du. UGCC: Social media user geolocation via cyclic coupling. IEEE Transactions on Big Data, 9(4):1128– 1141, August 2023. ISSN 2332-7790.

Lin:2024:LDL

[LLW^{+}24] Yaojin Lin, Yulin Li, Chenxi Wang, Lei Guo, and Jinkun Chen. Label distribution learning based on horizontal and vertical mining of label correlations. IEEE Transactions on Big Data, 10(3):275–287, June 2024. ISSN 2332-7790.

Lu:2017:VAM

[LLY^{+}17] M. Lu, C. Lai, T. Ye, J. Liang, and X. Yuan. Visual analysis of multiple route choices based on general GPS trajectories. IEEE Transactions on Big Data, 3(2): 234–247, June 2017. CODEN ???? ISSN 2332-7790.

Liu:2018:TAD

[LLY^{+}18] H. Liu, B. Liu, L. T. Yang, M. Lin, Y. Deng, K. Bilal, and S. U. Khan. Thermalaware and DVFS-enabled big data task scheduling for data centers. IEEE Transactions on Big Data, 4(2):177–190, June 2018. CODEN ???? ISSN 2332-7790.

Li:2019:CDH

[LLZ^{+}19] Yanhua Li, Guanxiong Liu, Zhi-Li Zhang, Jun Luo, and Fan Zhang. CityLines: Designing hybrid hub-and-spoke transit system with urban Big Data. IEEE Transactions on Big Data, 5(4):576–587, December 2019. CODEN ???? ISSN 2332-7790.

Ling:2023:CFS

[LLZ^{+}23] Zhaolong Ling, Bo Li, Yiwen Zhang, Qingren Wang, Kui Yu, and Xindong Wu. Causal feature selection with efficient spouses discovery. IEEE Transactions on Big Data, 9(2):555– 568, April 2023. ISSN 2332-7790.

LaGatta:2021:ENN

[LMPS21] Valerio La Gatta, Vincenzo Moscato, Marco Postiglione, and Giancarlo Sperlí. An epidemiological neural network exploiting dynamic graph structured data applied to the COVID-19 outbreak. IEEE Transactions on Big Data, 7(1): 45–55, March 2021. CODEN ???? ISSN 2332-7790.

Li:2021:LRL

[LMW21] Zengpeng Li, Chunguang Ma, and Ding Wang. Leakage resilient leveled FHE on multiple bits message. IEEE Transactions on Big Data, 7(5):845– 858, November 2021. CODEN ???? ISSN 2332-7790.

Liakos:2022:ULH

[LND22] Panagiotis Liakos, Alexandros Ntoulas, and Alex Delis. Uncovering local hierarchical overlapping communities at scale. IEEE Transactions on Big Data, 8(2):432–445, February 2022. CODEN ???? ISSN 2332-7790.

Le:2017:LGP

[LOLL17] T. V. Le, R. Oentaryo, S. Liu, and H. C. Lau. Local Gaussian processes for efficient finegrained traffic speed prediction. IEEE Transactions on Big Data, 3(2):194–207, June 2017. CODEN ???? ISSN 2332-7790.

Li:2023:OKS

[LPT^{+}23] Mingdao Li, Peng Peng, Zhen Tian, Zheng Qin, Zheng Huang, and Yi Liu. Optimizing keyword search over federated RDF systems. IEEE Transactions on Big Data, 9(3):918–935, June 2023. ISSN 2332-7790.

Ling:2021:NAE

[LQHC21] Zenan Ling, Robert C. Qiu, Xing He, and Lei Chu. A new approach of exploiting self-adjoint matrix polynomials of large random matrices for anomaly detection and fault location. IEEE Transactions on Big Data, 7(3):548–558, July 2021. CODEN ???? ISSN 2332-7790.

Li:2023:RLT

[LQZ23] Yao Li, Duo Qiu, and Xiongjun Zhang. Robust low transformed multi-rank tensor completion with deep prior regularization for multi-dimensional image recovery. IEEE Transactions on Big Data, 9(5):1288– 1301, October 2023. ISSN 2332-7790.

Li:2020:EPQ

[LRB^{+}20] Ruiyuan Li, Sijie Ruan, Jie Bao, Yanhua Li, Yingcai Wu, Liang Hong, and Yu Zheng. Efficient path query processing over massive trajectories on the cloud. IEEE Transactions on Big Data, 6(1):66–79, March 2020. CODEN ???? ISSN 2332-7790.

Lwowski:2020:GED

[LRC20] Brandon Lwowski, Paul Rad, and Kim-Kwang Raymond Choo. Geospatial event detection by grouping emotion contagion in social media. IEEE Transactions on Big Data, 6(1): 159–170, March 2020. CODEN ???? ISSN 2332-7790.

Liu:2017:GRE

[LRL^{+}17] F. Liu, J. Rosenberger, Y. Lou, R. Hosseini, J. Su, and S. Wang. Graph regularized EEG source imaging with in-class consistency and out-class discrimination. IEEE Transactions on Big Data, 3(4):378–391, December 2017. CODEN ???? ISSN 2332-7790.

Lin:2018:SQA

[LS18a] Y. Lin and H. Shen. SmartQ:

A question and answer system for supplying high-quality and trustworthy answers. IEEE Transactions on Big Data, 4(4): 600–613, December 2018. CO-DEN ???? ISSN 2332-7790.

Liu:2018:CED

[LS18b] Y. Liu and X. Sun. CaL: Extending data locality to consider concurrency for performance optimization. IEEE Transactions on Big Data, 4(2): 273–288, June 2018. CODEN ???? ISSN 2332-7790.

Li:2023:PPF

[LSH^{+}23] Wenjing Li, Xiaodan Shi, Dou Huang, Xudong Shen, Jinyu Chen, Hill Hiroki Kobayashi, Haoran Zhang, Xuan Song, and Ryosuke Shibasaki. PredLife: Predicting fine-grained future activity patterns. IEEE Transactions on Big Data, 9(6):1658– 1669, December 2023. ISSN 2332-7790.

Li:2024:FFD

[LTG^{+}24] Li Li, Chuanqi Tao, Hongjing Guo, Jingxuan Zhang, and Xiaobing Sun. FATS: Feature distribution analysis-based test selection for deep learning enhancement. IEEE Transactions on Big Data, 10(2):132– 145, April 2024. ISSN 2332-7790.

Lin:2016:GES

[LTTC16] Y. R. Lin, H. Tong, J. Tang, and K. Selçuk Candan. Guest editorial: Special issue on big scholar data discovery and collaboration (Continued). IEEE Transactions on Big Data, 2(2): 100, June 2016. CODEN ???? ISSN 2332-7790.

Lin:2017:GEB

[LTTC17] Y. Lin, H. Tong, J. Tang, and K. S. Candan. Guest editorial: Big scholar data discovery and collaboration. IEEE Transactions on Big Data, 3(1): 2, March 2017. CODEN ???? ISSN 2332-7790.

Li:2016:PMP

[LTX16] T. Li, J. Tang, and J. Xu. Performance modeling and predictive scheduling for distributed stream data processing. IEEE Transactions on Big Data, 2(4): 353–364, December 2016. CO-DEN ???? ISSN 2332-7790.

Liu:2018:LDR

[LW18] X. Liu and X. Wang. LSdecomposition for robust recovery of sensory big data. IEEE Transactions on Big Data, 4(4): 542–555, December 2018. CO-DEN ???? ISSN 2332-7790.

Lee:2018:VAV

[LWH18] P. Lee, J. D. West, and B. Howe. Viziometrics: Analyzing visual information in the scientific literature. IEEE Transactions on Big Data, 4(1):117– 129, March 2018. CODEN ???? ISSN 2332-7790.

Liu:2023:HRL

[LWL^{+}23] Tong Liu, Jinzhen Wang, Qing Liu, Shakeel Alibhai, Tao Lu, and Xubin He. High-ratio lossy compression: Exploring the autoencoder to compress scientific data. IEEE Transactions on Big Data, 9(1):22–36, February 2023. ISSN 2332-7790.

Luo:2023:DDA

[LWQ^{+}23] Huizhang Luo, Junqi Wang, Zhenlu Qin, Dan Huang, Qing Liu, Mengchu Zhou, and Hong Jiang. A data-driven approach to harvesting latent reduced models to precondition lossy compression for scientific data. IEEE Transactions on Big Data, 9(3):949–963, June 2023. ISSN 2332-7790.

Li:2024:MSO

[LWS^{+}24] Wenrui Li, Xiaoyu Wang, Yuetian Sun, Snezana Milanovic, Mark Kon, and Julio Enrique Castrillón-Candás. Multilevel stochastic optimization for imputation in massive medical data records. IEEE Transactions on Big Data, 10(2):122– 131, April 2024. ISSN 2332-7790.

Li:2022:LPK

[LWZ^{+}22] Manling Li, Yuanzhuo Wang, Denghui Zhang, Yantao Jia, and Xueqi Cheng. Link prediction in knowledge graphs: a hierarchy-constrained approach. IEEE Transactions on Big Data, 8(3):630–643, June 2022. ISSN 2332-7790.

Li:2024:HDR

[LWZL24] Xiaodong Li, Pangjing Wu, Chenxin Zou, and Qing Li. Hierarchical deep reinforcement learning for VWAP strategy optimization. IEEE Transactions on Big Data, 10(3):288–300, June 2024. ISSN 2332-7790.

Li:2023:CES

[LXLF23] Hongjian Li, Jianglin Xia, Wei Luo, and Hai Fang. Costefficient scheduling of streaming applications in Apache Flink on cloud. IEEE Transactions on Big Data, 9(4):1086–1101, August 2023. ISSN 2332-7790.

Li:2017:DOT

[LY17] M. Li and Z. Yin. Debugging object tracking by a recommender system with correction propagation. IEEE Transactions on Big Data, 3(4):429– 442, December 2017. CODEN ???? ISSN 2332-7790.

Liang:2024:LWG

[LYC^{+}24] Naiyao Liang, Zuyuan Yang, Junhang Chen, Zhenni Li, and Shengli Xie. Label-weighted graph-based learning for semisupervised classification under label noise. IEEE Transactions on Big Data, 10(1):55–65, February 2024. ISSN 2332-7790.

Liu:2022:WSC

[LYD^{+}22] Xuanwu Liu, Guoxian Yu, Carlotta Domeniconi, Jun Wang, Guoqiang Xiao, and Maozu Guo. Weakly supervised crossmodal hashing. IEEE Transactions on Big Data, 8(2):552– 563, February 2022. CODEN ???? ISSN 2332-7790.

Liu:2021:ITT

[LYG^{+}21] Huazhong Liu, Laurence T.

Yang, Yimu Guo, Xia Xie, and Jianhua Ma. An incremental tensor-train decomposition for cyber-physical-social big data. IEEE Transactions on Big Data, 7(2):341–354, June 2021. CODEN ???? ISSN 2332-7790.

Liu:2021:DUG

[LYL^{+}21] Jiaqiang Liu, Huan Yan, Yong Li, Dmytro Karamshuk, Nishanth Sastry, Di Wu, and Depeng Jin. Discovering and understanding geographical video viewing patterns in urban neighborhoods. IEEE Transactions on Big Data, 7(5):873– 884, November 2021. CODEN ???? ISSN 2332-7790.

Lin:2022:MDT

[LYLJ22] Longlong Lin, Pingpeng Yuan, Rong-Hua Li, and Hai Jin. Mining diversified top-r lasting cohesive subgraphs on temporal networks. IEEE Transactions on Big Data, 8(6):1537–1549, December 2022. ISSN 2332-

Li:2023:LGL

[LYLW23] Pengbo Li, Hang Yu, Xiangfeng Luo, and Jia Wu. LGM-GNN: a local and global aware memorybased graph neural network for fraud detection. IEEE Transactions on Big Data, 9(4):1116– 1127, August 2023. ISSN 2332-7790.

Liu:2022:HPO

[LYW^{+}22] Dong Liu, Guoliang Yang, Yanwei Wang, Hu Jin, and Enhong Chen. How to protect ourselves from overlapping community detection in social networks. IEEE Transactions on Big Data, 8(4):894–904, August 2022. ISSN 2332-7790.

Li:2023:MBL

[LZC^{+}23] Wenjing Li, Haoran Zhang, Jinyu Chen, Peiran Li, Yuhao Yao, Xiaodan Shi, Mariko Shibasaki, Hill Hiroki Kobayashi, Xuan Song, and Ryosuke Shibasaki. Metagraph-based life pattern clustering with big human mobility data. IEEE Transactions on Big Data, 9(1): 227–240, February 2023. ISSN 2332-7790.

Liang:2023:SMF

[LZD^{+}23] Xin Liang, Kai Zhao, Sheng Di, Sihuan Li, Robert Underwood, Ali M. Gok, Jiannan Tian, Junjing Deng, Jon C. Calhoun, Dingwen Tao, Zizhong Chen, and Franck Cappello. SZ3: a modular framework for composing prediction-based error-bounded lossy compressors. IEEE Transactions on Big Data, 9(2):485–498, April 2023. ISSN 2332-7790.

Liang:2024:SST

[LZH^{+}24] Huanghuang Liang, Zheng Zhang, Chuang Hu, Yili Gong, and Dazhao Cheng. A survey on spatio-temporal big data analytics ecosystem: Resource management, processing platform, and applications. IEEE Transactions on Big Data, 10 (2):174–193, April 2024. ISSN 2332-7790.

Luo:2021:AUN

[LZL^{+}21] Xin Luo, Mengchu Zhou, Shuai Li, Di Wu, Zhigang Liu, and Mingsheng Shang. Algorithms of unconstrained non-negative latent factor analysis for recommender systems. IEEE Transactions on Big Data, 7(1):227– 240, March 2021. CODEN ???? ISSN 2332-7790.

Li:2022:FCS

[LZL^{+}22] Ting Li, Yiming Zhang, Hao Liu, Guangtao Xue, and Ling Liu. Fast compressive spectral clustering for large-scale sparse graph. IEEE Transactions on Big Data, 8(1):193–202, February 2022. CODEN ???? ISSN 2332-7790.

Liao:2022:IBL

[LZLJ22] Xiaofei Liao, Zhan Zhang, Haikun Liu, and Hai Jin. Improving bank-level parallelism for in-memory checkpointing in hybrid memory systems. IEEE Transactions on Big Data, 8(2):

289–301, February 2022. CO-DEN ???? ISSN 2332-7790.

Li:2020:CAW

[LZMZ20] Yupeng Li, Jianhua Zhang, Zhanyu Ma, and Yu Zhang. Clustering analysis in the wireless propagation channel with a variational Gaussian mixture model. IEEE Transactions on Big Data, 6(2):223–232, June 2020. CODEN ???? ISSN 2332-7790.

Leng:2021:MDC

[LZS^{+}21] Yan Leng, Yujia Zhai, Shaojing Sun, Yifei Wu, Jordan Selzer, Sharon Strover, Hezhao Zhang, Anfan Chen, and Ying Ding. Misinformation during the COVID-19 outbreak in China: Cultural, social and political entanglements. IEEE Transactions on Big Data, 7(1): 69–80, March 2021. CODEN ???? ISSN 2332-7790.

Luo:2021:SSL

[LZSL21] Changqing Luo, Kaijin Zhang, Sergio Salinas, and Pan Li. Sec-Fact: Secure large-scale QR and LU factorizations. IEEE Transactions on Big Data, 7(4):796– 807, April 2021. CODEN ???? ISSN 2332-7790.

Lv:2022:UUV

[LZW^{+}22] Guangyi Lv, Kun Zhang, Le Wu, Enhong Chen, Tong Xu, Qi Liu, and Weidong He. Understanding the users and videos by mining a novel Danmu dataset. IEEE Transactions on Big Data, 8(2):535– 551, February 2022. CODEN ???? ISSN 2332-7790.

Liu:2023:IHU

[LZW^{+}23a] Haobing Liu, Yanmin Zhu, Chunyang Wang, Jianyu Ding, Jiadi Yu, and Feilong Tang. Incorporating heterogeneous user behaviors and social influences for predictive analysis. IEEE Transactions on Big Data, 9 (2):716–732, April 2023. ISSN 2332-7790.

Liu:2023:SCG

[LZW^{+}23b] Yanbei Liu, Shichuan Zhao, Xiao Wang, Lei Geng, Zhitao Xiao, and Jerry Chun-Wei Lin. Self-consistent graph neural networks for semi-supervised node classification. IEEE Transactions on Big Data, 9(4): 1186–1197, August 2023. ISSN 2332-7790.

Li:2023:SSN

[LZZ^{+}23] Mengran Li, Yong Zhang, Wei Zhang, Yi Chu, Yongli Hu, and Baocai Yin. Self-supervised nodes-hyperedges embedding for heterogeneous information network learning. IEEE Transactions on Big Data, 9(4):1210– 1224, August 2023. ISSN 2332-7790.

Liu:2023:RRG

[LZZF23] Zhen Liu, Wenbo Zuo, Dongning Zhang, and Xiaodong Feng. RGSE: Robust graph structure embedding for anomalous link detection. IEEE Transactions on Big Data, 9(5):1420–1429, October 2023. ISSN 2332-7790.

Matsuoka:2016:GAL

[Mat16]  Satoshi Matsuoka.  GPU-

accelerated large-scale distributed sorting coping with device memory capacity. IEEE Transactions on Big Data, 2(1): 57–69, March 2016. CODEN ???? ISSN 2332-7790.

Misra:2022:SFS

[MBG22] Chandan Misra, Sourangshu Bhattacharya, and Soumya K. Ghosh. Stark: Fast and scalable Strassen’s matrix multiplication using Apache Spark. IEEE Transactions on Big Data, 8(3): 699–710, June 2022. ISSN 2332-7790.

Malensek:2019:TDS

[MBS^{+}19] M. Malensek, W. Budgaga, R. Stern, S. Pallickara, and S. L. Pallickara. Trident: Distributed storage, analysis, and exploration of multidimensional phenomena. IEEE Transactions on Big Data, 5(2):252–265, June 2019. CODEN ???? ISSN 2332-7790.

Malandrino:2020:MCT

[MCA^{+}20] Francesco Malandrino, Carla-Fabiana Chiasserini, Giuseppe Avino, Marco Malinverno, and Scott Kirkpatrick. From megabits to CPU ticks: Enriching a demand trace in the age of MEC. IEEE Transactions on Big Data, 6(1):43–50, March 2020. CODEN ???? ISSN 2332-

Meng:2022:TIT

[MCH^{+}22] Chuishi Meng, Yu Cui, Qing He, Lu Su, and Jing Gao. Towards the inference of travel purpose with heterogeneous urban data. IEEE Transactions on Big Data, 8(1):166– 177, February 2022. CODEN ???? ISSN 2332-7790.

Mohan:2022:TCM

[MCL22] Ram Mohan, Santanu Chaudhury, and Brejesh Lall. Temporal causal modelling on large volume enterprise data. IEEE Transactions on Big Data, 8 (6):1678–1689, December 2022. ISSN 2332-7790.

Mavlyutov:2021:MBI

[MCM21] Ruslan Mavlyutov and Philippe Cudre-Mauroux. Managing big interval data with CINTIA: The checkpoint INTerval array. IEEE Transactions on Big Data, 7(2):285–298, June 2021. CODEN ???? ISSN 2332-7790.

Ma:2022:EIG

[MCO^{+}22] Mengyu Ma, Luo Chen, Xue Ouyang, Xiaoran Liu, Jun Li, and Ning Jing. Efficient interactive global cellular signal strength visualization. IEEE Transactions on Big Data, 8(5): 1209–1219, October 2022. ISSN 2332-7790.

Mavroudopoulos:2023:SSI

[MG23] Ioannis Mavroudopoulos and Anastasios Gounaris. SIESTA: a scalable infrastructure of sequential pattern analysis. IEEE Transactions on Big Data, 9(3): 975–990, June 2023. ISSN 2332-7790.

Mahmud:2023:ACE

[MHR^{+}23] Mohammad Sultan Mahmud, Joshua Zhexue Huang, Rukhsana Ruby, Alladoumbaye Ngueilbaye, and Kaishun Wu. Approximate clustering ensemble method for Big Data. IEEE Transactions on Big Data, 9(4): 1142–1155, August 2023. ISSN 2332-7790.

Mujtaba:2024:FGF

[MKJ^{+}24] Ghulam Mujtaba, Sunder Ali Khowaja, Muhammad Aslam Jarwar, Jaehyuk Choi, and Eun-Seok Ryu. FRC-GIF: Frame ranking-based personalized artistic media generation method for resource constrained devices. IEEE Transactions on Big Data, 10(4):343–355, August 2024. ISSN 2332-7790.

Myagmar:2021:HDL

[MLK21] Batsergelen Myagmar, Jie Li, and Shigetomo Kimura. Heterogeneous daily living activity learning through domain invariant feature subspace. IEEE Transactions on Big Data, 7(6): 922–929, December 2021. CO-DEN ???? ISSN 2332-7790.

Makkie:2019:DCP

[MLQ^{+}19] M. Makkie, X. Li, S. Quinn, B. Lin, J. Ye, G. Mon, and T. Liu. A distributed computing platform for fMRI big data analytics. IEEE Transactions on Big Data, 5(2):109– 119, June 2019. CODEN ???? ISSN 2332-7790.

Moustafa:2019:NGA

[MSC19] Nour Moustafa, Jill Slay, and Gideon Creech. Novel geometric area analysis technique for anomaly detection using trapezoidal area estimation on largescale networks. IEEE Transactions on Big Data, 5(4):481– 494, December 2019. CODEN ???? ISSN 2332-7790.

Minetto:2021:MHE

[MSRS21] Rodrigo Minetto, Maurício Pamplona Segundo, Gilbert Rotich, and Sudeep Sarkar. Measuring human and economic activity from satellite imagery to support city-scale decisionmaking during COVID-19 pandemic. IEEE Transactions on Big Data, 7(1):56–68, March 2021. CODEN ???? ISSN 2332-7790.

Maurya:2021:DSC

[MTV21] Chandresh Kumar Maurya, Durga Toshniwal, and Gopalan Vijendran Venkoparao. Distributed sparse class-imbalance learning and its applications. IEEE Transactions on Big Data, 7(5):832–844, November 2021. CODEN ???? ISSN 2332-7790.

Ma:2022:EEI

[MW22] Meng Ma and Ping Wang.

Efficient event inference and context-awareness in Internet of Things edge systems. IEEE Transactions on Big Data, 8(3):

658–670, June 2022. ISSN 2332-7790.

Ma:2020:DNN

[MXL^{+}20] Zhanyu Ma, Jiyang Xie, Hailong Li, Qie Sun, Fredrik Wallin, Zhongwei Si, and Jun Guo. Deep neural networkbased impacts analysis of multimodal factors on heat demand prediction. IEEE Transactions on Big Data, 6(3):594– 605, September 2020. CODEN ???? ISSN 2332-7790.

Ma:2024:CKL

[MZF24] Wenjun Ma, Yibing Zhao, and Xiaomao Fan. Cascaded knowledge-level fusion network for online course recommendation system. IEEE Transactions on Big Data, 10(4):457– 469, August 2024. ISSN 2332-7790.

Nie:2017:MTI

[NCS17] W. Nie, H. Cheng, and Y. Su. Modeling temporal information of mitotic for mitotic event detection. IEEE Transactions on Big Data, 3(4):458–469, December 2017. CODEN ???? ISSN 2332-7790.

Ng:2016:UUS

[Ng16] Wilfred Ng. USTF: A unified system of team formation. IEEE Transactions on Big Data, 2(1):70–84, March 2016. CODEN ???? ISSN 2332-7790. Nabavinejad:2016:MCR

[NGM16] S. M. Nabavinejad, M. Goudarzi, and S. Mozaffari. The memory challenge in reduce phase of MapReduce applications. IEEE Transactions on Big Data, 2(4): 380–386, December 2016. CO-DEN ???? ISSN 2332-7790.

Niu:2018:JTC

[NHL18] Z. Niu, B. He, and F. Liu.

JouleMR: Towards cost-effective and green-aware data processing frameworks. IEEE Transactions on Big Data, 4(2):258– 272, June 2018. CODEN ???? ISSN 2332-7790.

Ni:2015:SPS

[Ni15] Lionel M. Ni. SMC: A practical schema for privacy-preserved data sharing over distributed data streams. IEEE Transactions on Big Data, 1(2):68–81, June 2015. CODEN ???? ISSN 2332-7790.

Nguyen:2017:DCP

[NLC17] H. Nguyen, W. Liu, and F. Chen. Discovering congestion propagation patterns in spatio-temporal traffic data. IEEE Transactions on Big Data, 3(2):169–180, June 2017. CODEN ???? ISSN 2332-7790.

Nie:2023:SDM

[NLG^{+}23] Xiushan Nie, Xingbo Liu, Jie Guo, Letian Wang, and Yilong Yin. Supervised discrete multiple-length hashing for image retrieval. IEEE Transactions on Big Data, 9(1):312– 327, February 2023. ISSN 2332-7790.

Niu:2022:OTW

[NLY22] Xiaofei Niu, Guangchi Liu, and Qing Yang. OpinionRank:

Trustworthy website detection using three valued subjective logic. IEEE Transactions on Big Data, 8(3):855–866, June 2022. ISSN 2332-7790.

Nguyen:2022:FPN

[NYL^{+}22] Long Nguyen, Zhou Yang, Jia Li, Zhenhe Pan, Guofeng Cao, and Fang Jin. Forecasting people’s needs in hurricane events from social network. IEEE Transactions on Big Data, 8(1): 229–240, February 2022. CO-DEN ???? ISSN 2332-7790.

Oh:2024:MTW

[ONP^{+}24] Seungeun Oh, Hyelin Nam, Jihong Park, Praneeth Vepakomma, Ramesh Raskar, Mehdi Bennis, and Seong-Lyun Kim. Mix2SFL: Two-way mixup for scalable, accurate, and communication-efficient split federated learning. IEEE Transactions on Big Data, 10 (3):238–248, June 2024. ISSN 2332-7790.

Ota:2017:SST

[OVSF17] M. Ota, H. Vo, C. Silva, and J. Freire. STaRS: Simulating taxi ride sharing at scale. IEEE Transactions on Big Data, 3(3): 349–361, September 2017. CO-DEN ???? ISSN 2332-7790.

Perrot:2020:CTH

[PA20] Alexandre Perrot and David Auber. Cornac: Tackling huge graph visualization with Big Data infrastructure. IEEE Transactions on Big Data, 6(1):

80–92, March 2020. CODEN ???? ISSN 2332-7790.

Pejo:2023:QIF

[PB23] Balázs Pejó and Gergely Biczók. Quality inference in federated learning with secure aggregation. IEEE Transactions on Big Data, 9(5):1430–1437, October 2023. ISSN 2332-7790.

Phinyomark:2017:RSF

[PIMP17] A. Phinyomark, E. Ibáñez-Marcelo, and G. Petri. Restingstate fMRI functional connectivity: Big data preprocessing pipelines and topological data analysis. IEEE Transactions on Big Data, 3(4):415–428, December 2017. CODEN ???? ISSN 2332-7790.

Pan:2022:SSC

[PLZL22] Menghai Pan, Yanhua Li, Zhi-Li Zhang, and Jun Luo. SCCS: Smart cloud commuting system with shared autonomous vehicles. IEEE Transactions on Big Data, 8(5):1301–1311, October 2022. ISSN 2332-7790.

Poostchi:2022:BST

[PP22] Hanieh Poostchi and Massimo Piccardi. BiLSTM-SSVM: Training the BiLSTM with a structured hinge loss for namedentity recognition. IEEE Transactions on Big Data, 8(1):203– 212, February 2022. CODEN ???? ISSN 2332-7790.

Parmar:2022:SMR

[PRR22] Nishaal Parmar, Hazem H. Refai, and Thordur Runolfsson. A survey on the methods and results of data-driven Koopman analysis in the visualization of dynamical systems. IEEE Transactions on Big Data, 8(3): 723–738, June 2022. ISSN 2332-7790.

Peng:2021:CCC

[PTL^{+}21] Yifan Peng, Yuxing Tang, Sungwon Lee, Yingying Zhu, Ronald M. Summers, and Zhiyong Lu. COVID-19-CT-CXR: a freely accessible and weakly labeled chest X-ray and CT image collection on COVID-19 from biomedical literature. IEEE Transactions on Big Data, 7 (1):3–12, March 2021. CODEN ???? ISSN 2332-7790.

Pang:2022:SCP

[PWC^{+}22] Zhifei Pang, Sai Wu, Gang Chen, Lidan Shou, Ke Chen, and Bingsheng He. A stackcentric processing model for iterative processing. IEEE Transactions on Big Data, 8(2):318– 331, February 2022. CODEN ???? ISSN 2332-7790.

Pan:2016:ABS

[PWH16] E. Pan, D. Wang, and Z. Han. Analyzing big smart metering data towards differentiated user services: A sublinear approach. IEEE Transactions on Big Data, 2(3):249– 261, September 2016. CODEN ???? ISSN 2332-7790.

Puthal:2019:SSE

[PWS^{+}19] D. Puthal, X. Wu, N. Surya, R. Ranjan, and J. Chen. SEEN:

A selective encryption method to ensure confidentiality for big sensing data streams. IEEE Transactions on Big Data, 5(3): 379–392, September 2019. CO-DEN ???? ISSN 2332-7790.

Pal:2023:SMD

[PWW^{+}23] Amitangshu Pal, Junbo Wang, Yilang Wu, Krishna Kant, Zhi Liu, and Kento Sato. Social media driven big data analysis for disaster situation awareness: a tutorial. IEEE Transactions on Big Data, 9(1):1–21, February 2023. ISSN 2332-7790.

Qiao:2023:RTT

[QFW^{+}23] Ziyue Qiao, Yanjie Fu, Pengyang Wang, Meng Xiao, Zhiyuan Ning, Denghui Zhang, Yi Du, and Yuanchun Zhou. RPT: Toward transferable model on heterogeneous researcher data via pre-training. IEEE Transactions on Big Data, 9(1):186– 199, February 2023. ISSN 2332-7790.

Qi:2022:DDW

[QHC^{+}22] Lianyong Qi, Qiang He, Feifei Chen, Xuyun Zhang, Wanchun Dou, and Qiang Ni. Datadriven web APIs recommendation for building Web applications. IEEE Transactions on Big Data, 8(3):685–698, June 2022. ISSN 2332-7790.

Qin:2022:MSC

[QLW^{+}22] Hongchao Qin, Rong-Hua Li, Guoren Wang, Xin Huang, Ye Yuan, and Jeffrey Xu Yu. Mining stable communities in temporal networks by densitybased clustering. IEEE Transactions on Big Data, 8(3):671– 684, June 2022. ISSN 2332-7790.

Qiao:2023:MQC

[QLZ^{+}23] Lianpeng Qiao, Rong-Hua Li, Zhiwei Zhang, Ye Yuan, Guoren Wang, and Hongchao Qin. Maximal quasi-cliques mining in uncertain graphs. IEEE Transactions on Big Data, 9(1): 37–50, February 2023. ISSN 2332-7790.

Qin:2024:AAP

[QLZ24] Wen Qin, Xin Luo, and MengChu Zhou. Adaptivelyaccelerated parallel stochastic gradient descent for highdimensional and incomplete data representation learning. IEEE Transactions on Big Data, 10(1):92–107, February 2024. ISSN 2332-7790.

Qiu:2019:TGC

[QSC19] C. Qiu, H. Shen, and L. Chen. Towards green cloud computing: Demand allocation and pricing policies for cloud service brokerage. IEEE Transactions on Big Data, 5(2):238– 251, June 2019. CODEN ???? ISSN 2332-7790.

Qin:2023:IBE

[QWL^{+}23a] Yixiu Qin, Yizhao Wang, Jiawei Li, Shun Mao, He Wang, and Yuncheng Jiang. Improved box embeddings for fine-grained entity typing. IEEE Transactions on Big Data, 9(6):1631– 1642, December 2023. ISSN 2332-7790.

Qiu:2023:CSC

[QWL^{+}23b] Yu-Xuan Qiu, Dong Wen, Rong-Hua Li, Lu Qin, Michael Yu, and Xuemin Lin. Computing significant cliques in large labeled networks. IEEE Transactions on Big Data, 9(3):904– 917, June 2023. ISSN 2332-7790.

Qi:2018:SBT

[QXZ^{+}18] L. Qi, X. Xu, X. Zhang, W. Dou, C. Hu, Y. Zhou, and J. Yu. Structural balance theory-based e-commerce recommendation over big rating data. IEEE Transactions on Big Data, 4(3):301– 312, September 2018. CODEN ???? ISSN 2332-7790.

Qian:2023:ETA

[QYC^{+}23] Fulan Qian, Bei Yuan, Hai Chen, Jie Chen, Defu Lian, and Shu Zhao. Enhancing the transferability of adversarial examples based on Nesterov momentum for recommendation systems. IEEE Transactions on Big Data, 9(5):1276–1287, October 2023. ISSN 2332-7790.

Rammer:2022:EFE

[RBM^{+}22] Daniel Rammer, Thilina Buddhika, Matthew Malensek, Shrideep Pallickara, and Sangmi Lee Pallickara. Enabling fast exploratory analyses over voluminous spatiotemporal data using analytical engines. IEEE Transactions on Big Data, 8(1):213– 228, February 2022. CODEN ???? ISSN 2332-7790.

Raji:2021:MVS

[RDD^{+}21] Mohammad Raji, John Duggan, Blaise DeCotes, Jian Huang, and Bradley Vander Zanden. Modeling and visualizing student flow. IEEE Transactions on Big Data, 7(3):510– 523, July 2021. CODEN ???? ISSN 2332-7790.

Rezgui:2020:CSP

[RDM^{+}20] Abdelmounaam Rezgui, Nickolas Davis, Zaki Malik, Brahim Medjahed, and Hamdy S. Soliman. CloudFinder: a system for processing Big Data workloads on volunteered federated clouds. IEEE Transactions on Big Data, 6(2):347–358, June 2020. CODEN ???? ISSN 2332-7790.

Raghavan:2022:GTA

[RJS22] Krishnan Raghavan, Sarangapani Jagannathan, and V. A. Samaranayake. A game theoretic approach for addressing domain-shift in big-data. IEEE Transactions on Big Data, 8 (6):1610–1621, December 2022. ISSN 2332-7790.

Sainju:2020:PGB

[SAJP20] Arpan Man Sainju, Danial Aghajarian, Zhe Jiang, and Sushil Prasad. Parallel gridbased colocation mining algorithms on GPUs for big spatial event data. IEEE Transactions on Big Data, 6(1):107– 118, March 2020. CODEN ???? ISSN 2332-7790.

Sakurama:2020:CLS

[Sak20] Kazunori Sakurama. Control of large-scale cyber-physical systems with agents having various dynamics. IEEE Transactions on Big Data, 6(4):691– 701, December 2020. CODEN ???? ISSN 2332-7790.

Shah:2020:SDS

[SAM^{+}20] Zubair Shah, Adnan Anwar, Abdun Naser Mahmood, Zahir Tari, and Albert Y. Zomaya. A spatiotemporal data summarization approach for real-time operation of Smart Grid. IEEE Transactions on Big Data, 6(4): 624–637, December 2020. CO-DEN ???? ISSN 2332-7790.

Su:2022:TAP

[SBL^{+}22] Ziyi Su, Frédérique Biennier, Zhihan Lv, Yong Peng, Houbing Song, and Jingwei Miao. Toward architectural and protocol-level foundation for end-to-end trustworthiness in cloud/fog computing. IEEE Transactions on Big Data, 8(1): 35–47, February 2022. CODEN ???? ISSN 2332-7790.

Slo:2022:SAL

[SBR22] Ahmad Slo, Sukanya Bhowmik, and Kurt Rothermel. Stateaware load shedding from input event streams in complex event processing. IEEE Transactions on Big Data, 8(5):1340– 1357, October 2022. ISSN 2332-

Sarkar:2017:EUS

[SCA^{+}17] S. Sarkar, S. Chawla, S. Ahmad, J. Srivastava, H. Hammady, F. Filali, W. Znaidi, and J. Borge-Holthoefer. Effective urban structure inference from traffic flow dynamics. IEEE Transactions on Big Data, 3(2): 181–193, June 2017. CODEN ???? ISSN 2332-7790.

Sun:2023:HLM

[SCG^{+}23] Gan Sun, Yang Cong, Changjun Gu, Xu Tang, Zhengming Ding, and Haibin Yu. Hierarchical lifelong machine learning with “Watchdog”. IEEE Transactions on Big Data, 9(1):63–74, February 2023. ISSN 2332-7790.

Syrris:2020:MCS

[SCPS20] Vasileios Syrris, Christina Corbane, Martino Pesaresi, and Pierre Soille. Mosaicking Copernicus Sentinel-1 data at global scale. IEEE Transactions on Big Data, 6(3):547– 557, September 2020. CODEN ???? ISSN 2332-7790.

Shao:2018:SMC

[SCW18] Z. Shao, J. Cai, and Z. Wang. Smart monitoring cameras driven intelligent processing to big surveillance video data. IEEE Transactions on Big Data, 4(1):105–116, March 2018. CODEN ???? ISSN 2332-7790.

Sellami:2018:CQO

[SD18] R. Sellami and B. Defude. Complex queries optimization and evaluation over relational and NoSQL data stores in cloud environments. IEEE Transactions on Big Data, 4(2):217– 230, June 2018. CODEN ???? ISSN 2332-7790.

Shah:2022:EDT

[SD22] Zubair Shah and Adam G.

Dunn. Event detection on Twitter by mapping unexpected changes in streaming data into a spatiotemporal lattice. IEEE Transactions on Big Data, 8(2): 508–522, February 2022. CO-DEN ???? ISSN 2332-7790.

Shukla:2021:DSS

[SDR^{+}21] Manu Shukla, Dinesh Dharme, Pallavi Ramnarain, Ray Dos Santos, and Chang-Tien Lu. DIGDUG: Scalable separable dense graph pruning and join operations in MapReduce. IEEE Transactions on Big Data, 7(6):930–951, December 2021. CODEN ???? ISSN 2332-7790.

Servetnyk:2023:DDA

[SF23] Mykola Servetnyk and Carrson C.

Fung. Distributed dual averaging based data clustering. IEEE Transactions on Big Data, 9(1): 372–379, February 2023. ISSN 2332-7790.

Sidhanta:2021:DAC

[SGM21] Subhajit Sidhanta, Wojciech Golab, and Supratik Mukhopadhyay. Deadline-aware cost optimization for Spark. IEEE Transactions on Big Data, 7(1):

115–127, March 2021. CODEN ???? ISSN 2332-7790.

Sidhanta:2017:ASA

[SGMB17] S. Sidhanta, W. Golab, S. Mukhopadhyay, and S. Basu. Adaptable SLA-aware consistency tuning for quorum-replicated datastores. IEEE Transactions on Big Data, 3(3):248– 261, September 2017. CODEN ???? ISSN 2332-7790.

Seebacher:2021:VAS

[SHH^{+}21] Daniel Seebacher, Johannes Häußler, Michael Hundt, Manuel Stein, Hannes Müller, Ulrich Engelke, and Daniel A. Keim. Visual analysis of spatiotemporal event predictions: Investigating the spread dynamics of invasive species. IEEE Transactions on Big Data, 7(3):497– 509, July 2021. CODEN ???? ISSN 2332-7790.

Shi:2022:LSS

[SHL^{+}22] Xiaoyu Shi, Qiang He, Xin Luo, Yanan Bai, and Mingsheng Shang. Large-scale and scalable latent factor analysis via distributed alternative stochastic gradient descent for recommender systems. IEEE Transactions on Big Data, 8(2):420– 431, February 2022. CODEN ???? ISSN 2332-7790.

Schneider:2021:IDM

[SJS^{+}21] Bruno Schneider, Dominik Jäckle, Florian Stoffel, Alexandra Diehl, Johannes Fuchs, and Daniel Keim. Integrating data and model space in ensemble learning by visual analytics. IEEE Transactions on Big Data, 7(3):483–496, July 2021. CODEN ???? ISSN 2332-7790.

Salinas:2018:ESO

[SLC^{+}18] S. Salinas, C. Luo, X. Chen, W. Liao, and P. Li. Efficient secure outsourcing of large-scale sparse linear systems of equations. IEEE Transactions on Big Data, 4(1):26–39, March 2018. CODEN ???? ISSN 2332-7790. See [CM21].

Shen:2021:SAC

[SLL^{+}21] Jian Shen, Dengzhi Liu, Qi Liu, Xingming Sun, and Yan Zhang. Secure authentication in cloud big data with hierarchical attribute authorization structure. IEEE Transactions on Big Data, 7(4):668–677, April 2021. CODEN ???? ISSN 2332-7790.

Song:2022:DCN

[SLL^{+}22] Guojie Song, Qingqing Long, Yi Luo, Yiming Wang, and Yilun Jin. Deep convolutional neural network based medical concept normalization. IEEE Transactions on Big Data, 8(5): 1195–1208, October 2022. ISSN 2332-7790.

Shi:2022:VAA

[SLT^{+}22] Yang Shi, Yuyin Liu, Hanghang Tong, Jingrui He, Gang Yan, and Nan Cao. Visual analytics of anomalous user behaviors: a survey. IEEE Transactions on Big Data, 8(2):377–396, February 2022. CODEN ???? ISSN 2332-7790.

Shelestov:2020:CAA

[SLV^{+}20] Andrii Shelestov, Mykola Lavreniuk, Vladimir Vasiliev, Leonid Shumilo, Andrii Kolotii, Bohdan Yailymov, Nataliia Kussul, and Hanna Yailymova. Cloud approach to automated crop classification using Sentinel-1 imagery. IEEE Transactions on Big Data, 6(3): 572–582, September 2020. CO-DEN ???? ISSN 2332-7790.

Shen:2017:SOS

[SLWV17] H. Shen, G. Liu, H. Wang, and N. Vithlani. SocialQ&A: An online social network based question and answer system. IEEE Transactions on Big Data, 3(1): 91–106, March 2017. CODEN ???? ISSN 2332-7790.

Si:2022:GES

[SLZ^{+}22] Beibei Si, Yuxuan Liang, Jin Zhao, Yu Zhang, Xiaofei Liao, Hai Jin, Haikun Liu, and Lin Gu. GGraph: an efficient structure-aware approach for iterative graph processing. IEEE Transactions on Big Data, 8(5): 1182–1194, October 2022. ISSN 2332-7790.

Siriweera:2023:CDC

[SPH23] Akila Siriweera, Incheon Paik, and Huawei Huang. Constraintdriven complexity-aware data science workflow for AutoBDA. IEEE Transactions on Big Data, 9(6):1438–1457, December 2023. ISSN 2332-7790.

Srinivasan:2022:SMA

[SPN^{+}22] Sriram Srinivasan, Samuel D.

Pollard, Boyana Norris, Sajal K. Das, and Sanjukta Bhowmick. A shared-memory algorithm for updating treebased properties of large dynamic networks. IEEE Transactions on Big Data, 8(2):302– 317, February 2022. CODEN ???? ISSN 2332-7790.

Schmedding:2023:ESM

[SPYS23] Anna Schmedding, Riccardo Pinciroli, Lishan Yang, and Evgenia Smirni. Epidemic spread modeling for COVID-19 using cross-fertilization of mobility data. IEEE Transactions on Big Data, 9(5):1260–1275, October 2023. ISSN 2332-7790.

Singh:2017:DSD

[SRM17] D. Singh, D. Roy, and C. K.

Mohan. DiP-SVM: Distribution preserving kernel support vector machine for big data. IEEE Transactions on Big Data, 3(1): 79–90, March 2017. CODEN ???? ISSN 2332-7790.

Soda:2021:GIA

[SSAI21] Paolo Soda, Rosa Sicilia, Ludovica Acciai, and Giulio Iannello. Grasping inter-attribute and temporal variability in multivariate time series. IEEE Transactions on Big Data, 7(5): 885–892, November 2021. CO-DEN ???? ISSN 2332-7790.

Shao:2016:CBS

[SSSB16] W. Shao, F. D. Salim, A. Song, and A. Bouguettaya. Clustering big spatiotemporal-interval data. IEEE Transactions on Big Data, 2(3):190–203, September 2016. CODEN ???? ISSN 2332-7790.

Song:2022:GES

[SSSL22] Guojie Song, Chuan Shi, Yizhou Sun, and Zhiyuan Liu. Guest editorial: Special issue on social media computing. IEEE Transactions on Big Data, 8(4): 953–954, August 2022. ISSN 2332-7790.

Sun:2023:NED

[SSY^{+}23] Xin Sun, Zenghui Song, Yongbo Yu, Junyu Dong, Claudia Plant, and Christian Böhm. Network embedding via deep prediction model. IEEE Transactions on Big Data, 9(2):455–470, April 2023. ISSN 2332-7790.

Stivaktakis:2020:CNN

[STM^{+}20] Radamanthys Stivaktakis, Grigorios Tsagkatakis, Bruno Moraes, Filipe Abdalla, Jean-Luc Starck, and Panagiotis Tsakalides. Convolutional neural networks for spectroscopic redshift estimation on Euclid data. IEEE Transactions on Big Data, 6(3): 460–476, September 2020. CO-DEN ???? ISSN 2332-7790.

Sangat:2022:AGJ

[STM22] Prajwol Sangat, David Taniar, and Christopher Messom. ATrie group join: a parallel star group join and aggregation for inmemory column-stores. IEEE Transactions on Big Data, 8(4):

1020–1033, August 2022. ISSN 2332-7790.

Shi:2022:MLG

[STZL22] Min Shi, Yufei Tang, Xingquan Zhu, and Jianxun Liu. Multilabel graph convolutional network representation learning. IEEE Transactions on Big Data, 8(5):1169–1181, October 2022. ISSN 2332-7790.

Sun:2015:CCH

[Sun15] Zhenan Sun. Code consistent hashing based on informationtheoretic criterion. IEEE Transactions on Big Data, 1(3): 84–94, September 2015. CO-DEN ???? ISSN 2332-7790.

Sheng:2016:GEB

[SVYY16] M. Sheng, A. V. Vasilakos, Q. Yu, and L. You. Guest editorial: Big data analytics and the Web. IEEE Transactions on Big Data, 2(3):189, September 2016. CODEN ???? ISSN 2332-7790.

Shen:2022:TEB

[SWC^{+}22] Jian Shen, Chen Wang, Aniello Castiglione, Dengzhi Liu, and Christian Esposito. Trustworthiness evaluation-based routing protocol for incompletely predictable vehicular ad hoc networks. IEEE Transactions on Big Data, 8(1):48–59, February 2022. CODEN ???? ISSN 2332-7790.

Sun:2020:GEB

[SWDX20] Peng Sun, Yonggang Wen, Ta Nguyen Binh Duong, and Xiaokui Xiao. GraphMP: I/Oefficient big graph analytics on a single commodity machine. IEEE Transactions on Big Data, 6(4):816–829, December 2020. CODEN ???? ISSN 2332-7790.

Shao:2016:SNN

[SWF16] M. Shao, X. Wu, and Y. Fu. Scalable nearest neighbor sparse graph approximation by exploiting graph structure. IEEE Transactions on Big Data, 2(4): 365–380, December 2016. CO-DEN ???? ISSN 2332-7790.

Sun:2022:GON

[SWH^{+}22] Peng Sun, Yonggang Wen, Ruobing Han, Wansen Feng, and Shengen Yan. Gradient-Flow: Optimizing network performance for large-scale distributed DNN training. IEEE Transactions on Big Data, 8(2): 495–507, February 2022. CO-DEN ???? ISSN 2332-7790.

Sun:2018:MSM

[SWTX18] P. Sun, Y. Wen, D. N. B. Ta, and H. Xie. MetaFlow: A scalable metadata lookup service for distributed file systems in data centers. IEEE Transactions on Big Data, 4(2):203– 216, June 2018. CODEN ???? ISSN 2332-7790.

Su:2021:SAR

[SX21] Zhou Su and Qichao Xu. Security-aware resource allocation for mobile social big data: A matching-coalitional game solution. IEEE Transactions on Big Data, 7(4):632–642, April 2021. CODEN ???? ISSN 2332-7790.

Song:2023:SBB

[SXS^{+}23] Rui Song, Bin Xiao, Yubo Song, Songtao Guo, and Yuanyuan Yang. A survey of blockchainbased schemes for data sharing and exchange. IEEE Transactions on Big Data, 9(6):1477– 1495, December 2023. ISSN 2332-7790.

Sun:2023:BVS

[SXYL23] Yujing Sun, Hao Xiong, Siu Ming Yiu, and Kwok Yan Lam. BitAnalysis: a visualization system for bitcoin wallet investigation. IEEE Transactions on Big Data, 9(2):621–636, April 2023. ISSN 2332-7790.

Shuai:2018:QJS

[SYS^{+}18] H. Shuai, D. Yang, C. Shen, P. S. Yu, and M. Chen. QM-Sampler: Joint sampling of multiple networks with quality guarantee. IEEE Transactions on Big Data, 4(1):90– 104, March 2018. CODEN ???? ISSN 2332-7790.

SoltaniPanah:2019:CDG

[SYv^{+}19] A. Soltani Panah, A. Yavari, R. van Schyndel, D. Georgakopoulos, and X. Yi. Contextdriven granular disclosure control for Internet of Things applications. IEEE Transactions on Big Data, 5(3):408–422, September 2019. CODEN ???? ISSN 2332-7790.

Song:2022:BDE

[SZA^{+}22] Xuan Song, Haoran Zhang, Rajendra Akerkar, Huawei Huang, Song Guo, Lei Zhong, Yusheng Ji, Andreas L. Opdahl, Hemant Purohit, André Skupin, Akshay Pottathil, and Aron Culotta. Big data and emergency management: Concepts, methodologies, and applications. IEEE Transactions on Big Data, 8(2): 397–419, February 2022. CO-DEN ???? ISSN 2332-7790.

Shang:2021:IBD

[SZC^{+}21] Tao Shang, Feng Zhang, Xingyue Chen, Jianwei Liu, and Xinxi Lu. Identity-based dynamic data auditing for big data storage. IEEE Transactions on Big Data, 7(6):913–921, December 2021. CODEN ???? ISSN 2332-7790.

Song:2023:TEC

[SZG^{+}23] Jie Song, Xingchen Zhao, Chaopeng Guo, Yu Gu, and Ge Yu. Towards an energy complexity model for distributed data processing algorithms. IEEE Transactions on Big Data, 9(6):1510–1524, December 2023. ISSN 2332-7790.

Sun:2022:ETS

[SZJ^{+}22] Guodao Sun, Sujia Zhu, Qi Jiang, Wang Xia, and Ronghua Liang. EvoSets: Tracking the sensitivity of dimensionality reduction results across subspaces. IEEE Transactions on Big Data, 8(6):1566– 1579, December 2022. ISSN 2332-7790.

Shi:2024:DME

[SZSY24] Yu Shi, Xi Zhang, Yuming Shang, and Ning Yu. Don’t be misled by emotion! disentangle emotions and semantics for cross-language and crossdomain rumor detection. IEEE Transactions on Big Data, 10 (3):249–259, June 2024. ISSN 2332-7790.

Song:2022:DAN

[SZXL22] Guojie Song, Yizhou Zhang, Lingjun Xu, and Haibing Lu. Domain adaptive network embedding. IEEE Transactions on Big Data, 8(5):1220–1232, October 2022. ISSN 2332-7790.

Shi:2018:EPM

[SZY^{+}18] W. Shi, Y. Zhu, P. S. Yu, J. Zhang, T. Huang, C. Wang, and Y. Chen. Effective prediction of missing data on Apache Spark over multivariable time series. IEEE Transactions on Big Data, 4(4):473–486, December 2018. CODEN ???? ISSN 2332-7790.

Shen:2022:KEM

[SZY^{+}22] Ying Shen, Yilin Zhang, Kaiqi Yuan, Dagang Li, and Haitao Zheng. A knowledge-enhanced multi-view framework for drugtarget interaction prediction. IEEE Transactions on Big Data, 8(5):1387–1398, October 2022. ISSN 2332-7790.

Sun:2023:AMO

[SZZ^{+}23] Guodao Sun, Zihao Zhu, Gefei Zhang, Chaoqing Xu, Yunchao Wang, Sujia Zhu, Baofeng Chang, and Ronghua Liang. Application of mathematical optimization in data visualization and visual analytics: a survey. IEEE Transactions on Big Data, 9(4):1018–1037, August 2023. ISSN 2332-7790.

Su:2024:FSL

[SZZW24] Yuling Su, Hong Zhao, Yifeng Zheng, and Yu Wang. Few-shot learning with multi-granularity knowledge fusion and decisionmaking. IEEE Transactions on Big Data, 10(4):486–497, August 2024. ISSN 2332-7790.

Tian:2017:DIS

[TAL^{+}17] Y. Tian, I. Alagiannis, E. Liarou, A. Ailamaki, P. Michiardi, and M. Vukoli . DiNoDB: An interactive-speed query engine for ad-hoc queries on temporary data. IEEE Transactions on Big Data, 3(3):320– 333, September 2017. CODEN ???? ISSN 2332-7790.

Tang:2019:BDC

[TAL19] M. Tang, M. Alazab, and Y. Luo. Big data for cybersecurity: Vulnerability disclosure trends and dependencies. IEEE Transactions on Big Data, 5(3): 317–329, September 2019. CO-DEN ???? ISSN 2332-7790.

Tang:2020:MIE

[Tan20] Jie Tang. Message from the incoming Editor-in-Chief. IEEE Transactions on Big Data, 6(1): 2, March 2020. CODEN ???? ISSN 2332-7790.

Tao:2015:EIS

[Tao15]  Dacheng Tao.  Exploration

of image search results quality assessment. IEEE Transactions on Big Data, 1(3):95–108, September 2015. CODEN ???? ISSN 2332-7790.

Tsou:2021:RER

[TCC21] Yao-Tung Tsou, Hung-Li Chen, and Jia-Yang Chen. RoD: Evaluating the risk of data disclosure using noise estimation for differential privacy. IEEE Transactions on Big Data, 7(1): 214–226, March 2021. CODEN ???? ISSN 2332-7790.

Tang:2017:SLH

[TEOS17] X. Tang, E. Eftelioglu, D. Oliver, and S. Shekhar. Significant linear hotspot discovery. IEEE Transactions on Big Data, 3(2): 140–153, June 2017. CODEN ???? ISSN 2332-7790.

Tang:2022:WSC

[TF22] Bing Tang and Gilles Fedak.

WukaStore: Scalable, configurable and reliable data storage on hybrid volunteered cloud and desktop systems. IEEE Transactions on Big Data, 8(1):85– 98, February 2022. CODEN ???? ISSN 2332-7790.

Tang:2022:RRL

[TJZ^{+}22] Zhiqing Tang, Weijia Jia, Xiaojie Zhou, Wenmian Yang, and Yongjian You. Representation and reinforcement learning for task scheduling in edge computing. IEEE Transactions on Big Data, 8(3):795–808, June 2022. ISSN 2332-7790.

Tran:2023:PII

[TLL^{+}23] Ha Xuan Tran, Thuc Duy Le, Jiuyong Li, Lin Liu, Jixue Liu, Yanchang Zhao, and Tony Waters. Personalized interventions to increase the employment success of people with disability. IEEE Transactions on Big Data, 9(6):1561–1574, December 2023. ISSN 2332-7790.

Tong:2024:EMS

[TLS24] Guoxiang Tong, Quanquan Li, and Yan Song. Enhanced multiscale features mutual mapping fusion based on reverse knowledge distillation for industrial anomaly detection and localization. IEEE Transactions on Big Data, 10(4):498–513, August 2024. ISSN 2332-7790.

Tian:2023:DIH

[TNX23] Xing Tian, Wing W. Y. Ng, and Huihui Xu. Deep incremental hashing for semantic image retrieval with concept drift. IEEE Transactions on Big Data, 9(4): 1102–1115, August 2023. ISSN 2332-7790.

Thao:2021:OSS

[TRB^{+}21] Tran Phuong Thao, Mohammad Shahriar Rahman, Md Zakirul Alam Bhuiyan, Ayumu Kubota, Shinsaku Kiyomoto, and Kazumasa Omote. Optimizing share size in efficient and robust secret sharing scheme for big data. IEEE Transactions on Big Data, 7(4):703–716, April 2021. CODEN ???? ISSN 2332-7790.

Tamburri:2023:BDS

[TvMvdH23] Damian A. Tamburri, Vincent R. van Mierlo, and Willem-Jan van den Heuvel. Big data for the social good: The drought early-warning experience report. IEEE Transactions on Big Data, 9(3):773– 791, June 2023. ISSN 2332-7790.

Tong:2020:EDF

[TXH^{+}20] Xin-Yi Tong, Gui-Song Xia, Fan Hu, Yanfei Zhong, Mihai Datcu, and Liangpei Zhang. Exploiting deep features for remote sensing image retrieval: a systematic investigation. IEEE Transactions on Big Data, 6(3): 507–521, September 2020. CO-DEN ???? ISSN 2332-7790.

Teng:2022:LSA

[TZCC22] Dan Teng, Xiaowei Zhang, Li Cheng, and Delin Chu. Least squares approximation via sparse subsampled randomized Hadamard transform. IEEE Transactions on Big Data, 8(2):446–457, February 2022. CODEN ???? ISSN 2332-7790.

Vhaduri:2021:ODP

[VP21] Sudip Vhaduri and Christian Poellabauer. Opportunistic discovery of personal places using multi-source sensor data. IEEE Transactions on Big Data, 7(2): 383–396, June 2021. CODEN ???? ISSN 2332-7790.

Wang:2016:PCD

[Wan16] Jiajun Wang. Partial copy detection in videos: A benchmark and an evaluation of popular methods. IEEE Transactions on Big Data, 2(1):32–42, March 2016. CODEN ???? ISSN 2332-7790.

Wang:2020:BDP

[WBD^{+}20] Shuang Wang, Luca Bonomi, Wenrui Dai, Feng Chen, Cynthia Cheung, Cinnamon S. Bloss, Samuel Cheng, and Xiaoqian Jiang. Big data privacy in biomedical research. IEEE Transactions on Big Data, 6(2): 296–308, June 2020. CODEN ???? ISSN 2332-7790.

Wang:2023:SHG

[WBS^{+}23] Xiao Wang, Deyu Bo, Chuan Shi, Shaohua Fan, Yanfang Ye, and Philip S. Yu. A survey on heterogeneous graph embedding: Methods, techniques, applications and sources. IEEE Transactions on Big Data, 9 (2):415–436, April 2023. ISSN 2332-7790.

Wu:2022:SEP

[WCJI22] Sijie Wu, Hanhua Chen, Hai Jin, and Shadi Ibrahim. Shadow: Exploiting the power of choice for efficient shuffling in MapReduce. IEEE Transactions on Big Data, 8(1):253– 267, February 2022. CODEN ???? ISSN 2332-7790.

Wang:2023:AGB

[WCS^{+}23] Zhenya Wang, Xiang Cheng, Sen Su, Jintao Liang, and Haocheng Yang. ATLAS: GAN-Based differentially private multi-party data sharing. IEEE Transactions on Big Data, 9(4):1225–1237, August 2023. ISSN 2332-7790.

Wang:2022:IIR

[WCW^{+}22] Jia Wang, Jiannong Cao, Senzhang Wang, Zhongyu Yao, and Wengen Li. IRDA: Incremental reinforcement learning for dynamic resource allocation. IEEE Transactions on Big Data, 8(3):770–783, June 2022. ISSN 2332-7790.

Wen:2021:FPD

[WCYL21] Fei Wen, Lei Chu, Rendong Ying, and Peilin Liu. Fast and positive definite estimation of large covariance matrix for high-dimensional data analysis. IEEE Transactions on Big Data, 7(3):603–609, July 2021. CODEN ???? ISSN 2332-7790.

Wang:2023:RMD

[WFZ^{+}23] Chenxu Wang, Fuli Feng, Yang Zhang, Qifan Wang, Xunhan Hu, and Xiangnan He. Rethinking missing data: Aleatoric uncertainty-aware recommendation. IEEE Transactions on Big Data, 9(6):1607–1619, December 2023. ISSN 2332-7790.

Wang:2023:WBA

[WGG^{+}23] Shiqi Wang, Chongming Gao, Min Gao, Junliang Yu, Zongwei Wang, and Hongzhi Yin. Who are the best adopters? User selection model for free trial item promotion. IEEE Transactions on Big Data, 9(2):746– 757, April 2023. ISSN 2332-7790.

Wang:2023:EHG

[WGYW23] Junfu Wang, Yuanfang Guo, Liang Yang, and Yunhong Wang. Enabling homogeneous GNNs to handle heterogeneous graphs via relation embedding. IEEE Transactions on Big Data, 9(6):1697–1710, December 2023. ISSN 2332-7790.

Wan:2019:ALA

[WH19] Zhiqiang Wan and Haibo He. AnswerNet: Learning to answer questions. IEEE Transactions on Big Data, 5(4):540–549, December 2019. CODEN ???? ISSN 2332-7790.

Wang:2023:CRC

[WHGW23] Shuai Wang, Shijie Hu, Baoshen Guo, and Guang Wang. Crossregion courier displacement for on-demand delivery with multiagent reinforcement learning. IEEE Transactions on Big Data, 9(5):1321–1333, October 2023. ISSN 2332-7790.

Wan:2018:GMS

[WHT18] Z. Wan, H. He, and B. Tang.

A generative model for sparse hyperparameter determination. IEEE Transactions on Big Data, 4(1):2–10, March 2018. CODEN ???? ISSN 2332-7790.

Wang:2023:OPP

[WHX^{+}23] Zhuzhu Wang, Cui Hu, Bin Xiao, Yang Liu, Teng Li, Zhuo Ma, and Jianfeng Ma. Outsourced privacy-preserving data alignment on vertically partitioned database. IEEE Transactions on Big Data, 9(5):1408– 1419, October 2023. ISSN 2332-7790.

Wu:2016:TRT

[WJS^{+}16] L. Wu, K. John Wu, A. Sim, M. Churchill, J. Y. Choi, A. Stathopoulos, C. Chang, and S. Klasky. Towards realtime detection and tracking of spatio-temporal features: Blobfilaments in fusion plasma. IEEE Transactions on Big Data, 2(3):262–275, September 2016. CODEN ???? ISSN 2332-7790.

Wang:2018:AMC

[WJW^{+}18] B. Wang, J. Jiang, Y. Wu, G. Yang, and K. Li. Accelerating MapReduce on commodity clusters: An SSD-empowered approach. IEEE Transactions on Big Data, 4(3):396–407, September 2018. CODEN ???? ISSN 2332-7790.

Wang:2024:PSA

[WJW24] Zihao Wang, Ming Jiang, and Junli Wang. PHAED: a speaker-aware parallel hierarchical attentive encoderdecoder model for multi-turn dialogue generation. IEEE Transactions on Big Data, 10 (1):23–34, February 2024. ISSN 2332-7790.

Wang:2022:SAL

[WL22a] Li Wang and Ren-Cang Li. A scalable algorithm for largescale unsupervised multi-view partial least squares. IEEE Transactions on Big Data, 8(4): 1073–1083, August 2022. ISSN 2332-7790.

Wang:2022:DDB

[WL22b] Shengnan Wang and Chunguang Li. Discrete double-bit hashing. IEEE Transactions on Big Data, 8(2):482–494, February 2022. CODEN ???? ISSN 2332-7790.

Wang:2023:GDL

[WLD^{+}23] Dexian Wang, Tianrui Li, Ping Deng, Jia Liu, Wei Huang, and Fan Zhang. A generalized deep learning algorithm based on NMF for multi-view clustering. IEEE Transactions on Big Data, 9(1):328–340, February 2023. ISSN 2332-7790.

Wang:2022:CTD

[WLF^{+}22] Tian Wang, Yang Li, Weiwei Fang, Wenzheng Xu, Junbin Liang, Yewang Chen, and Xuxun Liu. A comprehensive trustworthy data collection approach in sensor-cloud systems. IEEE Transactions on Big Data, 8(1):140–151, February 2022. CODEN ???? ISSN 2332-7790.

Wang:2023:OPR

[WLXF23] Dongjie Wang, Kunpeng Liu, Hui Xiong, and Yanjie Fu. Online POI recommendation: Learning dynamic geo-human interactions in streams. IEEE Transactions on Big Data, 9(3): 832–844, June 2023. ISSN 2332-7790.

Wang:2020:PSO

[WLZ^{+}20] Guifeng Wang, Qi Liu, Hongke Zhao, Chuanren Liu, Tong Xu, and Enhong Chen. Product supply optimization for crowdfunding campaigns. IEEE Transactions on Big Data, 6(4): 741–756, December 2020. CO-DEN ???? ISSN 2332-7790.

Wu:2018:BDA

[WOD^{+}18] J. Wu, K. Ota, M. Dong, J. Li, and H. Wang. Big data analysis-based security situational awareness for smart grid. IEEE Transactions on Big Data, 4(3):408–417, September 2018. CODEN ???? ISSN 2332-7790.

Wang:2017:GEB

[WQSA17] J. Wang, G. Qi, N. Sebe, and C. Aggarwal. Guest editorial: Big media data: Understanding, search, and mining. IEEE Transactions on Big Data, 3(1):

36, March 2017. CODEN ???? ISSN 2332-7790.

Wu:2017:EED

[WS17] B. Wu and H. Shen. Exploiting efficient densest subgraph discovering methods for big data. IEEE Transactions on Big Data, 3(3):334– 348, September 2017. CODEN ???? ISSN 2332-7790.

Wang:2018:MEM

[WSW^{+}18] Y. Wang, Y. Shen, H. Wang, J. Cao, and X. Jiang. MtMR: Ensuring MapReduce computation integrity with Merkle tree-based verifications. IEEE Transactions on Big Data, 4(3): 418–431, September 2018. CO-DEN ???? ISSN 2332-7790.

Win:2018:BDB

[WTM18] T. Y. Win, H. Tianfield, and Q. Mair. Big data based security analytics for protecting virtualized infrastructures in cloud computing. IEEE Transactions on Big Data, 4(1):11–25, March 2018. CODEN ???? ISSN 2332-7790.

Wang:2016:KSP

[WTRK16] L. Wang, S. Tasoulis, T. Roos, and J. Kangasharju. Kvasir: Scalable provision of semantically relevant Web content on big data framework. IEEE Transactions on Big Data, 2(3): 219–233, September 2016. CO-DEN ???? ISSN 2332-7790.

Wang:2022:CFN

[WTX^{+}22] Wei Wang, Tao Tang, Feng Xia, Zhiguo Gong, Zhikui Chen, and Huan Liu. Collaborative filtering with network representation learning for citation recommendation. IEEE Transactions on Big Data, 8(5):1233–1246, October 2022. ISSN 2332-7790.

Wanyan:2021:RLI

[WVD^{+}21] Tingyi Wanyan, Akhil Vaid, Jessica K. De Freitas, Sulaiman Somani, Riccardo Miotto, Girish N. Nadkarni, Ariful Azad, Ying Ding, and Benjamin S. Glicksberg. Relational learning improves prediction of mortality in COVID-19 in the intensive care unit. IEEE Transactions on Big Data, 7(1): 38–44, March 2021. CODEN ???? ISSN 2332-7790.

Wang:2021:CEH

[WW21] Ning Wang and Jie Wu. Costefficient heterogeneous worker recruitment under coverage requirement in spatial crowdsourcing. IEEE Transactions on Big Data, 7(2):407–420, June 2021. CODEN ???? ISSN 2332-7790.

Wang:2023:PEO

[WWC^{+}23] Chao Wang, Shuo Wang, Xiaoman Cheng, Yunhua He, Ke Xiao, and Shujia Fan. A privacy and efficiency-oriented data sharing mechanism for IoTs. IEEE Transactions on Big Data, 9(1):174–185, February 2023. ISSN 2332-7790.

Wang:2022:MRM

[WWCK22] Rui Wang, Xiao-Jun Wu, Kai-Xuan Chen, and Josef Kittler. Multiple Riemannian manifoldvalued descriptors based image set classification with multikernel metric learning. IEEE Transactions on Big Data, 8(3): 753–769, June 2022. ISSN 2332-7790.

Wong:2022:DTC

[WWCZ22] Hok Shing Wong, Li Wang, Raymond Chan, and Tieyong Zeng. Deep tensor CCA for multi-view learning. IEEE Transactions on Big Data, 8 (6):1664–1677, December 2022. ISSN 2332-7790.

Wu:2021:GTB

[WWK^{+}21] Xiaotong Wu, Taotao Wu, Maqbool Khan, Qiang Ni, and Wanchun Dou. Game theory based correlated privacy preserving analysis in big data. IEEE Transactions on Big Data, 7(4):643–656, April 2021. CODEN ???? ISSN 2332-7790.

Wang:2023:EMD

[WWL^{+}23] Zhu Wang, Zilong Wang, Xiaona Li, Zhiwen Yu, Bin Guo, Liming Chen, and Xingshe Zhou. Exploring multidimension user-item interactions with attentional knowledge graph neural networks for recommendation. IEEE Transactions on Big Data, 9(1):212– 226, February 2023. ISSN 2332-

West:2016:RSB

[WWSB16] Jevin D. West, Ian Wesley-Smith, and Carl T. Bergstrom. A recommendation system based on hierarchical clustering of an article-level citation network. IEEE Transactions on Big Data, 2(2):113–123, June 2016. CODEN ???? ISSN 2332-7790.

Wang:2024:AAM

[WWYH24] Lei Wang, Leon Wong, Zhu-Hong You, and De-Shuang Huang. AMDECDA: Attention mechanism combined with data ensemble strategy for predicting CircRNA-disease association. IEEE Transactions on Big Data, 10(4):320–329, August 2024. ISSN 2332-7790.

Wang:2022:MSC

[WXW^{+}22] Xinying Wang, Cong Xu, Ke Wang, Feng Yan, and Dongfang Zhao. Memory scaling of cloud-based big data systems: a hybrid approach. IEEE Transactions on Big Data, 8(5):1259– 1272, October 2022. ISSN 2332-7790.

Wang:2019:RBD

[WXZ^{+}19] K. Wang, C. Xu, Y. Zhang, S. Guo, and A. Y. Zomaya. Robust big data analytics for electricity price forecasting in the smart grid. IEEE Transactions on Big Data, 5(1):34–45, March 2019. CODEN ???? ISSN 2332-7790.

Wang:2024:SLC

[WYBH24] Meng Wang, Yanhao Yang, David Bindel, and Kun He. Streaming local community detection through approximate conductance. IEEE Transactions on Big Data, 10(1):12–22, February 2024. ISSN 2332-7790.

Wang:2020:MOD

[WYC^{+}20] Xiaokang Wang, Laurence T. Yang, Xingyu Chen, Lizhe Wang, Rajiv Ranjan, Xiaodao Chen, and M. Jamal Deen. A multi-order distributed HOSVD with its incremental computing for big services in cyberphysical-social systems. IEEE Transactions on Big Data, 6(4): 666–678, December 2020. CO-DEN ???? ISSN 2332-7790.

Wang:2019:NSM

[WYK^{+}19] H. Wang, H. Yao, D. Kifer, C. Graif, and Z. Li. Nonstationary model for crime rate inference using modern urban data. IEEE Transactions on Big Data, 5(2):180–194, June 2019. CODEN ???? ISSN 2332-7790.

Wang:2018:BDS

[WYLD18] X. Wang, L. T. Yang, H. Liu, and M. J. Deen. A big data-as-a-service framework: State-of-the-art and perspectives. IEEE Transactions on Big Data, 4(3): 325–340, September 2018. CO-DEN ???? ISSN 2332-7790.

Wang:2021:PAA

[WYLG21] Kun Wang, Jiahui Yu, Xiulong Liu, and Song Guo. A pre-authentication approach to proxy re-encryption in big data context. IEEE Transactions on Big Data, 7(4):657–667, April 2021. CODEN ???? ISSN 2332-7790.

Wan:2021:YSI

[WYXL21] Liangtian Wan, Yuyuan Yuan, Feng Xia, and Huan Liu. To your surprise: Identifying serendipitous collaborators. IEEE Transactions on Big Data, 7(3):574–589, July 2021. CODEN ???? ISSN 2332-7790.

Wang:2023:LFP

[WZH^{+}23] Cheng Wang, Hangyu Zhu, Ruixin Hu, Rui Li, and Changjun Jiang. LongArms: Fraud prediction in online lending services using sparse knowledge graph. IEEE Transactions on Big Data, 9(2):758– 772, April 2023. ISSN 2332-7790.

Wang:2024:DHS

[WZL^{+}24a] Shun Wang, Yong Zhang, Xuanqi Lin, Yongli Hu, Qingming Huang, and Baocai Yin. Dynamic hypergraph structure learning for multivariate time series forecasting. IEEE Transactions on Big Data, 10(4):556– 567, August 2024. ISSN 2332-7790.

Wu:2024:MAN

[WZL^{+}24b] Han Wu, Guanqi Zhu, Qi Liu, Hengshu Zhu, Hao Wang, Hongke Zhao, Chuanren Liu, Enhong Chen, and Hui Xiong. A multi-aspect neural tensor factorization framework for patent litigation prediction. IEEE Transactions on Big Data, 10(1):35–54, February 2024. ISSN 2332-7790.

Wu:2018:HCF

[WZLS18a] D. Wu, L. Zhu, Q. Lu, and S. Sakr. HDM: A composable framework for big data processing. IEEE Transactions on Big Data, 4(2):150–163, June 2018. CODEN ???? ISSN 2332-7790.

Wu:2018:ECL

[WZLS18b] J. Wu, W. Zheng, J. Lai, and C. Y. Suen. Euler clustering on large-scale dataset. IEEE Transactions on Big Data, 4(4): 502–515, December 2018. CO-DEN ???? ISSN 2332-7790.

Wang:2018:SBD

[WZY^{+}18] J. Wang, X. Zhang, J. Yin, R. Wang, H. Wu, and D. Han. Speed up big data analytics by unveiling the storage distribution of sub-datasets. IEEE Transactions on Big Data, 4(2): 231–244, June 2018. CODEN ???? ISSN 2332-7790.

Wang:2019:NRS

[WZZ^{+}19] Binfeng Wang, Jun Zhang, Zili Zhang, Lei Pan, Yang Xiang, and Dawen Xia. Noiseresistant statistical traffic classification. IEEE Transactions on Big Data, 5(4):454–466, December 2019. CODEN ???? ISSN 2332-7790.

Xiong:2022:RIB

[XCV22] Hu Xiong, Kim-Kwang Raymond Choo, and Athanasios V. Vasilakos. Revocable identitybased access control for big data with verifiable outsourced computing. IEEE Transactions on Big Data, 8(1):1–13, February 2022. CODEN ???? ISSN 2332-7790.

Xu:2023:MHI

[XDH23] Kejie Xu, Peifang Deng, and Hong Huang. Mining hierarchical information of CNNs for scene classification of VHR remote sensing images. IEEE Transactions on Big Data, 9 (2):542–554, April 2023. ISSN 2332-7790.

Xu:2020:BPC

[XDZ20] Yonghao Xu, Bo Du, and Liangpei Zhang. Beyond the patchwise classification: Spectralspatial fully convolutional networks for hyperspectral image classification. IEEE Transactions on Big Data, 6(3):492– 506, September 2020. CODEN ???? ISSN 2332-7790.

Xue:2021:REB

[XHZW21] Xiao Xue, Shuai Huangfu, Lejun Zhang, and Shufang Wang. Research on escaping the Big-Data traps in O2O service recommendation strategy. IEEE Transactions on Big Data, 7(1):199–213, March 2021. CODEN ???? ISSN 2332-7790.

Xu:2021:PAT

[XJQ^{+}21] Lei Xu, Chunxiao Jiang, Yi Qian, Jianhua Li, Youjian Zhao, and Yong Ren. Privacy-accuracy trade-off in differentially-private distributed classification: a game theoretical approach. IEEE Transactions on Big Data, 7(4):770– 783, April 2021. CODEN ???? ISSN 2332-7790.

Xu:2020:OTT

[XLC20] Ran Xu, Yong Li, and Sheng Chen. On the opportunistic topology of taxi networks in urban mobility environment. IEEE Transactions on Big Data, 6(1):171–188, March 2020. CODEN ???? ISSN 2332-7790.

Xu:2022:RLP

[XLC22] Xiangxi Xu, Zhihui Lai, and Yudong Chen. Relaxed locality preserving supervised discrete hashing. IEEE Transactions on Big Data, 8(4):1118– 1128, August 2022. ISSN 2332-7790.

Xu:2018:LTO

[XLL^{+}18] Z. Xu, X. Luo, Y. Liu, K. R. Choo, V. Sugumaran, N. Yen, L. Mei, and C. Hu. From latency, through outbreak, to decline: Detecting different states of emergency events using Web resources. IEEE Transactions on Big Data, 4(2):245–257, June 2018. CODEN ???? ISSN 2332-7790.

Xie:2022:IAD

[XLL^{+}22] Ruobing Xie, Qi Liu, Shukai Liu, Ziwei Zhang, Peng Cui, Bo Zhang, and Leyu Lin. Improving accuracy and diversity in matching of recommendation with diversified preference network. IEEE Transactions on Big Data, 8(4):955–967, August 2022. ISSN 2332-7790.

Xia:2016:SAR

[XLLC16] Feng Xia, Haifeng Liu, Ivan Lee, and Longbing Cao. Scientific article recommendation: Exploiting common author relations and historical preferences. IEEE Transactions on Big Data, 2(2):101–112, June 2016. CODEN ???? ISSN 2332-7790.

Xue:2018:SNN

[XLP^{+}18] W. Xue, H. Li, Y. Peng, J. Cui, and Y. Shi. Secure k nearest neighbors query for high-dimensional vectors in outsourced environments. IEEE Transactions on Big Data, 4(4): 586–599, December 2018. CO-DEN ???? ISSN 2332-7790.

Xie:2022:ASS

[XLQ^{+}22] Yu Xie, Shengze Lv, Yuhua Qian, Chao Wen, and Jiye Liang. Active and semisupervised graph neural networks for graph classification. IEEE Transactions on Big Data, 8(4):920–932, August 2022. ISSN 2332-7790.

Xu:2019:MMD

[XML^{+}19] Zheng Xu, Lin Mei, Zhihan Lv, Chuanping Hu, Xiangfeng Luo, Hui Zhang, and Yunhuai Liu. Multi-modal description of public safety events using surveillance and social media. IEEE Transactions on Big Data, 5(4): 529–539, December 2019. CO-DEN ???? ISSN 2332-7790.

Xu:2023:NDF

[XPC^{+}23] Weihua Xu, Yanzhou Pan, Xiuwei Chen, Weiping Ding, and Yuhua Qian. A novel dynamic fusion approach using information entropy for interval-valued ordered datasets. IEEE Transactions on Big Data, 9(3):845– 859, June 2023. ISSN 2332-7790.

Xu:2022:TBI

[XSYW22] Qichao Xu, Zhou Su, Shui Yu, and Ying Wang. Trust based incentive scheme to allocate big data tasks with mobile social cloud. IEEE Transactions on Big Data, 8(1):113–124, February 2022. CODEN ???? ISSN 2332-7790.

Xia:2017:BSD

[XWBL17] F. Xia, W. Wang, T. M. Bekele, and H. Liu. Big scholarly data: A survey. IEEE Transactions on Big Data, 3(1):18–35, March 2017. CODEN ???? ISSN 2332-7790.

Xiao:2023:SSG

[XWG23a] Shunxin Xiao, Shiping Wang, and Wenzhong Guo. SGAE:

Stacked graph autoencoder for deep clustering. IEEE Transactions on Big Data, 9(1):254– 266, February 2023. ISSN 2332-7790.

Xu:2023:STC

[XWG^{+}23b] Xovee Xu, Zhiyuan Wang, Qiang Gao, Ting Zhong, Bei Hui, Fan Zhou, and Goce Trajcevski. Spatial-temporal contrasting for fine-grained urban flow inference. IEEE Transactions on Big Data, 9(6):1711– 1725, December 2023. ISSN 2332-7790.

Xiao:2021:APL

[XWI^{+}21] Cao Xiao, Shouyi Wang, Leon Iasemidis, Stephen Wong, and Wanpracha Art Chaovalitwongse. An adaptive pattern learning framework to personalize online seizure prediction. IEEE Transactions on Big Data, 7(5):819–831, November 2021. CODEN ???? ISSN 2332-7790.

Xu:2023:ESP

[XWZ^{+}23] Chang Xu, Ruijuan Wang, Liehuang Zhu, Chuan Zhang, Rongxing Lu, and Kashif Sharif. Efficient strong privacypreserving conjunctive keyword search over encrypted cloud data. IEEE Transactions on Big Data, 9(3):805–817, June 2023. ISSN 2332-7790.

Xu:2023:DUG

[XYX^{+}23] Chenchu Xu, Yuan Yang, Zhiqiang Xia, Boyan Wang, Dong Zhang, Yanping Zhang, and Shu Zhao. Dual uncertaintyguided mixing consistency for semi-supervised 3D medical image segmentation. IEEE Transactions on Big Data, 9(4):1156– 1170, August 2023. ISSN 2332-7790.

Xu:2021:ALR

[XZGW21] Junhong Xu, Shangyue Zhu, Hanqing Guo, and Shaoen Wu. Automated labeling for robotic autonomous navigation through multi-sensory semi-supervised learning on Big Data. IEEE Transactions on Big Data, 7(1): 93–101, March 2021. CODEN ???? ISSN 2332-7790.

Xue:2024:FTP

[XZJT24] Meiting Xue, Zian Zhou, Pengfei Jiao, and Huijun Tang. Fine-tuned personality federated learning for graph data. IEEE Transactions on Big Data, 10(3):313–319, June 2024. ISSN 2332-7790.

Yang:2022:WMA

[YAG^{+}22] Zhengyu Yang, Manu Awasthi, Mrinmoy Ghosh, Janki Bhimani, and Ningfang Mi. I/O workload management for allflash datacenter storage systems based on total cost of ownership. IEEE Transactions on Big Data, 8(2):332–345, February 2022. CODEN ???? ISSN 2332-7790.

Yang:2015:LAA

[Yan15a] Laurence T. Yang. LS-AMS: An adaptive indexing structure for realtime search on microblogs. IEEE Transactions on Big Data, 1(4):125–137, December 2015. CODEN ???? ISSN 2332-7790.

Yang:2015:IIT

[Yan15b]  Qiang Yang.  Introduction

to the IEEE Transactions on Big Data. IEEE Transactions on Big Data, 1(1):2– 15, March 2015. CODEN ???? ISSN 2332-7790.

URL http://www.computer.
org/csdl/trans/bd/2015/01/
07265156.pdf.

Yang:2017:SJE

[Yan17] Q. Yang. State of the journal editorial. IEEE Transactions on Big Data, 3(1):1, March 2017. CODEN ???? ISSN 2332-7790.

Yang:2018:SJE

[Yan18] Q. Yang. State of the journal editorial. IEEE Transactions on Big Data, 4(1):1, March 2018. CODEN ???? ISSN 2332-7790.

Yang:2019:SJ

[Yan19] Q. Yang. State of the journal. IEEE Transactions on Big Data, 5(1):1, March 2019. CO-DEN ???? ISSN 2332-7790.

Yang:2020:MOE

[Yan20] Qiang Yang. Message from the outgoing Editor-in-Chief. IEEE Transactions on Big Data, 6(1): 1, March 2020. CODEN ???? ISSN 2332-7790.

Yang:2023:APS

[Yan23] Zhuang Yang. Adaptive powerball stochastic conjugate gradient for large-scale learning. IEEE Transactions on Big Data, 9(6):1598–1606, December 2023. ISSN 2332-7790.

Yao:2015:GEB

[Yao15]  Lina Yao.  Guest edito-

rial: Big data analytics and the Web. IEEE Transactions on Big Data, 1(4):123– 124, December 2015. CO-DEN ???? ISSN 2332-7790.

URL http://www.computer.
org/csdl/trans/bd/2015/04/
07395032.pdf.

Yang:2024:LCC

[YCYL24] Shujing Yang, Fuyuan Cao, Kui Yu, and Jiye Liang. Learning causal chain graph structure via alternate learning and double pruning. IEEE Transactions on Big Data, 10(4):442–456, August 2024. ISSN 2332-7790.

Yi:2022:PFG

[YDL^{+}22] Xiuwen Yi, Zhewen Duan, Ruiyuan Li, Junbo Zhang, Tianrui Li, and Yu Zheng. Predicting fine-grained air quality based on deep neural networks. IEEE Transactions on Big Data, 8(5):1326–1339, October 2022. ISSN 2332-7790.

Yang:2023:ADC

[YDM^{+}23] Chen Yang, Zhihui Du, Xiaofeng Meng, Xukang Zhang, Xinli Hao, and David A. Bader. Anomaly detection in catalog streams. IEEE Transactions on Big Data, 9(1):294–311, February 2023. ISSN 2332-7790.

Yan:2016:DEB

[YDY^{+}16] Zheng Yan, Wenxiu Ding, Xixun Yu, Haiqi Zhu, and Robert H. Deng. Deduplication on encrypted big data in cloud. IEEE Transactions on Big Data, 2(2):138–150, June 2016. CODEN ???? ISSN 2332-7790.

Yuan:2022:MRL

[YHLS22] Ye Yuan, Qiang He, Xin Luo, and Mingsheng Shang. A multilayered-and-randomized latent factor model for highdimensional and sparse matrices. IEEE Transactions on Big Data, 8(3):784–794, June 2022. ISSN 2332-7790.

Yazdanparast:2021:LTC

[YJH21] Sakineh Yazdanparast, Mohsen Jamalabdollahi, and Timothy C. Havens. Linear time community detection by a novel modularity gain acceleration in label propagation. IEEE Transactions on Big Data, 7(6):961– 966, December 2021. CODEN ???? ISSN 2332-7790.

Yao:2023:MVM

[YKL^{+}23] Zijun Yao, Deguang Kong, Miao Lu, Xiao Bai, Jian Yang, and Hui Xiong. Multi-view multi-task campaign embedding for cold-start conversion rate forecasting. IEEE Transactions on Big Data, 9(1):280– 293, February 2023. ISSN 2332-7790.

Yu:2019:RRT

[YLB^{+}19] Weiren Yu, Jianxin Li, Md Zakirul Alam Bhuiyan, Richong Zhang, and Jinpeng Huai. Ring: Real-time emerging anomaly monitoring system over text streams. IEEE Transactions on Big Data, 5(4):506–519, December 2019. CODEN ???? ISSN 2332-7790.

Yang:2022:AEP

[YLC^{+}22] Xue Yang, Rongxing Lu, Kim Kwang Raymond Choo, Fan Yin, and Xiaohu Tang. Achieving efficient and privacypreserving cross-domain big data deduplication in cloud. IEEE Transactions on Big Data, 8(1):73–84, February 2022. CODEN ???? ISSN 2332-7790.

Yu:2022:MAP

[YLL22] Yangwen Yu, Victor O. K. Li, and Jacqueline C. K. Lam. Missing air pollution data recovery based on long-short term context encoder. IEEE Transactions on Big Data, 8(3):711– 722, June 2022. ISSN 2332-7790.

Yu:2023:HRM

[YLL23] Yangwen Yu, Victor O. K. Li, and Jacqueline C. K. Lam. Hierarchical recovery of missing air pollution data via improved long-short term context encoder network. IEEE Transactions on Big Data, 9(1):93–105, February 2023. ISSN 2332-7790.

Yao:2018:HAR

[YLLC18] Y. Yao, Y. Liu, Z. Liu, and H. Chen. Human activity recognition with posture tendency descriptors on action snippets. IEEE Transactions on Big Data, 4(4):530–541, December 2018. CODEN ???? ISSN 2332-7790.

Yu:2023:GSM

[YLLC23] Yangwen Yu, Victor O. K. Li, Jacqueline C. K. Lam, and Kelvin Chan. GCN-ST-MDIR: Graph convolutional networkbased spatial-temporal missing air pollution data pattern identification and recovery. IEEE Transactions on Big Data, 9(5): 1347–1364, October 2023. ISSN 2332-7790.

Yang:2020:SIW

[YLW^{+}20] Yang Yang, Jie Li, Cheng-Xiang Wang, Olav Tirkkonen, and Ming-Tuo Zhou. Special issue on wireless Big Data. IEEE Transactions on Big Data, 6(2): 209–210, June 2020. CODEN ???? ISSN 2332-7790.

Yu:2019:BRS

[YLZ^{+}19] Shengkang Yu, Xi Li, Xueyi Zhao, Zhongfei Zhang, Fei Wu, Jingdong Wang, Yueting Zhuang, and Xuelong Li. A bilinear ranking SVM for knowledge based relation prediction and classification. IEEE Transactions on Big Data, 5(4):588– 600, December 2019. CODEN ???? ISSN 2332-7790.

Yao:2022:IET

[YLZ^{+}22] Haipeng Yao, Chong Liu, Peiying Zhang, Sheng Wu, Chunxiao Jiang, and Shui Yu. Identification of encrypted traffic through attention mechanism based long short term memory. IEEE Transactions on Big Data, 8(1):241–252, February 2022. CODEN ???? ISSN 2332-7790.

Yang:2024:GCL

[YM24] Yifei Yang and Xiaoke Ma. Graph contrastive learning for clustering of multi-layer networks. IEEE Transactions on Big Data, 10(4):429–441, August 2024. ISSN 2332-7790.

Yi:2023:TLT

[YMJ^{+}23] Si-Yu Yi, Zhengyang Mao, Wei Ju, Yong-Dao Zhou, Luchen Liu, Xiao Luo, and Ming Zhang. Towards long-tailed recognition for graph classification via collaborative experts. IEEE Transactions on Big Data, 9(6):1683–1696, December 2023. ISSN 2332-7790.

Yu:2024:BEI

[YRL^{+}24] Changyong Yu, Tianmei Ren, Wenyu Li, Huimin Liu, Haitao Ma, and Yuhai Zhao. BL: an efficient index for reachability queries on large graphs. IEEE Transactions on Big Data, 10 (2):108–121, April 2024. ISSN 2332-7790.

Yang:2023:MAM

[YSM^{+}23] Jun Yang, Yaoru Sun, Maoyu Mao, Lizhi Bai, Siyu Zhang, and Fang Wang. Modelagnostic method: Exposing deepfake using pixel-wise spatial and temporal fingerprints. IEEE Transactions on Big Data, 9(6):1496–1509, December 2023. ISSN 2332-7790.

Yu:2021:CTV

[YTY^{+}21] Yanwei Yu, Xianfeng Tang, Huaxiu Yao, Xiuwen Yi, and Zhenhui Li. Citywide traffic volume inference with surveillance camera records. IEEE Transactions on Big Data, 7(6):900– 912, December 2021. CODEN ???? ISSN 2332-7790.

Yu:2015:PNP

[Yu15] Yaoliang Yu. Petuum: A new platform for distributed machine learning on big data. IEEE Transactions on Big Data, 1(2):49–67, June 2015. CODEN ???? ISSN 2332-7790.

Yan:2021:CCC

[YWG^{+}21] Qingsen Yan, Bo Wang, Dong Gong, Chuan Luo, Wei Zhao, Jianhu Shen, Jingyang Ai, Qinfeng Shi, Yanning Zhang, Shuo Jin, Liang Zhang, and Zheng You. COVID-19 chest CT image segmentation network by multi-scale fusion and enhancement operations. IEEE Transactions on Big Data, 7(1):13– 24, March 2021. CODEN ???? ISSN 2332-7790.

Yu:2022:ETM

[YWL^{+}22] Jiahui Yu, Kun Wang, Peng Li, Rui Xia, Song Guo, and Minyi Guo. Efficient trustworthiness management for malicious user detection in big data collection. IEEE Transactions on Big Data, 8(1):99–112, February 2022. CODEN ???? ISSN 2332-7790.

Yue:2021:FLD

[YWM^{+}21] Chaoqun Yue, Shweta Ware, Reynaldo Morillo, Jin Lu, Chao Shang, Jinbo Bi, Jayesh Kamath, Alexander Russell, Athanasios Bamis, and Bing Wang. Fusing location data for depression prediction. IEEE Transactions on Big Data, 7(2): 355–370, June 2021. CODEN ???? ISSN 2332-7790.

Yang:2021:SID

[YWRY21] Xinyu Yang, Teng Wang, Xuebin Ren, and Wei Yu. Survey on improving data utility in differentially private sequential data publishing. IEEE Transactions on Big Data, 7(4):729– 749, April 2021. CODEN ???? ISSN 2332-7790.

Yan:2021:NNS

[YWW^{+}21a] Donghui Yan, Yingjie Wang, Jin Wang, Honggang Wang, and Zhenpeng Li. K-nearest neighbor search by random projection forests. IEEE Transactions on Big Data, 7(1):147– 157, March 2021. CODEN ???? ISSN 2332-7790.

Yan:2021:FCE

[YWW^{+}21b] Donghui Yan, Yingjie Wang, Jin Wang, Guodong Wu, and Honggang Wang. Fast communication-efficient spectral clustering over distributed data. IEEE Transactions on Big Data, 7(1):158–168, March 2021. CODEN ???? ISSN 2332-7790.

Yang:2022:TEL

[YWY^{+}22] Shuai Yang, Hao Wang, Kui Yu, Fuyuan Cao, and Xindong Wu. Towards efficient local causal structure learning. IEEE Transactions on Big Data, 8(6):1592– 1609, December 2022. ISSN 2332-7790.

Yue:2023:AGA

[YXZL23] Guowei Yue, Rui Xiao, Zhongying Zhao, and Chao Li. AF-GCN: Attribute-fusing graph convolution network for recommendation. IEEE Transactions on Big Data, 9(2):597– 607, April 2023. ISSN 2332-7790.

Yang:2023:BDF

[YYD^{+}23] Jia-Quan Yang, Qi-Te Yang, Ke-Jing Du, Chun-Hua Chen, Hua Wang, Sang-Woon Jeon, Jun Zhang, and Zhi-Hui Zhan. Bi-directional feature fixationbased particle swarm optimization for large-scale feature selection. IEEE Transactions on Big Data, 9(3):1004–1017, June 2023. ISSN 2332-7790.

Yan:2019:HDS

[YZDZ19] Z. Yan, L. Zhang, W. DING, and Q. Zheng. Heterogeneous data storage management with deduplication in cloud computing. IEEE Transactions on Big Data, 5(3):393–407, September 2019. CODEN ???? ISSN 2332-7790.

Yang:2023:SSF

[YZW^{+}23a] Qiming Yang, Qi Zhu, Mingming Wang, Wei Shao, Zheng Zhang, and Daoqiang Zhang. Self-supervised federated adaptation for multi-site brain disease diagnosis. IEEE Transactions on Big Data, 9(5):1334– 1346, October 2023. ISSN 2332-7790.

Yuan:2023:SST

[YZW^{+}23b] Ye Yuan, Yong Zhang, Boyue Wang, Yuan Peng, Yongli Hu, and Baocai Yin. STGAN: Spatio-temporal generative adversarial network for traffic data imputation. IEEE Transactions on Big Data, 9(1):200– 211, February 2023. ISSN 2332-7790.

Zhang:2023:PUS

[ZAL^{+}23] Yupei Zhang, Rui An, Shuhui Liu, Jiaqi Cui, and Xuequn Shang. Predicting and understanding student learning performance using multi-source sparse attention convolutional neural networks. IEEE Transactions on Big Data, 9(1):118– 132, February 2023. ISSN 2332-7790.

Zinno:2020:NSS

[ZBB^{+}20] I. Zinno, M. Bonano, S. Buonanno, F. Casu, C. De Luca, M. Manunta, M. Manzo, and R. Lanari. National scale surface deformation time series generation through advanced DInSAR processing of Sentinel-1 data within a cloud computing environment. IEEE Transactions on Big Data, 6(3):558– 571, September 2020. CODEN ???? ISSN 2332-7790.

Zhang:2023:SDP

[ZBL23] Mengxiao Zhang, Fernando Beltrán, and Jiamou Liu. A survey of data pricing for data marketplaces. IEEE Transactions on Big Data, 9(4):1038– 1056, August 2023. ISSN 2332-7790.

Zhang:2020:KPA

[ZBN^{+}20] Zhao Zhang, Kyle Barbary, Frank Austin Nothaft, Evan R. Sparks, Oliver Zahn, Michael J. Franklin, David A. Patterson, and Saul Perlmutter. Kira: Processing astronomy imagery using Big Data technology. IEEE Transactions on Big Data, 6(2): 369–381, June 2020. CODEN ???? ISSN 2332-7790.

Zhao:2018:DFS

[ZCH^{+}18] L. Zhao, Z. Chen, Y. Hu, G. Min, and Z. Jiang. Distributed feature selection for efficient economic big data analysis. IEEE Transactions on Big Data, 4(2):164–176, June 2018. CODEN ???? ISSN 2332-7790.

Zhang:2023:TAU

[ZCSW23] Pengfei Zhang, Xiang Cheng, Sen Su, and Ning Wang. Task allocation under geoindistinguishability via groupbased noise addition. IEEE Transactions on Big Data, 9(3): 860–877, June 2023. ISSN 2332-7790.

Zhang:2023:PLD

[ZCSZ23] Pengfei Zhang, Xiang Cheng, Sen Su, and Binyuan Zhu. PrivTDSI: a local differentially private approach for truth discovery via sampling and inference. IEEE Transactions on Big Data, 9(2):471–484, April 2023. ISSN 2332-7790.

Zhang:2023:TTH

[ZDD^{+}23] Chunkai Zhang, Quanjian Dai, Zilin Du, Wensheng Gan, Jian Weng, and Philip S. Yu. TUSQ: Targeted high-utility sequence querying. IEEE Transactions on Big Data, 9(2):512–527, April 2023. ISSN 2332-7790.

Zhao:2023:SDI

[ZDZC23] Jiachen Zhao, Fang Deng, Jiaqi Zhu, and Jie Chen. Searching density-increasing path to local density peaks for unsupervised anomaly detection. IEEE Transactions on Big Data, 9(4): 1198–1209, August 2023. ISSN 2332-7790.

Zhang:2019:UUD

[ZFLC19] M. Zhang, H. Fu, Y. Li, and S. Chen. Understanding urban dynamics from massive mobile traffic data. IEEE Transactions on Big Data, 5(2):266– 278, June 2019. CODEN ???? ISSN 2332-7790.

Zhang:2023:LIF

[ZFQ^{+}23] Yushu Zhang, Zhibin Fu, Shuren Qi, Mingfu Xue, Zhongyun Hua, and Yong Xiang. Localization of inpainting forgery with feature enhancement network. IEEE Transactions on Big Data, 9(3):936– 948, June 2023. ISSN 2332-7790.

Zeng:2024:LTS

[ZGC^{+}24] Qingtian Zeng, Shuai Guo, Rui Cao, Ziqi Zhao, and Hua Duan. Legal transition sequence recognition of a bounded petri net using a gate recurrent unit. IEEE Transactions on Big Data, 10 (1):66–76, February 2024. ISSN 2332-7790.

Zhao:2022:DEB

[ZGH^{+}22] Feng Zhao, Xiangyu Gui, Yafan Huang, Hai Jin, and Laurence T. Yang. Dynamic entitybased named entity recognition under unconstrained tagging schemes. IEEE Transactions on Big Data, 8(4):1059– 1072, August 2022. ISSN 2332-7790.

Zhang:2022:MBD

[ZGS^{+}22] Yang Zhang, Hanqi Guo, Lanyu Shang, Dong Wang, and Tom Peterka. A multi-branch decoder network approach to adaptive temporal data selection and reconstruction for big scientific simulation data. IEEE Transactions on Big Data, 8 (6):1637–1649, December 2022. ISSN 2332-7790.

Zheng:2015:MCD

[Zhe15] Yu Zheng. Methodologies for cross-domain data fusion: An overview. IEEE Transactions on Big Data, 1(1):16–34, March 2015. CODEN ???? ISSN 2332-7790.

Zhu:2024:SHO

[ZHKL24] Yanhui Zhu, Fang Hu, Lei Hsin Kuo, and Jia Liu. SCOREH+: a high-order node proximity spectral clustering on ratios-of-eigenvectors algorithm for community detection. IEEE Transactions on Big Data, 10(3):301– 312, June 2024. ISSN 2332-7790.

Zhang:2017:TPD

[ZHL^{+}17] D. Zhang, T. He, S. Lin, S. Munir, and J. A. Stankovic. Taxi-passenger-demand modeling based on big data from a roving sensor network. IEEE Transactions on Big Data, 3(3): 362–374, September 2017. CO-DEN ???? ISSN 2332-7790.

Zhang:2024:BSI

[ZHL^{+}24] Xiao Zhang, Zhaoqian He, Jinhai Li, Changlin Mei, and Yanyan Yang. Bi-selection of instances and features based on neighborhood importance degree. IEEE Transactions on Big Data, 10(4):415–428, August 2024. ISSN 2332-7790.

Zhuang:2015:WSS

[Zhu15] Yueting Zhuang. Weakly semisupervised deep learning for multi-label image annotation. IEEE Transactions on Big Data, 1(3):109–122, September 2015. CODEN ???? ISSN 2332-7790.

Zhou:2021:ELS

[ZHW21] Wei Zhou, Jiankun Hu, and Song Wang. Enhanced localitysensitive hashing for fingerprint forensics over large multi-sensor databases. IEEE Transactions on Big Data, 7(4):759– 769, April 2021. CODEN ???? ISSN 2332-7790.

Zhang:2020:STR

[ZJG^{+}20] Pengcheng Zhang, Yangyang Jia, Jerry Gao, Wei Song, and Hareton Leung. Short-term rainfall forecasting using multilayer perceptron. IEEE Transactions on Big Data, 6(1):93– 106, March 2020. CODEN ???? ISSN 2332-7790.

Zhong:2022:MIS

[ZJSL22] Yurong Zhong, Long Jin, Mingsheng Shang, and Xin Luo. Momentum-incorporated symmetric non-negative latent factor models. IEEE Transactions on Big Data, 8(4):1096–1106, August 2022. ISSN 2332-7790.

Zhang:2019:KIL

[ZJZ^{+}19] Z. Zhang, L. Jia, M. Zhao, G. Liu, M. Wang, and S. Yan. Kernel-induced label propagation by mapping for semisupervised classification. IEEE Transactions on Big Data, 5(2): 148–165, June 2019. CODEN ???? ISSN 2332-7790.

Zhu:2017:SIB

[ZLC^{+}17] J. Zhu, A. Liu, M. Chen, T. Tasdizen, and H. Su. Special issue on biomedical big data: Understanding, learning and applications. IEEE Transactions on Big Data, 3(4):375–377, December 2017. CODEN ???? ISSN 2332-7790.

Zheng:2021:EGP

[ZLG^{+}21] Long Zheng, Xianliang Li, Xi Ge, Xiaofei Liao, Zhiyuan Shao, Hai Jin, and Qiang-Sheng Hua. Efficient graph processing with invalid update filtration. IEEE Transactions on Big Data, 7(3):590–602, July 2021. CODEN ???? ISSN 2332-7790.

Zhou:2020:STD

[ZLT20] Xiaoping Zhou, Xun Liang, and Zhi Tang. Scalable triangle discovery algorithm for large scalefree network with limited internal memory. IEEE Transactions on Big Data, 6(4):757– 769, December 2020. CODEN ???? ISSN 2332-7790.

Zhang:2022:UAA

[ZLY^{+}22] Mingyang Zhang, Tong Li, Yue Yu, Yong Li, Pan Hui, and Yu Zheng. Urban anomaly analytics: Description, detection, and prediction. IEEE Transactions on Big Data, 8(3):809– 826, June 2022. ISSN 2332-7790.

Zhang:2020:DMB

[ZLZ^{+}20] Wenlu Zhang, Rongjian Li, Tao Zeng, Qian Sun, Sudhir Kumar, Jieping Ye, and Shuiwang Ji. Deep model based transfer and multi-task learning for biological image analysis. IEEE Transactions on Big Data, 6(2):322– 333, June 2020. CODEN ???? ISSN 2332-7790.

Zhang:2022:DTE

[ZLZ^{+}22] Yingxue Zhang, Yanhua Li, Xun Zhou, Xiangnan Kong, and Jun Luo. Off-deployment traffic estimation: a traffic generative adversarial networks approach. IEEE Transactions on Big Data, 8(4):1084–1095, August 2022. ISSN 2332-7790.

Zhang:2022:CCG

[ZLZL22] Xin Zhang, Yanhua Li, Xun Zhou, and Jun Luo. cGAIL: Conditional generative adversarial imitation learning an application in taxi drivers strategy learning. IEEE Transactions on Big Data, 8(5):1288–1300, October 2022. ISSN 2332-7790.

Zurita-Milla:2020:ESO

[ZMGIVO20] Raul Zurita-Milla, Romulo Goncalves, Emma Izquierdo-Verdiguier, and Frank O. Ostermann. Exploring spring onset at continental scales: Mapping phenoregions and correlating temperature and satellite-based phenometrics. IEEE Transactions on Big Data, 6(3):583– 593, September 2020. CODEN ???? ISSN 2332-7790.

Zheng:2017:GEU

[ZMS17] Y. Zheng, C. Mascolo, and C. T. Silva. Guest editorial: Urban computing. IEEE Transactions on Big Data, 3(2):124– 125, June 2017. CODEN ???? ISSN 2332-7790.

Zheng:2021:VBS

[ZOLP21] Yan Zheng, Yi Ou, Alexander Lex, and Jeff M. Phillips. Visualization of big spatial data using coresets for kernel density estimates. IEEE Transactions on Big Data, 7(3):524–534, July 2021. CODEN ???? ISSN 2332-7790.

Zucker:2021:LSB

[ZPMT^{+}21] Jeremy Zucker, Kaushal Paneri, Sara Mohammad-Taheri, Somya Bhargava, Pallavi Kolambkar, Craig Bakker, Jeremy Teuton, Charles Tapley Hoyt, Kristie Oxford, Robert Ness, and Olga Vitek. Leveraging structured biological knowledge for counterfactual inference: a case study of viral pathogenesis. IEEE Transactions on Big Data, 7(1):25–37, March 2021. CODEN ???? ISSN 2332-7790.

Zhang:2023:PEC

[ZPT^{+}23] Zongshun Zhang, Andrea Pinto, Valeria Turina, Flavio Esposito, and Ibrahim Matta. Privacy and efficiency of communications in federated split learning. IEEE Transactions on Big Data, 9(5):1380–1391, October 2023. ISSN 2332-7790.

Zhang:2022:MSM

[ZQD^{+}22] Xuyun Zhang, Lianyong Qi, Wanchun Dou, Qiang He, Christopher Leckie, Ramamohanarao Kotagiri, and Zoran Salcic. MRMondrian: Scalable multidimensional anonymisation for big data privacy preservation. IEEE Transactions on Big Data, 8(1):125–139, February 2022. CODEN ???? ISSN 2332-7790.

Zhao:2017:SRP

[ZQK17] G. Zhao, X. Qian, and C. Kang.

Service rating prediction by exploring social mobile users geographical locations. IEEE Transactions on Big Data, 3(1): 67–78, March 2017. CODEN ???? ISSN 2332-7790.

Zhao:2017:TEF

[ZQZ^{+}17] D. Zhao, K. Qiao, Z. Zhou, T. Li, Z. Lu, and X. Xu. Toward efficient and flexible metadata indexing of big data systems. IEEE Transactions on Big Data, 3(1):107–117, March 2017. CODEN ???? ISSN 2332-7790.

Zhang:2022:SES

[ZRL22] Songnian Zhang, Suprio Ray, and Rongxing Lu. SOREL: Efficient and secure ORE-based range query over outsourced data. IEEE Transactions on Big Data, 8(6):1702–1715, December 2022. ISSN 2332-7790.

Zhang:2023:ELS

[ZRLZ23] Songnian Zhang, Suprio Ray, Rongxing Lu, and Yandong Zheng. Efficient learned spatial index with interpolation function based learned model. IEEE Transactions on Big Data, 9 (2):733–745, April 2023. ISSN 2332-7790.

Zhou:2020:OTD

[ZRY^{+}20] Xun Zhou, Huigui Rong, Chang Yang, Qun Zhang, Amin Vahedian Khezerlou, Hui Zheng, Zubair Shafiq, and Alex X. Liu. Optimizing taxi driver profit efficiency: a spatial networkbased Markov decision process approach. IEEE Transactions on Big Data, 6(1):145– 158, March 2020. CODEN ???? ISSN 2332-7790.

Zhang:2022:RRC

[ZS22] Yaying Zhang and Xinyuan Sui. RCIVMM: a route choice-based interactive voting map matching approach for complex urban road networks. IEEE Transactions on Big Data, 8(5):1414– 1427, October 2022. ISSN 2332-7790.

Zheng:2022:MPB

[ZSC^{+}22] Yuyan Zheng, Chuan Shi, Xiaohuan Cao, Xiaoli Li, and Bin Wu. A meta path based method for entity set expansion in knowledge graph. IEEE Transactions on Big Data, 8(3): 616–629, June 2022. ISSN 2332-7790.

Zhu:2017:LRG

[ZSHS17] X. Zhu, H. Suk, H. Huang, and D. Shen. Low-rank graphregularized structured sparse regression for identifying genetic biomarkers. IEEE Transactions on Big Data, 3(4):405– 414, December 2017. CODEN ???? ISSN 2332-7790.

Zheng:2023:POC

[ZSJ23] Zhigao Zheng, Xuanhua Shi, and Hai Jin. Parallel overlapping community detection algorithm on GPU. IEEE Transactions on Big Data, 9(2):677– 687, April 2023. ISSN 2332-7790.

Zhang:2019:QEB

[ZSK19] Peng Zhang, Xiang Shi, and Samee U. Khan. QuantCloud: Enabling Big Data complex event processing for quantitative finance through a datadriven execution. IEEE Transactions on Big Data, 5(4):564– 575, December 2019. CODEN ???? ISSN 2332-7790.

Zhu:2017:EST

[ZSL17] J. Y. Zhu, C. Sun, and V. O. K.

Li. An extended spatiotemporal Granger causality model for air quality estimation with heterogeneous urban big data. IEEE Transactions on Big Data, 3(3):307– 319, September 2017. CODEN ???? ISSN 2332-7790.

Zhang:2022:TUI

[ZSL^{+}22] Zongpu Zhang, Tao Song, Liwei Lin, Yang Hua, Xufeng He, Zhengui Xue, Ruhui Ma, and Haibing Guan. Towards ubiquitous intelligent computing: Heterogeneous distributed deep neural networks. IEEE Transactions on Big Data, 8(3): 644–657, June 2022. ISSN 2332-7790.

Zhang:2021:UOT

[ZSS^{+}21] Wei Emma Zhang, Ali Shemshadi, Quan Z. Sheng, Yongrui Qin, Xiujuan Xu, and Jian Yang. A user-oriented taxi ridesharing system with large-scale urban GPS sensor data. IEEE Transactions on Big Data, 7(2):327– 340, June 2021. CODEN ???? ISSN 2332-7790.

Zhang:2019:SAM

[ZSWZ19] Zhiyong Zhang, Ranran Sun, Xiaoxue Wang, and Changwei Zhao. A situational analytic method for user behavior pattern in multimedia social networks. IEEE Transactions on Big Data, 5(4):520–528, December 2019. CODEN ???? ISSN 2332-7790.

Zheng:2016:VAU

[ZWC^{+}16] Y. Zheng, W. Wu, Y. Chen, H. Qu, and L. M. Ni. Visual analytics in urban computing: An overview. IEEE Transactions on Big Data, 2(3):276– 296, September 2016. CODEN ???? ISSN 2332-7790.

Zhao:2023:BBA

[ZWD^{+}23] Shu Zhao, Wenyu Wang, Ziwei Du, Jie Chen, and Zhen Duan. A black-box adversarial attack method via Nesterov accelerated gradient and rewiring towards attacking graph neural networks. IEEE Transactions on Big Data, 9(6):1586–1597, December 2023. ISSN 2332-7790.

Zhang:2022:UCD

[ZWH^{+}22] Yanghua Zhang, Jice Wang, Hexiang Huang, Yuqing Zhang, and Peng Liu. Understanding and conquering the difficulties in identifying third-party libraries from millions of Android apps. IEEE Transactions on Big Data, 8(6):1511–1523, December 2022. ISSN 2332-7790.

Zhao:2022:MNG

[ZWLN22] Wan-Lei Zhao, Hui Wang, Peng-Cheng Lin, and Chong-Wah Ngo. On the merge of k-NN graph. IEEE Transactions on Big Data, 8(6):1496–1510, December 2022. ISSN 2332-7790.

Zhou:2021:MDB

[ZWS21] Zhili Zhou, Q. M. Jonathan Wu, and Xingming Sun. Multiple distance-based coding: Toward scalable feature matching for large-scale Web image search. IEEE Transactions on Big Data, 7(3):559–573, July 2021. CODEN ???? ISSN 2332-7790.

Zhang:2022:HGA

[ZWS^{+}22] Yiding Zhang, Xiao Wang, Chuan Shi, Xunqiang Jiang, and Yanfang Ye. Hyperbolic graph attention network. IEEE Transactions on Big Data, 8 (6):1690–1701, December 2022. ISSN 2332-7790.

Zhang:2019:SRT

[ZWV^{+}19] D. Zhang, D. Wang, N. Vance, Y. Zhang, and S. Mike. On scalable and robust truth discovery in big data social media sensing applications. IEEE Transactions on Big Data, 5(2):195– 208, June 2019. CODEN ???? ISSN 2332-7790.

Zhang:2022:PRP

[ZWZ^{+}22] Lei Zhang, Xinpeng Wu, Hongke Zhao, Fan Cheng, and Qi Liu. Personalized recommendation in P2P lending based on risk-return management: a multi-objective perspective. IEEE Transactions on Big Data, 8(4):1141–1154, August 2022. ISSN 2332-7790.

Zhang:2023:SAG

[ZWZ^{+}23] Yong Zhang, Xiulan Wei, Xinyu Zhang, Yongli Hu, and Baocai Yin. Self-attention graph convolution residual network for traffic data completion. IEEE Transactions on Big Data, 9 (2):528–541, April 2023. ISSN 2332-7790.

Zhao:2016:DGP

[ZYB^{+}16] Y. Zhao, K. Yoshigoe, J. Bian, M. Xie, Z. Xue, and Y. Feng. A distributed graph-parallel computing system with lightweight communication overhead. IEEE Transactions on Big Data, 2(3): 204–218, September 2016. CO-DEN ???? ISSN 2332-7790.

Zhang:2020:IDC

[ZYCL20] Qingchen Zhang, Laurence T.

Yang, Zhikui Chen, and Peng Li. Incremental deep computation model for wireless Big Data feature learning. IEEE Transactions on Big Data, 6(2):248– 257, June 2020. CODEN ???? ISSN 2332-7790.

Zhang:2022:PPP

[ZYCL22] Qingchen Zhang, Laurence T.

Yang, Zhikui Chen, and Peng Li. PPHOPCM: Privacypreserving high-order possibilistic c-means algorithm for big data clustering with cloud computing. IEEE Transactions on Big Data, 8(1):25–34, February 2022. CODEN ???? ISSN 2332-7790.

Zhu:2023:MDA

[ZYW^{+}23] Qi Zhu, Qiming Yang, Mingming Wang, Xiangyu Xu, Yuwu Lu, Wei Shao, and Daoqiang Zhang. Multi-discriminator active adversarial network for multi-center brain disease diagnosis. IEEE Transactions on Big Data, 9(6):1575–1585, December 2023. ISSN 2332-7790.

Zhou:2023:LMF

[ZYX^{+}23] Hao Zhou, Geng Yang, Yang Xiang, Yunlu Bai, and Weiya Wang. A lightweight matrix factorization for recommendation with local differential privacy in big data. IEEE Transactions on Big Data, 9(1):160–173, February 2023. ISSN 2332-7790.

Zhang:2018:QBD

[ZYYK18] P. Zhang, K. Yu, J. J. Yu, and S. U. Khan. Quant-Cloud: Big data infrastructure for quantitative finance on the cloud. IEEE Transactions on Big Data, 4(3):368– 380, September 2018. CODEN ???? ISSN 2332-7790.

Zhao:2023:MTS

[ZYZ^{+}23] Qinpei Zhao, Guangda Yang, Kai Zhao, Jiaming Yin, Weixiong Rao, and Lei Chen. Multivariate time-series forecasting model: Predictability analysis and empirical study. IEEE Transactions on Big Data, 9 (6):1536–1548, December 2023. ISSN 2332-7790.

Zhang:2020:NRL

[ZYZZ20] Daokun Zhang, Jie Yin, Xingquan Zhu, and Chengqi Zhang. Network representation learning: a survey. IEEE Transactions on Big Data, 6(1):3– 28, March 2020. CODEN ???? ISSN 2332-7790.

Zheng:2023:DDS

[ZZ23] Qi Zheng and Yaying Zhang. DSTAGCN: Dynamic spatialtemporal adjacent graph convolutional network for traffic forecasting. IEEE Transactions on Big Data, 9(1):241–253, February 2023. ISSN 2332-7790.

Zhang:2022:EDG

[ZZL^{+}22] Fan Zhang, Long Zheng, Xiaofei Liao, Xinqiao Lv, Hai Jin, and Jiang Xiao. An effective 2-dimension graph partitioning for work stealing assisted graph processing on multi-FPGAs. IEEE Transactions on Big Data, 8(5):1247–1258, October 2022. ISSN 2332-7790.

Zhou:2024:OHS

[ZZL^{+}24] Peng Zhou, Yunyun Zhang, Zhaolong Ling, Yuanting Yan, Shu Zhao, and Xindong Wu. Online heterogeneous streaming feature selection without feature type information. IEEE Transactions on Big Data, 10 (4):470–485, August 2024. ISSN 2332-7790.

Zhang:2021:SUC

[ZZLW21] Daniel Zhang, Yang Zhang, Qi Li, and Dong Wang. Sparse user check-in venue prediction by exploring latent decision contexts from location-based social networks. IEEE Transactions on Big Data, 7(5):859– 872, November 2021. CODEN ???? ISSN 2332-7790.

Zaghloul:2020:PMS

[ZZR20] Ehab Zaghloul, Kai Zhou, and Jian Ren. P-MOD: Secure privilege-based multilevel organizational data-sharing in cloud computing. IEEE Transactions on Big Data, 6(4):804–815, December 2020. CODEN ???? ISSN 2332-7790.

Zhang:2021:COM

[ZZX^{+}21] Yushu Zhang, Jiantao Zhou, Yong Xiang, Leo Yu Zhang, Fei Chen, Shaoning Pang, and Xiaofeng Liao. Computation outsourcing meets lossy channel: Secure sparse robustness decoding service in multiclouds. IEEE Transactions on Big Data, 7(4):717–728, April 2021. CODEN ???? ISSN 2332-7790.

Zheng:2023:MOM

[ZZY^{+}23] Furong Zheng, Juanjuan Zhao, Jiexia Ye, Xitong Gao, Kejiang Ye, and Chengzhong Xu. Metro OD matrix prediction based on multi-view passenger flow evolution trend modeling. IEEE Transactions on Big Data, 9 (3):991–1003, June 2023. ISSN 2332-7790.

Zhu:2018:PCI

[ZZZ^{+}18] J. Y. Zhu, C. Zhang, H. Zhang, S. Zhi, V. O. K. Li, J. Han, and Y. Zheng. pg-causality: Identifying spatiotemporal causal pathways for air pollutants with urban big data. IEEE Transactions on Big Data, 4(4):571– 585, December 2018. CODEN ???? ISSN 2332-7790.

Zou:2022:LTO

[ZZZ^{+}22] Xiexin Zou, Shiyao Zhang, Chenhan Zhang, James J. Q. Yu, and Edward Chung. Longterm origin-destination demand prediction with graph deep learning. IEEE Transactions on Big Data, 8(6):1481–1495, December 2022. ISSN 2332-7790.